### v1

# Private From Whom? Minimal Information Leakage in Auctions

Abstract
In many auctions, bidders may be reluctant to reveal private information to the auctioneer and other bidders. Among deterministic bilateral communication protocols, reducing what bidders learn requires increasing what the auctioneer learns. A protocol implementing a given social choice rule is on the privacy Pareto frontier if no alternative protocol reveals less to both bidders and the auctioneer. For first-price auctions, the descending protocol and the sealed-bid protocol are both on the privacy Pareto frontier. For second-price auctions, the ascending protocol and the ascending-join protocol of haupt2025contextually are both on the privacy Pareto frontier, but the sealed-bid protocol is not. A designer can flexibly trade off between what bidders learn and what the auctioneer learns by “stitching” different protocols together.
Keywords: Extensive-Form Mechanism Design, Privacy, Auctions
JEL Codes: D44, D47, D82.
1 Introduction
In many high-stakes economic settings, unwanted disclosure of agents’ private information can be extremely costly. In the 1960s, the Walt Disney Company extensively used shell corporations to conceal its valuation for the land it purchased to build Walt Disney World.222“[Disney] used aliases and code names, as well as shell and dummy corporations like Project Winter, Project X and Project Future. If word leaked that Walt Disney was buying up huge tracts, prices would shoot sky-high, dooming the project. Eighteen months of secrecy to buy up 27,400 acres.” (Payne 2021). In FCC Spectrum Auction 66, SpectrumCo successfully used early bids to estimate rival bidders’ budgets, allowing SpectrumCo to adjust its auction strategy and save billions of dollars. [^3] In 2006, the Federal Communications Commission conducted Auction 66: Advanced Wireless Services (AWS), netting over 13 billion dollars in revenue. In round nine, “SpectrumCo made a jump bid doubling the prices of the large REAG licenses covering the Northeast and Western United States. This move, which we referred to at the time as the ‘shake-out tactic’ was intended to resolve competitive uncertainty and favorably align relative prices in the auction.” Later on in “round 15, it was possible to forecast that cumulative high bids on the REAG licenses were so high relative to the total budgets in the auction that the smaller licenses would sell for a steep discount. This allowed SpectrumCo, alone among the major bidders, to make an early switch to the smaller licenses.” Such analysis hinged on observing other bidders’ behavior in early rounds: “total winning bids will be forecast early and well by total exposure.” SpectrumCo was able to save over \$1.1 billion to acquire licenses covering the United States, compared to the average price paid by its competitors. (bulow2009winning). As rothkopf1990vickrey and milgrom2020clock describe, bidders may be concerned that their bids reveal their technology level to rivals or reduce the bidder’s leverage in post-auction negotiations; they highlight the importance of privacy in auctions for construction contracts, oil leases, and spectrum. [^4] See rothkopf1990vickrey: “People of our acquaintance with experience in conducting business are reluctant to reveal their true costs or valuations. … A truth-revealing strategy may give away valuable information. It could reveal to potential competitors the extent to which a firm’s technology was superior. Most important, it could reveal to others with whom the firm must subsequently negotiate precisely how much it can yield. … Successful oil lease bidders must deal with drilling contractors, rig owners, and so forth. Successful coal lease bidders must deal with equipment suppliers, railroads, and coal purchasers. Successful construction contract bidders must deal with subcontractors and labor unions.” See milgrom2020clock on spectrum auctions: “The practical significance of UWP [Unconditional Winner Privacy] is that it may alleviate winners’ concerns about misuse of the revealed information, encourage participation by bidders who find it costly to figure out their exact values, and allow the auctioneer to conceal winners’ politically sensitive windfalls.” Faced with these concerns, a designer may wish to choose an auction protocol that preserves bidders’ privacy from both the auctioneer and other bidders. Can any auction format preserve privacy in both dimensions? We identify a stark tradeoff between these objectives: For most choice rules, either the auctioneer or bidders must learn unnecessary information.
We consider a designer, external to the interaction, who constructs a deterministic bilateral communication protocol among an auctioneer and multiple bidders to implement a given choice function. The auctioneer learns information about bidders’ types by sequentially communicating with them, which partitions the space of type profiles. Bidders also learn about other bidders’ types: a bidder’s observed history is only consistent with some realization of other bidders’ types, yielding an analogous partition. Protocol $P$ reveals more to the auctioneer than protocol $P^{\prime}$ if the auctioneer learns a finer partition of the type space under protocol $P$ than protocol $P^{\prime}$. Similarly, protocol $P$ reveals more to bidder $i$ than protocol $P^{\prime}$ if bidder $i$ learns a finer partition under $P$ than $P^{\prime}$.
Under a general condition (the indistinguishable corners condition, a weakening of the corners condition previously seen in the literature), we show that the designer faces a tradeoff between protecting bidders’ privacy from one another or from the auctioneer. Intuitively, to protect bidders’ privacy from the auctioneer, the protocol must refrain from asking bidders questions about their type when doing so is unnecessary to compute the outcome. However, which questions are necessary depends on what the auctioneer knows so far; by conditioning such questions on other bidders’ types, bidders infer what the auctioneer has learned and, hence, other bidders’ private information.
We then use this framework to compare the privacy tradeoffs between different protocols. Protocol $P$ privacy dominates protocol $P^{\prime}$ if protocol $P$ reveals less to both the auctioneer and other bidders than protocol $P^{\prime}$ does, with at least one of these holding strictly. Under these orderings, a protocol lies on the privacy Pareto frontier if no protocol that implements the same underlying choice rule privacy dominates it. We then apply our results to first-price and second-price auctions.
In a first-price auction setting, the sealed-bid first-price auction reveals more information to the auctioneer than the descending auction, but reveals less information to bidders. In a sealed-bid first-price auction, the auctioneer learns every bidder’s value, but each bidder learns only whether he won or lost. In a descending auction, the auctioneer learns less information by no longer querying any bidder once some bidder claims the good. However, each losing bidder now learns the winning bidder’s value based on the final question the auctioneer asks them before the auction ends. We show that both protocols lie on the privacy Pareto frontier.
In a second-price auction setting, the ascending-join protocol, proposed by haupt2025contextually, is one protocol that reveals minimal information to the auctioneer. However, other protocols reveal less information to bidders. In particular, the ascending-join protocol does not privacy-dominate the standard ascending protocol. This complicates the privacy rationale for using the ascending-join protocol: a privacy-conscious designer may prefer the ascending protocol. Privacy dominance still rules out some protocols: in particular, the ascending auction privacy dominates the sealed-bid second-price auction. We then quantify violations of the indistinguishable corners condition to show that the ascending auction lies on the privacy Pareto frontier.
Our results highlight that the designer often must trade off between revealing minimal information to the auctioneer and revealing minimal information to bidders. The designer’s preferred choice may depend on the specific interaction taking place. In some settings, such as a one-time auction for a procurement contract, preventing bidders from learning about rival bidders’ technological capabilities may be more important. In other settings, such as repeated online ad auctions (in which the auctioneer may set future reserve prices based on bidders’ values), preserving privacy from the auctioneer may be more important. As such, privacy concerns alone may not distinguish among the auction formats that lie on the frontier: protocols that initially seem to be “more private” may actually release more private information to participants. The designer may also value other criteria, such as preferring a static protocol (which requires less time to run) or an obviously strategyproof protocol (which requires less reasoning from participants). We do not take a stand on balancing these criteria with privacy concerns.
Our framework restricts the usage of alternative tools such as cryptography or an external mediator. [^5] Formally, our designer does not rely on a trusted randomization device, as in izmalkov2011perfect, we make no computational hardness assumptions, and assume that the designer also chooses bidder strategies leading to them being invertible. Alternative mechanisms come with their own drawbacks. [^6] Participants must understand or trust that a cryptographic scheme works as intended for the mechanism to be successful. Many alternative mechanisms also require players to communicate in a different manner, such as through an underlying blockchain or by encoding messages in specific ways themselves. Relying on a separate auctioneer or platform requires bidders to trust that the auction host will not divulge any information that they learn. Finally, alternative tools may not exist in certain settings (many auctions are run in certain ways due to historical significance), may become obsolete (for example, if quantum computers are able to break cryptographic schemes), or are too costly for other reasons (in spectrum auctions, where a graph coloring problem must be solved quickly in every iteration, any increase in time complexity could be prohibitive). With that in mind, our results provide a calculus for deciding when the costs of moving to an alternative scheme are worthwhile. A designer may first wish to understand what privacy protections are feasible in the absence of cryptography before deciding whether to use cryptographic protocols; our work provides this characterization.
The remainder of the paper proceeds as follows. Section˜2 reviews related literature. Section˜3 introduces our model, presents our notions of bidder and auctioneer privacy, and comments on the incentive properties of protocols. Section˜4 contains our main characterizations: an impossibility theorem identifying the tradeoff between bidders and the auctioneer learning unnecessary information, and a stitching lemma demonstrating that the tradeoff is flexible. Section˜5 and Section˜6 apply our results to first and second-price auctions respectively, situating common auction formats relative to the frontier. Section˜7 concludes.
2 Related Literature
Our work is related to the notions of unconditional privacy (which has the strict requirement of zero privacy violations) and contextual privacy (which only analyzes privacy from the auctioneer). brandt2008privacy consider implementing auctions via bilateral communication among bidders without revealing any information, a condition known as unconditional privacy (and considered earlier by chor1991boolean). We instead consider settings with an auctioneer and characterize privacy tradeoffs even when unconditional privacy is unattainable. milgrom2020clock and ausubel2004auction consider the related notion of unconditional winner privacy, which reveals no unnecessary information about the winner’s value; in contrast, we consider the privacy of all bidders. haupt2025contextually characterizes protocols that implement choice rules while revealing minimal unnecessary information to the auctioneer, which they call contextual privacy. We show that contextual privacy typically conflicts with the desire to reveal minimal unnecessary information to bidders. In turn, protocols which are not maximally contextually private may be better at protecting bidders’ privacy, leading us to consider a distinct privacy Pareto frontier. Finally, Liu2025WorstCasePrivacy quantifies the worst-case privacy risk of information structures and studies optimal mechanisms subject to a “budget” of privacy violations. We instead study the maximally privacy-preserving way to implement a given allocation rule. Additionally, Liu2025WorstCasePrivacy considers inferences an outside observer can make in non-auction settings, which differs from information leakages to the bidders and the auctioneer (as they gain additional information by participating in the auction itself) and are a larger concern in auctions.
A broader literature considers protocols that use trusted randomization or computational hardness assumptions. izmalkov2011perfect provide a way to implement any choice rule in the presence of a trusted randomization device that fully preserves privacy. [^7] The implementation in izmalkov2011perfect is, informally speaking, somewhat complex. This may be another motivation for the designer to choose protocols that do not rely on trusted randomization. In contrast, we restrict attention to protocols that do not require trusted randomization, and we study the tradeoffs between protocols commonly used in practice. This provides a benchmark for what privacy protections are possible without resorting to cryptographic protocols. See alvarez2020survey for a survey of cryptographic protocols in auctions and see bogetoft2009sugarbeets for applications to an agricultural auction.
Other literature discusses the importance of bidders’ privacy in real-world settings, typically nesting a mechanism in a larger model. rothkopf1990vickrey consider bidders who are reluctant to reveal their true valuations in Vickrey auctions due to post-auction bargaining. milgrom2020clock, ausubel2004auction and mcmillan1994spectrum discuss the importance of protecting winner privacy in spectrum auctions and multi-unit auctions. dworczak2020aftermarkets models an aftermarket which is affected by the information revealed in an auction. We are motivated by similar concerns as this literature, and we provide a framework to compare whose privacy is compromised by a given protocol, and who their information is revealed to.
Finally, our work relates to the literature on mechanism design with partial commitment. AkbarpourLi2020CredibleAuctions and komo2025shill consider auctioneer deviations from a specified mechanism: Whether such deviations are profitable are informed by what the auctioneer learns while whether such deviations are detectable are disciplined by what bidders learn. grigoryan2023auditability propose an index to measure the extent to which agents can detect deviations from the mechanism. In contrast to this literature, we consider an auctioneer who faithfully executes a specified protocol: Privacy concerns persist even if the auctioneer can commit to the protocol used.
3 Model
3.1 Primitives
Our model resembles that of haupt2025contextually and the messaging game of AkbarpourLi2020CredibleAuctions. To analyze bidder privacy, we utilize the notion of an bidder’s experience developed by AkbarpourLi2020CredibleAuctions. Our characterizations are type profile by type profile and abstract away from (Bayesian) incentive constraints, so we do not need to work with beliefs. The model consists of the following:
  1. 1.
    A finite set of bidders $N=\{1,2,...,n\}$ with generic element $i$.
  2. 2.
    An auctioneer $A$ that communicates with the bidders and chooses the final outcome.
  3. 3.
    A finite type space $\bm{\Theta}=\times_{i\in N}\Theta_{i}$.
  4. 4.
    A set of outcomes $X$.
  5. 5.
    Bidder utilities $u_{i}:X\times\Theta_{i}\to\mathbb{R}$.
  6. 6.
    A partition $\Omega_{i}$ of $X$, describing which outcomes bidder $i$ can inherently distinguish. [^8] In auction settings, we assume that a bidder observes their allocation and payment at the conclusion of the auction. Suppose $\{(q_{i},t_{i})\}_{i=1}^{N}\in X$ represents an outcome with allocation $q_{i}\in\{0,1\}$ and transfer $t_{i}$ for bidder $i$. Then $\{(q_{i},t_{i})\}_{i=1}^{N}\sim_{\Omega_{i}}\{(q_{i}^{\prime},t_{i}^{\prime})\}_{i=1}^{N}$ if $q_{i}=q_{i}^{\prime}$ and $t_{i}=t_{i}^{\prime}$. For example, bidder $i$ cannot inherently distinguish between two different outcomes which both result in $i$ losing the auction.
  7. 7.
    A social choice rule $\phi:\bm{\Theta}\to X$ that the designer would like to implement.
We consider a fixed choice rule $\phi$, such as a second-price auction rule or a first-price auction rule. The designer may account for many tradeoffs in choosing $\phi$ (such as efficiency, revenue, or desirable incentive properties). For a given choice rule $\phi$, the designer still has considerable latitude to choose an extensive-form implementation of $\phi$; we describe this frontier of implementing the chosen choice rule.
We model the designer as external to the interaction between bidders and the auctioneer. In many mechanism design settings, the economist that advises organizations on what mechanism to use is not a participant themselves, and simply seeks to identify tradeoffs, bounds on what is inherently feasible, and how to best achieve certain objectives. To implement $\phi$, the designer chooses an extensive-form game $G=(H,i,\{\mathcal{I}\}_{i\in N\cup\{A\}},\{\mathcal{A}_{i}\}_{i\in N\cup\{A\}},\mathcal{A},I)$ of finite length and perfect recall with notation specified as follows:
Table 1: Notation for Extensive-Form Games
Name Notation Representative Element
Histories $H$ $h$
Player called to play at $h$ $i(h)$
Information sets of player $i$ $\mathcal{I}_{i}$ $I_{i}$
Actions available to $i$ $A_{i}$
Actions available at $I_{i}$ $\mathcal{A}(I_{i})$ $a$
Information set at history $h$ $I(h)$
Terminal histories $Z$ $z$
Outcome at terminal node $g(z)$
Interim strategy of player $i$ $\sigma_{i}(\mathcal{I}_{i})$
Following AkbarpourLi2020CredibleAuctions, a game $G$ generates a bilateral communication game for the auctioneer to implement the choice function by repeatedly querying successive bidders and observing their actions before moving to successor information sets until choosing an outcome at a terminal history. [^9] We can formally write down this bilateral communication game as follows: Let the auctioneer’s message space be the union of all players’ information sets, $\cup_{i}\mathcal{I}_{i}$. Then at each stage, according to $G$, 1. The auctioneer either selects an outcome (finishing the game) or proceeds to step 2. 2. The auctioneer sends message $I_{i}\in\mathcal{I}_{i}$ to bidder $i$. 3. Bidder $i$ observes $I_{i}$ and chooses action $a\in A(I_{i})$. 4. The auctioneer observes $a$. 5. Return to step 1. This setup embeds an assumption that the auctioneer cannot “commit to forget” the content of a bidder’s message.
We limit our attention to deterministic strategy profiles. An interim strategy for bidder $i$ is a map $\sigma_{i}:\mathcal{I}_{i}\to A_{i}$ satisfying $\sigma_{i}(I_{i})\in A(I_{i})$. Let $\Sigma_{i}$ denote the space of interim strategies for bidder $i$. An ex-ante strategy for bidder $i$ is a map $S_{i}:\Theta_{i}\to\Sigma_{i}$ mapping from types to interim strategies. A protocol $P$ is a pair $P=(G,\bm{S})$ where $\bm{S}=\{S_{i}\}_{i\in N}$ is some fixed profile of deterministic strategies for $G$.
Each type profile $\bm{\theta}$ yields a sequence of information sets and bidder actions at those information sets. Let $I_{i}^{k}(\bm{\theta})$ denote the $k$’th information set reached at which bidder $i$ is called to play, and let $a_{i}^{k}(\bm{\theta})$ denote their action chosen at this information set. Let $I^{k}(\bm{\theta})$ denote the $k$’th information set reached in the game, and let $a^{k}(\bm{\theta})$ denote the action chosen at this information set. Let $z(\{\sigma_{i}\})$ denote the terminal node reached via playing interim strategies $\{\sigma_{i}\}$ so $z(\bm{S}(\bm{(}\theta))$ is the terminal reached via playing ex-ante strategies $\bm{S}$ at type profile $\bm{\theta}$; protocol $P$ implements social choice rule $\phi$ if for all $\bm{\theta}$, $\phi(\bm{\theta})=g(z(\bm{S}(\bm{\theta}))).$
What do the bidders and auctioneer learn along the path of play of the mechanism? The auctioneer observes the full sequence of information sets reached and actions taken during the play of the game: For a given protocol $P$ and type profile $\bm{\theta}$, the auctioneer observes the game history $(I^{k}(\bm{\theta}),a^{k}(\bm{\theta}))_{k=1}^{T}$. Let $o_{A}(P,\bm{\theta})$ denote this observed history under protocol $P$ at type profile $\theta$. We say $\bm{\theta}\sim_{(P,A)}\bm{\theta}^{\prime}$ if $o_{A}(P,\bm{\theta})=o_{A}(P,\bm{\theta}^{\prime})$: If the auctioneer cannot distinguish between type profiles $\bm{\theta}$ and $\bm{\theta}^{\prime}$, then they are equivalent under $\sim_{(P,A)}$. To implement $\phi$, the auctioneer must at least be able to distinguish between type profiles that prescribe different outcomes. Letting $\mathcal{P}(\sim)$ denote the partition induced by an equivalence relation $\sim$, this condition can be stated in the language of partitions: $\mathcal{P}(\sim_{(P,A)})$ must be finer than $\mathcal{P}(\sim_{\phi})$ where $a\sim_{\phi}b$ if and only if $\phi(a)=\phi(b)$. [^10] Elements $a,b\in\Theta$ are defined to be in the same cell of $\mathcal{P}(\sim)$ if and only if $a\sim b$. Partition $\mathcal{P}$ is finer than partition $\mathcal{P}^{\prime}$ if for any cell $X$ in $\mathcal{P}$, there exists a cell $Y$ in $\mathcal{P}^{\prime}$ such that $X\subseteq Y$.
What about bidders? Bidder $i$ observes $(\theta_{i},(I_{i}^{k}(\bm{\theta}),a_{i}^{k}(\bm{\theta}))_{k=1}^{T_{i}},\omega_{i})$, which is the history of the information sets at which he has been called to play, the actions he has taken at those information sets, and what she observes about the final outcome, $\omega_{i}\in\Omega_{i}$. [^11] Under this formulation, the bidder can observe their sequence of play, but not calendar time. This is also the definition of an experience, restricted to the end of the game, in li2025obvious. Let $o_{i}(P,\bm{\theta})$ denote bidder $i$’s observation in protocol $P$ when bidders have type profile $\bm{\theta}\in\bm{\Theta}$. We may then define what bidder $i$ learns about other bidders under protocol $P$ by defining the following equivalence relation. We say $\bm{\theta}\sim_{(P,i)}\bm{\theta}^{\prime}$ if $o_{i}(P,\bm{\theta})=o_{i}(P,\bm{\theta}^{\prime})$. That is, the bidder can distinguish between $\bm{\theta}$ and $\bm{\theta}^{\prime}$ in protocol $P$ if and only if they induce different experiences for bidder $i$. Similarly, $\sim_{(P,i)}$ induces a partition $\mathcal{P}(\sim_{(P,i)})$ over $\bm{\Theta}$ for $i$.
3.2 Privacy
Using the above setup, we can now introduce our notions of privacy.
Definition 1 (Bidder Privacy).
Protocol $P$ reveals (strictly) more to bidder $i$ than protocol $P^{\prime}$ if $\mathcal{P}(\sim_{(P,i)})$ is (strictly) finer than $\mathcal{P}(\sim_{(P^{\prime},i)})$.
Definition 2 (Auctioneer Privacy).
Protocol $P$ reveals (strictly) more to the auctioneer than protocol $P^{\prime}$ if $\mathcal{P}(\sim_{(P,A)})$ is (strictly) finer than $\mathcal{P}(\sim_{(G^{\prime},A)})$.
Bidder privacy for each $i$ and auctioneer privacy form their own partial orders on the space of all protocols in a given environment (identified with the usual ordering over partitions, restricted to the set of partitions induced by feasible protocols). With the competing objectives of bidder and auctioneer privacy, we can define notions of when a protocol is “better” than another and “best” from a pure privacy perspective.
Definition 3 (Privacy Domination).
A protocol $P$ is (strictly) privacy dominated by protocol $P^{\prime}$ if $P$ reveals more information to each bidder than $P^{\prime}$ and $P$ reveals more to the auctioneer than $P^{\prime}$ (with at least one strict).
Definition 4 (Privacy Pareto Frontier).
A protocol $P$ which implements $\phi$ lies on the privacy Pareto frontier for $\phi$ if there does not exist a protocol that implements $\phi$ which strictly privacy dominates $P$.
At the endpoints of this frontier are maximally auctioneer private and maximally bidder private protocols. Since partitions are only partially ordered by fineness, ordering protocols by bidder or auctioneer privacy is also a partial ordering. As such, the privacy Pareto frontier will generally have many protocols on the frontier being incomparable to one another.
We adopt the language of haupt2025contextually to further refine auctioneer privacy. Some privacy violations may be necessary to implement a given $\phi$, captured in the following:
Definition 5 (Contextual Privacy).
Protocol $P$ produces a contextual privacy violation for bidder $i$ at type profile $(\theta_{i},\bm{\theta}_{-i})$ if there exists $\theta_{i}^{\prime}\in\Theta_{i}$ such that
$(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})\text{ but }\phi(\theta_{i},\bm{\theta}_{-i})=\phi(\theta_{i}^{\prime},\bm{\theta}_{-i}).$
Protocols can be partially ordered via inclusion on the type profiles where contextual privacy violations occur. A protocol produces no contextual privacy violations for bidder $i$ if the auctioneer only learns from bidder $i$ exactly what is needed to compute the choice function. For many choice functions, no protocol produces no contextual privacy violations. However, there are still maximally contextually private protocols; as some violations are inevitable, the set of maximally contextually private protocols is often large, each producing contextual privacy violations at different type profiles. Maximal contextual privacy coincides with our notion of maximal auctioneer privacy: If $P$ reveals less to the auctioneer than $P^{\prime}$ then protocol $P$ must also induce less contextual privacy violations.
Definition 6 (Maximal auctioneer Privacy).
A protocol $P$ implementing $\phi$ is maximally contextually private if there does not exist a protocol $P^{\prime}$ implementing $\phi$ which reveals strictly less to the auctioneer than $P$.
In contrast, minimal bidder informativeness captures our notion of privacy from bidders. A minimally bidder-informative protocol reveals no information to any bidder beyond what the bidder can infer from the outcome itself (and observing their own type).
Definition 7 (Minimal Bidder-Informativeness).
A protocol $P$ implementing $\phi$ is minimally bidder-informative if for all $i\in\mathcal{I}$ and $\bm{\theta},\bm{\theta}^{\prime}\in\bm{\Theta}$,
$\theta_{i}=\theta^{\prime}_{i}\text{ and }\phi(\bm{\theta})\sim_{\Omega_{i}}\phi(\bm{\theta}^{\prime})\text{ implies }\bm{\theta}\sim_{(P,i)}\bm{\theta}^{\prime}.$
For any choice function $\phi$, one way to attain minimal bidder-informativeness is via a sealed-bid protocol, in which each bidder directly reports their type to the auctioneer, and the auctioneer then chooses the outcome. Such a protocol reveals no information to bidders beyond what they observe of the outcome.
Definition 8 (Sealed-Bid Protocol).
In a sealed-bid protocol for $\phi$, each bidder $i$ has one possible information set and is queried to directly reveal their type: $\mathcal{I}_{i}=\{I_{i}\}$, and $A(I_{i})=\Theta_{i}$. At each terminal node, the designer selects the outcome $\phi(a_{1}(I_{1}),a_{2}(I_{2}),...,a_{n}(I_{n}))$.
Lemma 1 (Minimal Bidder-Informativeness is Attained).
For any choice function $\phi$, the sealed-bid protocol achieves minimal bidder-informativeness. No protocol $P^{\prime}$ can reveal strictly less to bidder $i$ than the sealed-bid protocol.
Strictly speaking, multiple sealed-bid protocols exist since the auctioneer can query bidders in any arbitrary order. These protocols are all equivalent, and under any sealed-bid protocol, the auctioneer perfectly learns bidders’ values, and bidders are minimally informed. As such, sealed-bid protocols are minimally contextually private, foreshadowing our impossibility results in the next section. Sealed-bid protocols also are not the unique minimally bidder-informative protocols, as we will show in the second-price auction setting. [^12] Protocols that enable bidders to make inferences may also be minimally bidder-informative if those inferences could have been made anyways based on what bidders observe about the outcome. In the second-price auction case, for example, ascending auctions are also minimally bidder-informative. As a result, a sealed-bid protocol may not lie on the privacy frontier, although it is minimally bidder-informative.
3.3 Pruning and Incentive Compatibility
We briefly comment on the incentive properties of protocols. If the choice function the designer seeks to implement is dominant-strategy incentive compatible, then it is without loss to restrict to protocols $P=(G,\bm{S})$ which are also dominant-strategy incentive compatible (that is, $\bm{S}$ is an equilibrium of $G$ in dominant strategies). Recall that $g(z)$ is the outcome at terminal history $z$ and $z(\bm{S})$ is the terminal history reached at strategies $S$.
Definition 9 (Incentive Compatibility).
A choice function $\phi$ is dominant-strategy incentive compatible if it is a dominant strategy for each bidder to truthfully reveal their type: for all bidders $i$, $\theta_{i},\theta_{i}^{\prime}\in\Theta_{i},\bm{\theta}_{-i}\in\bm{\Theta}_{-i}$,
$u_{i}(\phi(\theta_{i},\bm{\theta}_{-i}),\theta_{i})\geq u_{i}(\phi(\theta_{i}^{\prime},\bm{\theta}_{-i}),\theta_{i}).$
A protocol $P=(G,\{S_{i}(\theta_{i})\}_{i\in N})$ is dominant-strategy incentive compatible if it is a dominant strategy for each bidder to follow the designer’s strategy recommendation: for all bidders $i$, $\theta_{i}\in\Theta_{i},\bm{\theta}_{-i}\in\bm{\Theta}_{-i},\sigma_{i}^{\prime}\in\Sigma_{i}$,
$u_{i}(g(z(S_{i}(\theta_{i}),\bm{S}_{-i}(\bm{\theta}_{-i}))),\theta_{i})\geq u_{i}(g(z(\sigma_{i}^{\prime},\bm{S}_{-i}(\bm{\theta}_{-i}))),\theta_{i}).$
We will restrict attention to protocols where each history is reached at some type profile. Protocols where some histories are never reached at any type profile can be “pruned” to produce an alternative protocol such that the preceding condition is satisfied without changing privacy. Pruning also simplifies strategic analysis: At non-pruned protocols, deviations to histories that are never relevant may be profitable, but the designer can simply prune these branches of the game tree.
Definition 10 (Pruning).
[…]
  1. 1.
[…]
  2. 2.
    $\tilde{\mathcal{I}}_{i}=\{I\cap\tilde{H}\}_{I\in\mathcal{I}_{i}}$ for each $i\in N$.
  3. 3.
[…]
Given $P=(G,\bm{S})$ let $\tilde{P}=(\tilde{G},\bm{S})$ denote the protocol where $G$ is pruned and $\bm{S}$ remains unchanged.
Since $P$ implements $\phi$ on-path and $P=\tilde{P}$ for each realized path of play, $\tilde{P}$ implements $\phi$ as well. As such, it is without loss for the purpose of implementation to restrict to pruned protocols. It is also without loss for the purpose of analyzing privacy. Privacy violations are determined by on-path behavior so pruning off-path histories has no effect.
Proposition 1 (Pruning Preserved Privacy).
Let $\tilde{P}=(\tilde{G},\bm{S})$ be the protocol attained from pruning $P=(G,\bm{S})$. Then:
  1. 1.
    Auctioneer Privacy Equivalence: $\bm{\theta}\sim_{(P,A)}\bm{\theta}^{\prime}$ if and only if $\bm{\theta}\sim_{(\tilde{P},A)}\bm{\theta}^{\prime}$.
  2. 2.
    Bidder Privacy Equivalence: For each bidder $i$, $\bm{\theta}\sim_{(P,i)}\bm{\theta}^{\prime}$ if and only if $\bm{\theta}\sim_{(\tilde{P},i)}\bm{\theta}^{\prime}$.
This result is of independent interest beyond enabling analysis of incentive compatibility. It allows the designer to significantly simplify the search for protocols on the frontier to protocols where each history is reached at some path of play; in particular it places a finite bound on the maximal length of protocols which do not ask redundant questions.
In any pruned protocol, each history is reached on-path for some type, so deviations in the game $G$ are essentially deviations where a bidder pretends to be a different type. As such, incentive compatibility of the choice function is able to rule out all possible deviations.
Proposition 2 (Incentive Compatibility of Pruned Protocols).
Suppose $P=(G,\bm{S})$ is pruned and implements $\phi$. If $\phi$ is dominant-strategy incentive compatible, then $P$ is dominant-strategy incentive compatible.
Protocols do not necessarily preserve Bayesian incentive compatibility, however. A protocol could provide certain bidders with pertinent information before they make a choice. As such, interim beliefs may be different under the protocol implementing the mechanism compared to its normal form game, so Bayesian incentive compatibility does not translate. This may provide another novel justification for dominant strategy incentive compatibility: Dominant strategy incentive compatibility is robust to the details of how a mechanism is actually run whereas Bayesian incentive compatibility is not. Furthermore, our impossibility theorem only gets stronger if Bayesian incentive compatibility is required as it restricts the set of potential protocols the designer may use. In our application to first-price auctions, for which the choice function $\phi^{FPA}$ is not dominant-strategy incentive compatible, we show that certain auctions are not privacy dominated by any protocol implementing $\phi^{FPA}$. This immediately implies that these auctions are not privacy-dominated by any incentive-compatible protocol implementing $\phi^{FPA}$.
4 Characterizations
4.1 Impossibility Theorem
We provide a condition on the choice function $\phi$ under which any minimally bidder-informative protocol must distinguish type profiles that are not pivotal to the outcome. We begin by stating a condition analogous to the Corners Lemma of chor1991boolean, which is sufficient for non-existence of a fully contextually private protocol.
Definition 11.
A choice function $\phi$ violates the corners condition for bidder $i$ at type profiles $\{\theta_{i},\theta_{i}^{\prime}\}\times\{\bm{\theta}_{-i},\bm{\theta}_{-i}^{\prime}\}$ if
$\phi(\theta_{i},\bm{\theta}_{-i})=\phi(\theta_{i}^{\prime},\bm{\theta}_{-i})=\phi(\theta_{i},\bm{\theta}_{-i}^{\prime})\neq\phi(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime}).$
Figure˜1 displays a violation of the corners condition, with $\alpha,\gamma\in X$ representing the choice function outcome $\phi(\bm{\theta})$ at type profile $\bm{\theta}$. To implement $\phi$, the auctioneer must distinguish the top-right corner, but need not distinguish between the other three type profiles.
$\alpha$$\alpha$$\alpha$$\gamma$$\theta_{i}$$\theta_{i}^{\prime}$$\bm{\theta}_{-i}$$\bm{\theta}_{-i}^{\prime}$
Figure 1: Corners Condition Violation
As we described in the model section, bidders are often unable to distinguish distinct outcomes (such as different outcomes when they lose the auction). As such, the equality between $\phi(\theta_{i},\bm{\theta_{-i}})$ and $\phi(\theta_{i},\bm{\theta_{-i}}^{\prime})$ can be weakened to equivalence under $\sim_{\Omega_{i}}$. This motivates our indistinguishable corners condition, corresponding to cases where bidder $i$ cannot distinguish between three of the outcomes at the corners, but the auctioneer must distinguish the final corner to implement the choice function.
Definition 12.
A choice function $\phi$ violates the indistinguishable corners condition for bidder $i$ at type profiles $\{\theta_{i},\theta_{i}^{\prime}\}\times\{\bm{\theta}_{-i},\bm{\theta}_{-i}^{\prime}\}$ if:
  1. 1.
    $\theta_{i},\theta_{i}^{\prime}$ should not be distinguished at $\bm{\theta}_{-i}$:
    $\phi(\theta_{i},\bm{\theta_{-i}})=\phi(\theta_{i}^{\prime},\bm{\theta_{-i}});$
  2. 2.
    At $\theta_{i}$, $i$ should not distinguish between $\bm{\theta}_{-i},\bm{\theta}_{-i}^{\prime}$:
    $\phi(\theta_{i},\bm{\theta_{-i}})\sim_{\Omega_{i}}\phi(\theta_{i},\bm{\theta_{-i}}^{\prime});$
  3. 3.
    $\theta_{i},\theta_{i}^{\prime}$ needs to be distinguished at $\bm{\theta}_{-i}^{\prime}$:
    $\phi(\theta_{i},\bm{\theta_{-i}}^{\prime})\neq\phi(\theta_{i}^{\prime},\bm{\theta_{-i}}^{\prime}).$
We say that $\phi$ violates the indistinguishable corners condition for bidder $i$ if there exists some type profile where the above holds.
[…]
$\alpha$$\alpha$$\beta$$\gamma$$\theta_{i}$$\theta_{i}^{\prime}$$\bm{\theta}_{-i}$$\bm{\theta}_{-i}^{\prime}$
Figure 2: Indistinguishable Corners Condition Violation
Note that a violation of the (standard) corners condition is also a violation of the indistinguishable corners condition as $\phi(\theta_{i},\theta_{j},\bm{\theta}_{-ij})=\phi(\theta_{i},\theta_{j}^{\prime},\bm{\theta}_{-ij})$ implies $\phi(\theta_{i},\theta_{j},\bm{\theta}_{-ij})\sim_{\Omega_{i}}\phi(\theta_{i},\theta_{j}^{\prime},\bm{\theta}_{-ij})$ since $\sim_{\Omega_{i}}$ is defined solely as a partition over the outcome space (and each outcome is trivially in the same partition as itself). In many cases, there are strictly more violations of the indistinguishable corners condition than the corners condition. In the first-price auction, there are no violations of the corners condition but many violations of the indistinguishable corners condition (Lemma 5). In the second-price auction, there are violations of the corners condition, but also distinct violations of the indistinguishable corners condition (Lemma 7).
Suppose the designer implements some choice function $\phi$ that violates the indistinguishable corners condition. The driving force behind our impossibility theorem is that to prevent bidders from learning information about other bidders’ types, the auctioneer must ask superfluous questions, thus learning unnecessary information. That is, any minimally bidder-informative protocol which implements $\phi$ must distinguish type profile $(\theta_{i},\bm{\theta}_{-i})$ from type profile $(\theta_{i}^{\prime},\bm{\theta}_{-i})$.
Lemma 2.
Suppose $\phi$ violates the indistinguishable corners condition for bidder $i$ at $\{\theta_{i},\theta_{i}^{\prime}\}\times\{\bm{\theta}_{-i},\bm{\theta}_{-i}^{\prime}\}$. If protocol $P$ is minimally bidder-informative and implements $\phi$, then
$(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i}).$
As $\phi(\theta_{i},\bm{\theta_{-i}})=\phi(\theta_{i}^{\prime},\bm{\theta}_{-i})$, the auctioneer learns information which is not necessary to compute the outcome. For intuition, consider Figure˜2. Suppose bidder $i$ has type $\theta_{i}$. If other bidders have type profile $\bm{\theta}_{-i}^{\prime}$, then the auctioneer must eventually ask bidder $i$ to distinguish type $\theta_{i}$ and $\theta_{i}^{\prime}$, in order for the auctioneer to determine the outcome. This means that if the auctioneer never asks bidder $i$ to distinguish $\theta_{i}$ and $\theta_{i}^{\prime}$, then bidder $i$ understands the type profile is not $(\theta_{i},\bm{\theta}_{-i}^{\prime})$. This is additional information that bidder $i$ obtains from the protocol being run. A auctioneer can provide each bidder strictly less information by running a sealed-bid protocol instead. In such a sealed-bid protocol, bidder $i$ cannot distinguish between $\phi(\theta_{i},\bm{\theta}_{-i})$ and $\phi(\theta_{i},\bm{\theta}_{-i}^{\prime})$ based on neither their observation about the outcome nor their experience in the protocol.
To show Lemma˜2, we use the following lemma formalizing the intuition above: bidder $i$ learns about $\bm{\theta}_{-i}$ if the auctioneer only asks them questions for certain values of $\bm{\theta}_{-i}$.
Lemma 3.
Consider $\theta_{i},\theta_{i}^{\prime}\in\Theta_{i}$ and $\bm{\theta}_{-i},\bm{\theta}_{-i}^{\prime}$. If $(\theta_{i},\bm{\theta}_{-i})\sim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$ and $(\theta_{i},\bm{\theta}_{-i}^{\prime})\nsim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$ then $o_{i}(\theta_{i},\bm{\theta}_{-i})\neq o_{i}(\theta_{i},\bm{\theta}_{-i}^{\prime})$ and $o_{i}(\theta_{i}^{\prime},\bm{\theta}_{-i})\neq o_{i}(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$.
Proof.
(Lemma˜3) Because $(\theta_{i},\bm{\theta}_{-i}^{\prime})\nsim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$, there exists an information set $I_{i}$ on the path of play for $(\theta_{i},\bm{\theta}_{-i}^{\prime})$ at which $S_{i}(\theta_{i})(I_{i})\neq S_{i}(\theta_{i}^{\prime})(I_{i})$. Because $(\theta_{i},\bm{\theta}_{-i})\sim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$, at every information set $I_{i}$ on the path of play for $(\theta_{i},\bm{\theta}_{-i})$, we have $S_{i}(\theta_{i})(I_{i})=S_{i}(\theta_{i}^{\prime})(I_{i})$. Hence there exists some information set $I_{i}$ on the path of play for $(\theta_{i},\bm{\theta}_{-i}^{\prime})$ which is not on the path of play for $(\theta_{i},\bm{\theta}_{-i})$, and therefore $o_{i}(\theta_{i},\bm{\theta}_{-i})\neq o_{i}(\theta_{i},\bm{\theta}_{-i}^{\prime})$. By an identical argument, there exists some information set on the path of play for $(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$ which is not on the path of play for $(\theta_{i}^{\prime},\bm{\theta}_{-i})$, and therefore $o_{i}(\theta_{i}^{\prime},\bm{\theta}_{-i})\neq o_{i}(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$. ∎
We may then use Lemma˜3 to prove Lemma˜2.
Proof.
(Lemma˜2) By Lemma 1, in any minimally bidder-informative protocol, bidders learn nothing except what they observe of the outcome. Now suppose $G$ is minimally bidder-informative and implements $\phi$. Consider any $\theta_{i},\theta_{i}^{\prime}\in\Theta_{i}$ and $\bm{\theta}_{-i},\bm{\theta}_{-i}^{\prime}\in\bm{\Theta}_{-i}$ which violate the indistinguishable corners condition in Definition˜12. Because protocol $P$ implements $\phi$, we have $(\theta_{i},\bm{\theta}_{-i}^{\prime})\nsim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$. By the indistinguishable corners condition, $\phi(\theta_{i},\bm{\theta}_{-i})\sim_{\Omega_{i}}\phi(\theta_{i},\bm{\theta}_{-i}^{\prime})$, and hence by $P$ being minimally bidder-informative, $o_{i}(\theta_{i},\bm{\theta}_{-i})=o_{i}(\theta_{i},\bm{\theta}_{-i}^{\prime})$. By the contrapositive of Lemma˜3, we conclude that $(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$. ∎
This now gives us a set of types who must be “unnecessarily” queried in any minimally bidder-informative protocol. This inherently conflicts with protecting bidders’ privacy from the auctioneer.
Theorem 1 (Impossibility).
Suppose $\phi$ violates the indistinguishable corners condition for bidder $i$. No protocol $P$ that implements $\phi$ can satisfy both:
  1. (i)
    $P$ produces no contextual privacy violations for bidder $i$.
  2. (ii)
    $P$ is minimally bidder-informative.
The intuition is straightforward, given Lemma˜2: any minimally bidder-informative protocol must learn unnecessary information about bidder $i$, meaning the privacy of bidder $i$ is compromised.
Proof.
(Theorem˜1) Suppose $\phi$ violates the indistinguishable corners condition at some type profiles $\{\theta_{i},\theta_{i}^{\prime}\}\times\{\bm{\theta}_{-i},\bm{\theta}_{-i}\}$. Consider any minimally bidder-informative protocol $P$ that implements $\phi$. By Lemma˜2, $(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$. But $\phi(\theta_{i},\bm{\theta}_{-i})=\phi(\theta_{i}^{\prime},\bm{\theta}_{-i})$ so $P$ produces a contextual privacy violation for bidder $i$. ∎
Our results indicate that designers need not search for a protocol that preserves both desired privacy properties; such a protocol does not exist. We may also now use Theorem˜1 to show that several protocols which preserve privacy from the auctioneer may not be preferred by a privacy-conscious designer. In particular, in a setting that violates the indistinguishable corners condition, if a protocol $P$ prevents the auctioneer from learning any unnecessary information about bidder $i$, Theorem˜1 implies that we can produce a protocol that is strictly less bidder-informative than $P$. This complicates the argument for using such protocols from purely a privacy perspective without additionally considering privacy from whom.
4.2 Stitching
Theorem 1 shows that the two endpoints of the Privacy Pareto Frontier are distinct. What happens between the two endpoints—are there protocols that interpolate between preserving privacy from bidders and preserving privacy from the auctioneer? Generally, yes: The designer can take the problem of implementing $\phi$, break it down into simpler implementation problems, choose protocols on the frontier for each sub-problem, and “stitch” the chosen protocols together to form a protocol that implements the original $\phi$. [^13] Since we can always prune a protocol that arises from stitching, every protocol that arises from stitching also has a privacy-equivalent and incentive-compatible pruned version. This gives the auctioneer flexibility along the two dimensions of privacy: for example, the auctioneer can choose a maximally contextually private protocol for one subproblem and a minimally bidder informative protocol for the other subproblem. A sufficient condition for stitching to work is for the protocols used in each sub-problem to be “separable” in the following sense.
Lemma 4 (Stitching).
[…]
where $\hat{x}$ is some new outcome not in $X$. Suppose $P^{2}$ is on the frontier for $\phi$ restricted to $\hat{\bm{\Theta}}$ and $P^{1}$ is on the frontier for $\phi^{\prime}$ and produces no contextual privacy violations or bidder privacy violations within $\hat{\bm{\Theta}}$:
  1. 1.
    For any $\bm{\theta},\bm{\theta}^{\prime}\in\hat{\bm{\Theta}}$,
    $\bm{\theta}\sim_{(P^{1},A)}\bm{\theta}^{\prime};$
  2. 2.
    For any $i$ and $\bm{\theta},\bm{\theta}^{\prime}\in\hat{\bm{\Theta}}$ with $\theta_{i}=\theta_{i}^{\prime}$,
    $(\theta_{i},\bm{\theta}_{-i})\sim_{(P^{1},i)}(\theta_{i},\bm{\theta}_{-i}^{\prime}).$
Then, the protocol $P$ of running $P^{1}$ and then $P^{2}$ if $P^{1}$ outputs $\hat{x}$ is on the frontier for $\phi$.
A sufficient condition for the two conditions above is for $P^{1}=(G^{1},S^{1})$ to have $S_{i}^{1}(\theta_{i})=S_{i}^{1}(\theta_{i}^{\prime})$ for all $\theta_{i},\theta_{i}^{\prime}\in\hat{\Theta}_{i}.$ This stronger condition is often simpler to check and will be used when applying the stitching lemma to first-price auctions in the next section.
A stitched protocol inherits the privacy properties of its sub-protocols. Suppose $P$ is obtained from stitching $P^{1},P^{2}$ and $\hat{P}$ is obtained from stitching $\hat{P}^{1},\hat{P}^{2}$ where both stitchings were done with respect to the same subproblems induced by the same $\hat{\Theta}$. If $P^{1}$ is more bidder-private (auctioneer-private) than $\hat{P}^{1}$ and $P^{2}$ is more bidder-private (auctioneer-private) than $\hat{P}^{2}$, then $P$ is more bidder-private (auctioneer-private) than $\hat{P}$. Otherwise, $P$ cannot be ordered against $\hat{P}$.
5 First-Price Auctions
Consider an auctioneer conducting a first-price auction. In a sealed-bid first-price auction, the auctioneer learns every bidder’s bid and hence learns every bidder’s value (by inverting their bid functions). In contrast, a descending auction implements the first-price auction choice rule with no contextual privacy violations since the auctioneer learns nothing about losing bidders’ values, except that they are not the highest value. However, the descending auction does not privacy dominate the sealed-bid first-price auction. In a descending auction, all bidders learn the value of the winning bidder: Each bidder observes the last price the auctioneer offers, and hence can perfectly back out the winning bidder’s bid. In contrast, in a sealed-bid auction, each losing bidder learns only that their value is not the highest. By conditioning questions on the reports of previous bidders, the auctioneer reveals information about the winner’s value in a descending auction.
In this section, we first show that a first-price auction violates the indistinguishable corners condition for every bidder. This immediately implies that the descending auction is not minimally bidder-informative. We then show that both the descending auction and sealed-bid first-price auction lie on the privacy Pareto frontier. Privacy arguments alone are therefore insufficient for a designer to select between the descending auction and sealed-bid auction protocols.
Given our focus on deterministic protocols, we consider a first-price auction rule with lexicographic tie-breaking. The type spaces are $\Theta_{i}\subset\mathbb{R}_{+}$ for $i\in\{1,\dots,N\}$.
Definition 13.
Given strictly increasing bidding functions $b_{i}:\Theta_{i}\to\mathbb{R}_{+}$ for $i\in\{1,\dots N\}$, the first-price auction choice rule $\phi^{FPA}$ is defined by
$\phi^{FPA}(\bm{\theta})=(\phi_{1}^{FPA},\dots,\phi_{n}^{FPA})(\bm{\theta)}=((q_{1},t_{1}),\dots,(q_{n},t_{n}))(\bm{\theta})$
with
$(q_{j},t_{j})(\bm{\theta})=\begin{cases}(1,b_{i}(\theta_{i}))&\text{ if }i=\min\{k:b_{k}(\theta_{k})=\max_{j\in N}b_{j}(\theta_{j})\}\\ 0&\text{ otherwise}\end{cases}$
Defining the first-price auction allocation rule with respect to given bid functions disentangles incentive compatibility and privacy concerns. Whether $\phi^{FPA}$ is incentive-compatible depends on the choice of bid functions $\{b_{i}\}$, while for any $\{b_{i}\}$, the same privacy concerns persist. We show that $\phi^{FPA}$ violates the indistinguishable corners condition for every bidder $i$, implying that our impossibility theorem applies for each bidder.
Lemma 5.
Suppose $\Theta_{i}=\Theta$ for all $i$ and $|\Theta|>3$. [^14] This condition ensures that there is sufficient overlap between bidders’ type spaces. Then $\phi^{FPA}$ violates the indistinguishable corners condition for every bidder.
To see why $\phi^{FPA}$ violates the indistinguishable corners, consider the pair of values $\theta_{i},\theta_{i}^{\prime}\in\Theta_{i}$ with $\theta_{i}<\theta_{i}^{\prime}$. If the highest bid among other bidders is some $b_{j}(\theta_{j})>b_{i}(\theta_{i}^{\prime})$, then bidder $i$ is not pivotal to the outcome at these types: she will lose the auction regardless. However, if the highest bid among other bidders is some $b_{j}(\theta_{j})\in(b_{i}(\theta_{i}),b_{i}(\theta_{i}^{\prime}))$, then bidder $i$ is pivotal to the outcome at these types: whether she has value $\theta_{i}$ or $\theta_{i}^{\prime}$ determines the winner of the auction. Finally, if bidder $i$ loses the auction, all she observes is that she loses the auction: this is why the indistinguishable corners condition is violated, despite the corners condition not being violated. [^15] The existence of a protocol that produces no contextual privacy violations (the descending auction protocol) is consistent with the absence of corners condition violations in $\phi^{FPA}$. However, because $\phi^{FPA}$ violates the indistinguishable corners condition, we may apply Theorem 1 Given that our impossibility theorem applies, what protocols are on the privacy frontier? Two natural candidates are the descending protocol and the sealed-bid protocol, which are commonly used in practice.
Definition 14.
The descending auction protocol for $\phi^{FPA}$ successively queries bidders as follows. Let $\tilde{\bm{\Theta}}=\bigtimes_{i=1}^{n}\tilde{\Theta}_{i}$ denote the potential type profiles remaining at each stage. At each stage, there exists bidder $i$ and a minimum threshold $\hat{\theta}_{i}\in\tilde{\Theta}_{i}$ for which $\bm{\theta}_{-i}\in\tilde{\bm{\Theta}}_{-i}$ and $\theta_{i}\geq\hat{\theta}_{i}$ implies that $i$ wins the good. Ask bidder $i$ if their value is at least $\hat{\theta}_{i}$. If so, award the object to bidder $i$ at price $b_{i}(\hat{\theta}_{i})$; otherwise, update $\tilde{\Theta}$ and continue.
The descending protocol only reveals the winner’s value and that losers’ values were below the winner’s value to the auctioneer. This is exactly the information necessary to compute the outcome. Hence, as haupt2025contextually note, it produces no contextual privacy violations.
Lemma 6.
(haupt2025contextually.) The descending auction protocol for a first-price auction produces no contextual privacy violations for any bidder.
By our impossibility theorem (Theorem˜1) and Lemma˜5, this then implies that the descending auction protocol is not minimally bidder-informative: When a losing bidder $i$ observes that the clock has stopped at some $b_{i}(\theta)$, she can immediately infer that the winner has value $\theta$.
Corollary 1.
The descending auction protocol for a first-price auction is not minimally bidder-informative. Any protocol on the privacy frontier implementing $\phi_{FPA}$ either reveals strictly more information to bidders than the sealed-bid FPA protocol, or reveals strictly more information to the auctioneer than the descending FPA protocol.
The second sentence of the corollary holds for the same reason: no auction format can simultaneously be minimally bidder-informative and produce no contextual privacy violations. Despite this, the descending protocol is on the frontier, being the only fully contextually private protocol. Sealed-bid auctions are also on the frontier (at the other end), being the only minimally bidder-informative protocol.
Proposition 3.
The descending protocol lies on the privacy Pareto frontier for $\phi_{FPA}$.
Proposition 4.
The sealed-bid protocol lies on the privacy Pareto frontier for $\phi_{FPA}$.
Although the sealed-bid protocol is minimally bidder-informative, this does not immediately imply that it lies on the privacy Pareto frontier. [^16] For example, in the second-price auction setting, we will show that the sealed-bid second-price auction protocol does not lie on the privacy Pareto frontier. Instead, we must show that all information learned by the auctioneer in the sealed-bid protocol cannot be omitted without giving bidders additional information. We use Lemma˜3 to show that the unique way to preserve minimal bidder-informativeness requires the auctioneer fully distinguishing bidders’ types.
5.1 Stitching and Self-Similarity in the First-Price Auction
Suppose there were two bidders, $A$ and $B$, each with values in $\{1,2,3,4,5\}$. The following figure plots $\phi^{FPA}$ under truthful bidding (for the purpose of illustration, fully abstract away from incentive concerns) where $i,p$ denotes bidder $i$ winning at a price of $p$:
12345Bidder AA12345Bidder BBA,1A,2A,3A,4A,5B,2B,3B,4B,5
Figure 3: Choice function for $\phi^{FPA}$.
Let $\hat{X}$ be the set of outcomes corresponding to the price being less than or equal to $3$, so $(\phi^{FPA})^{-1}(\hat{X})=\hat{\bm{\Theta}}=\{1,2,3\}\times\{1,2,3\}$. [^17] In the general, multi-bidder case, we can take $\hat{X}$ to still be the set of outcomes corresponding to the price being less than or equal to some threshold $p$, leading $\hat{\Theta}=\prod_{i}\{\theta_{i}:b_{i}(\theta_{i})<p\}$. All the structure illustrated in this example generalizes. Construct $\phi^{\prime}$ by setting $\phi^{\prime}(\bm{\theta})=\hat{x}$ for all $\bm{\theta}\in\hat{\bm{\Theta}}$. Splitting up the original implementation problem into implementing $\phi^{\prime}$ on $\bm{\Theta}$ and $\phi$ on $\hat{\bm{\Theta}}$ yields the following:
[…]
Figure 4: Splitting $\phi^{FPA}$
When considering $\phi^{\prime}$, suppose values of $1,2$, and $3$ were homogenized into one type, called $\leq 3$.” Then, $\phi^{\prime}$ has the exact same structure as a first-price auction on this simpler type space; furthermore, if $\leq 3$ is treated as one type, then no protocol distinguishes between types $1,2,$ and $3$. As such, the assumptions of the stitching lemma are satisfied, and the designer can choose any protocol on the frontiers of $\phi^{\prime}$ with the simpler type space and $\phi$ restricted to $\hat{\bm{\Theta}}$.
Suppose the designer cares about bidder privacy more than auctioneer privacy at small valuations (perhaps because post-auction competition is more intense at these valuations) but cares more about auctioneer privacy than bidder privacy at high valuations (perhaps because bidder with especially high valuations in this auction are more likely to win future auctions, which the auctioneer can use to their advantage). The designer may then choose to use a descending protocol on $\phi^{\prime}$ with the simpler type space and a sealed-bid protocol on $\phi$ restricted to $\hat{\bm{\Theta}}$. This produces the following “descending then sealed-bid” protocol:
12345Bidder AA12345Bidder BBA,4A,5B,4B,5A,1A,2A,2B,2A,3A,3A,3B,3B,3
Figure 5: Partition Induced by the “descending then sealed-bid” protocol.
By choosing $\hat{\Theta}$, the designer can change when the auction switches from the descending stage to the sealed-bid stage. If one descending then sealed-bid protocol descends for longer than another, then it preserves more privacy from the auctioneer at the cost of bidders learning more. At the two extremes of this tradeoff are the descending protocol and sealed-bid protocol themselves. The privacy frontier for first-price auctions looks like the following (in a simplified two-dimensional representation):
Auctioneer PrivacyBidder Privacydescendingsealed-biddescending then sealed-bid
Figure 6: Privacy Pareto Frontier for First-Price Auctions
6 Second-Price Auctions
Consider an auctioneer implementing a second-price auction. Despite having corners condition violations, the second-price auction has additional violations of the indistinguishable corners condition so our impossibility theorem applies in this setting. The ascending auction and the ascending-join auction lie on the privacy Pareto frontier whereas the sealed-bid second-price auction is privacy dominated by the ascending auction.
Definition 15.
The second-price auction $\phi^{SPA}$ is defined by
$\phi^{SPA}(\bm{\theta})=(\phi_{1}^{SPA},\dots,\phi_{n}^{SPA})(\bm{\theta)}=((q_{1},t_{1}),\dots,(q_{n},t_{n}))(\bm{\theta})$
with
$(q_{j},t_{j})(\bm{\theta})=\begin{cases}(1,\max_{j\neq i}\theta_{j})&\text{ if }i=\min\{k:\theta_{k}=\max_{j\in N}\theta_{j}\}\\ 0&\text{ otherwise}\end{cases}$
Unlike the first-price auction, the second-price auction violates the standard corners condition. [^18] Our impossibility theorem still has implications beyond haupt2025contextually, even in settings with only violations of the standard corners condition. In such a setting, the designer can construct a protocol (in this case, the ascending-join protocol) which produces no contextual privacy violations for one protected bidder; our impossibility theorem shows that such a protocol cannot be minimally bidder-informative. $\phi^{SPA}$ also has indistinguishable corners conditions violations which are not corners conditions violations. Intuitively, whenever a bidder is relevant in determining the price but not who receives the good, they become pivotal between outcomes indistinguishable to them.
Lemma 7.
Suppose $\Theta_{i}=\Theta$ for all $i$ and $|\Theta|>5$. Then for each bidder, there exists type profiles that violate the indistinguishable corners condition for $\phi^{SPA}$ but do not violate the corners condition.
Two protocols that the designer could consider using are the sealed-bid protocol and the ascending protocol. The ascending protocol increases the price offered to each bidder until all but one of the bidders drops out.
Definition 16.
The ascending auction protocol has information sets $\mathcal{I}_{i}=\{I_{\theta_{i}}\}_{\theta_{i}\in\Theta_{i}}$ for bidder $i$ with $A(I_{\theta_{i}})=\{\text{Stay},\text{Drop}\}$. Let $p(i)$ denote the highest information set that bidder $i$ has reached in bidder auction. At each stage, the auctioneer queries the bidder with the lowest $p(i)$ who is still in the auction. If the bidder has value $\theta_{i}\leq p(i)$, they leave the auction. Otherwise, bidder $i$’s $p(i)$ is incremented. The last remaining bidder wins the good at the highest dropout price among other bidders.
The ascending-join protocol from haupt2025contextually instead reveals strictly less information to the auctioneer than a sealed-bid auction (for any $N$) or ascending auction (for $N>2$). The ascending-join protocol begins by increasing the clocks for two bidders until one bidder drops out, then successively asks new bidders whether their value is above the current second-highest value. (See the appendix for a formal definition of the ascending-join protocol.) This protocol is one way to minimize what the auctioneer learns about losers’ values. However, it reveals more information to bidders than the sealed-bid auction: if a bidder’s first query is “Is your value above pp?” the bidder immediately learns that the second-highest value among all previous bidders is $p$. The ascending-join protocol therefore fails to be minimally bidder-informative; this also follows immediately from our impossibility theorem by noting that there is no contextual privacy violation for the final bidder who joins the auction.
Proposition 5.
For $N>2$, the ascending-join auction is not minimally bidder-informative. In particular, the ascending-join auction does not privacy dominate the sealed-bid second-price auction.
This result shows that a privacy-conscious designer need not prefer the ascending-join auction format. If privacy from bidders is more important than privacy from the auctioneer, she may instead prefer the sealed-bid second-price auction protocol. However, the sealed-bid protocol is in turn privacy dominated by the ascending auction, and thus does not lie on the privacy Pareto frontier. That is, a designer who solely values privacy for all parties involved should always weakly prefer using the ascending auction instead of the sealed-bid second-price auction.
Proposition 6.
The sealed-bid second-price-auction is privacy-dominated by the ascending auction.
The ascending auction may initially appear to reveal more to the winning bidder than the sealed-bid auction does. However, the only additional information that the ascending auction reveals to the winner is the second-highest value. Since the winner learns this information via their payment regardless of the protocol used, the ascending auction reveals the same information to bidders as the sealed-bid auction does. We then show that the ascending protocol and the ascending-join protocol lie on the privacy Pareto frontier.
Proposition 7.
The ascending auction is on the privacy Pareto frontier.
As described above, the ascending auction is also minimally bidder-informative. However, as Proposition˜6 illustrates, a minimally bidder-informative protocol need not lie on the privacy frontier. The ascending auction reveals all losing bidders’ values to the auctioneer, and reveals no unnecessary information about the winning bidder to the auctioneer. The difficulty in the proof is therefore to show that in any minimally bidder-informative protocol which implements $\phi^{SPA}$, the auctioneer must learn the value of all losing bidders.
In our proof, we use Lemma˜2 to show that the auctioneer is forced to distinguish all losers’ values. Concretely, consider a type profile $\bm{\theta}_{-i}$ and $\theta_{i},\theta_{i}^{\prime}\in\Theta_{-i}$ such that bidder $i$ loses when she has value $\theta_{i}$ or $\theta_{i}^{\prime}$. We construct a type profile $\bm{\theta}_{-i}^{\prime}$ such that bidder $i$ is pivotal at type profile $\bm{\theta}_{-i}^{\prime}$, that is, that these type profiles violate the indistinguishable corners condition. Then by Lemma˜2, to maintain minimal bidder-informativeness, the auctioneer must distinguish $\theta_{i}$ and $\theta_{i}^{\prime}$ at type profile $\bm{\theta}$. Finally:
Proposition 8.
The ascending-join protocol is on the privacy Pareto frontier.
This statement is straightforward to show, though non-trivial because we must show that the designer cannot reorder the queries in the ascending-join protocol to obtain the same partition $\mathcal{P}(\sim_{(P,A)})$ while reducing the information revealed to bidders. In sum, the privacy frontier for $\phi^{SPA}$ can be represented by Figure˜7.
Auctioneer PrivacyBidder PrivacyMinimally Bidder-Informativeascending-joinascendingsealed-bid
Figure 7: Privacy Pareto Frontier for Second-Price Auctions
7 Conclusion
We study the problem a designer faces when constructing a deterministic bilateral communication protocol among an auctioneer and multiple bidders to implement a given choice rule. Information leakages to the auctioneer or bidders may have unintended and negative consequences in a variety of economic settings. However, the designer faces an inherent tradeoff between preventing the auctioneer from unnecessarily learning bidders’ private information and preventing bidders from unnecessarily learning each others’ private information. Protocols can be partially ordered based on how much information they reveal to the auctioneer or to other bidders; protocols for which no alternative protocols reveal less to both the auctioneer and to other bidders are said to lie on the privacy Pareto frontier. For first-price auctions, the sealed-bid protocol (which reveals minimal information to bidders) and the descending protocol (which reveals minimal information to the auctioneer) lie on the frontier. For second-price auctions, the ascending protocol (which reveals minimal information to bidders) and the ascending-join protocol (which reveals minimal information to the auctioneer) lie on the frontier. However, our privacy considerations have bite: The sealed-bid protocol is not on the frontier for second-price auctions. Beyond the extreme points of the frontier, the designer can interpolate between these (and other) formats to trade off between preserving privacy from the auctioneer or from other bidders at different regions in the type space. These tradeoffs persist beyond auction settings: All our characterizations hold for arbitrary choice functions. Our results highlight that the criterion of simply preserving privacy is necessarily incomplete without first answering the question of privacy from whom.
Several avenues for further research remain. First is establishing a sharp characterization of protocols on the frontier. In general, the frontier is ill-behaved for several reasons. For some choice functions, there exist protocols on the frontier that violate both bidder and auctioneer privacy at certain type profiles. Similarly, protocols which induce a coarser partition of the type space unambiguously reveal less information to the auctioneer, but the effects of moving to a coarser partition on bidder privacy are ambiguous. Second is endogenizing privacy and studying optimal design under such concerns. If informational advantages are payoff-relevant to participants in the protocol, then there may be incentives to deviate purely for the purpose of information acquisition. In a descending auction, for example, there is an additional informational benefit to bid-shading: If the winner reduces the price at which they claim the object, that lowers the upper bound on every losing bidder’s value (giving them more information) while reducing the lower bound other bidders have about their value (giving competitors less information). As such, privacy concerns may put downward pressure on auction prices. Incorporating such concerns into equilibrium analysis may further understanding of bidder behavior. Finally, future work could connect privacy to other concerns that the extensive-form mechanism design literature has studied. If bidders learn more along the path of play of a protocol, the protocol is less private from bidders. However, this could result in (1) bidders needing less contingent reasoning, resulting in a protocol that is strategically simpler and (2) bidders having more information to detect potential auctioneer deviations. This interplay between privacy and other economic forces is important yet underdeveloped.
Appendix A: Omitted Proofs
PROOF OF LEMMA 1
Proof.
One possible protocol is the sealed-bid protocol, where the auctioneer asks each bidder to report their type in a sealed bid and then computes the resulting outcome from $\phi$. bidders are always queried in exactly the same way, regardless of other bidders’ types, so they have the same experience regardless of others’ types, and hence learn nothing besides what they learn about the outcome. That is, if $\phi(\theta_{i},\bm{\theta}_{-i})$ and $\phi(\theta_{i},\bm{\theta}_{-i}^{\prime})$ are in the same element of the partition $\Omega_{i}$, then $o_{i}(\theta_{i},\bm{\theta}_{-i})=o_{i}(\theta_{i},\bm{\theta}_{-i}^{\prime})$. Thus, if $\bm{\theta},\bm{\theta}^{\prime}$ satisfy $\theta_{i}=\theta_{i}^{\prime}$ then $o_{i}(\bm{\theta})=o_{i}(\bm{\theta}^{\prime})$ and $\bm{\theta}\sim_{(P,i)}\bm{\theta}^{\prime}$. Otherwise, if $\bm{\theta},\bm{\theta}^{\prime}$ have $\theta_{i}\neq\theta_{i}^{\prime}$ then $i$ can already distinguish between the two type vectors regardless of $P,\Omega_{i},\phi$. This leads to the desired partition in the definition of minimally bidder-informative.
Next, no protocol $P^{\prime}$ can lead to a coarser information partition for any bidder $j$. Towards a contradiction, suppose there some $P^{\prime},j$ exists. Then, there must exist $\bm{\theta},\bm{\theta^{\prime}}$ such that $\bm{\theta}\not\sim_{(G,j)}\bm{\theta^{\prime}}$ but $\bm{\theta}\sim_{(P^{\prime},j)}\bm{\theta^{\prime}}$. The only way for this to happen is if
$\bm{\theta}\sim_{(P^{\prime},j)}\bm{\theta}^{\prime}\text{ but }\theta_{j}\neq\theta^{\prime}_{j}\text{ or }\phi(\bm{\theta})\not\sim_{\Omega_{j}}\phi(\bm{\theta}^{\prime}).$
If $\theta_{j}\neq\theta_{j}^{\prime}$ then $j$ always knows their own type and thus can tell $\bm{\theta},\bm{\theta}^{\prime}$ apart. If $\phi(\bm{\theta})\not\sim_{\Omega_{i}}\phi(\bm{\theta}^{\prime})$, then by the definition of $\Omega_{j}$, $j$ can tell $\bm{\theta},\bm{\theta}^{\prime}$ apart. ∎
PROOF OF PROPOSITION 1
Proof.
First, auctioneer privacy. By definition of $\sim_{(P,A)}$, we have $\bm{\theta}\sim_{(P,A)}\bm{\theta}^{\prime}$ is equivalent to $o_{A}(P,\bm{\theta})=o_{A}(P,\bm{\theta}^{\prime})$. Pruning leaves strategies unchanged and does not remove the on-path history $o_{A}(P,\bm{\theta})$, so $o_{A}(P,\bm{\theta})=o_{A}(P,\bm{\theta}^{\prime})$ holds if and only if $o_{A}(\tilde{P},\bm{\theta})=o_{A}(\tilde{P},\bm{\theta}^{\prime})$. However, $o_{A}(\tilde{P},\bm{\theta})=o_{A}(\tilde{P},\bm{\theta}^{\prime})$ is in turn equivalent to $\bm{\theta}\sim_{(\tilde{P},A)}\bm{\theta}^{\prime}$. Thus $\bm{\theta}\sim_{(P,A)}\bm{\theta}^{\prime}$ if and only if $\bm{\theta}\sim_{(\tilde{P},A)}\bm{\theta}^{\prime}$.
Next, bidder privacy. Suppose $\bm{\theta}\sim_{(P,i)}\bm{\theta}^{\prime}$. This happens if and only if:
  1. 1.
    $\theta_{i}=\theta^{\prime}_{i}$;
  2. 2.
    $\phi(\bm{\theta})\sim_{\Omega_{i}}\phi(\bm{\theta^{\prime}})$;
  3. 3.
    bidder $i$ observes the same sequence of information sets and takes the same actions along the path of play of $\bm{\theta}$ and $\bm{\theta}^{\prime}$ under protocol $P$.
Since conditions one and two do not depend on the game $G$ or $\tilde{G}$ being played, it suffice to focus on condition three. Let $\{h_{k}(\bm{\theta})\}_{k}$ be the sequence of histories at type profile $\bm{\theta}$ and let $\{h_{k}(\bm{\theta})\}_{k:i(h_{k})=i}$ be the subsequence of histories for which bidder $i$ is called to play. Then, condition three can be stated as
$\{I(h_{k}(\bm{\theta}))\}_{k:i(h_{k})=i}=\{I(h_{k}(\bm{\theta}^{\prime}))\}_{k:i(h_{k})=i}.$ (1)
For any on-path histories $h,h^{\prime}\in\tilde{H}$, we have $I(h)=I(h^{\prime})\Leftrightarrow\tilde{I}(h)=\tilde{I}(h^{\prime})$, as $\tilde{I}$ is simply a restriction of $I$ to on-path histories. Thus, 1 is equivalent to
$\{\tilde{I}(h_{k}(\bm{\theta}))\}_{k:i(h_{k})=i}=\{\tilde{I}(h_{k}(\bm{\theta}^{\prime}))\}_{k:i(h_{k})=i}$ (2)
which is then equivalent to bidder $i$ observing the same sequence of information sets along the path of play of $\bm{\theta}$ and $\bm{\theta}^{\prime}$ under $\tilde{P}$. Since the strategies $S$ are unchanged under pruning, the actions bidder $i$ plays are similarly unchanged. Thus, if bidder $i$ observes the same sequence of information sets and takes the same actions along the path of play of $\bm{\theta}$ and $\bm{\theta}^{\prime}$ under protocol $P$, they will also observe the same sequence of information sets and takes the same actions along the path of play of $\bm{\theta}$ and $\bm{\theta}^{\prime}$ under protocol $\tilde{P}$. ∎
PROOF OF PROPOSITION 2
Proof.
Suppose $\phi$ is dominant-strategy incentive compatible. Towards a contradiction, suppose $P$ is not dominant-strategy incentive compatible. Then, there exists $i,\theta_{i},\bm{\theta_{-i}},\sigma_{i}^{\prime}$ such that
$u_{i}(G(S_{i}(\theta_{i}),\bm{S}_{-i}(\bm{\theta}_{-i}),\theta_{i})<u_{i}(G(\sigma_{i}^{\prime},\bm{S}_{-i}(\bm{\theta}_{-i}),\theta_{i}).$ (3)
As $G$ is pruned, the sequence of histories reached under $(\sigma_{i}^{\prime},\bm{S}_{-i}(\bm{\theta}_{-i}))$ must have been reachable via $\bm{S}(\bm{\theta}^{\prime})$ for some $\bm{\theta}^{\prime}$. Without loss, we can take $\bm{\theta}^{\prime}_{-i}=\bm{\theta}_{-i}$; furthermore, $\sigma_{i}^{\prime}=S_{i}(\theta_{i}^{\prime})$. As $P$ implements, $\phi$, we have that
$G(S_{i}(\theta_{i}),\bm{S}_{-i}(\bm{\theta}_{-i}))=\phi(\theta_{i},\bm{\theta}_{-i})\text{ and }G(\sigma_{i}^{\prime},\bm{S}_{-i}(\bm{\theta}_{-i})=G(S_{i}(\theta_{i}^{\prime}),\bm{S}_{-i}(\bm{\theta}_{-i})=\phi(\theta_{i}^{\prime},\bm{\theta}_{-i}).$
Plugging this back into (3) gives
$u_{i}(\phi(\theta_{i},\bm{\theta}_{-i}),\theta_{i})<u_{i}(\phi(\theta_{i}^{\prime},\bm{\theta}_{-i}),\theta_{i})$
which contradicts $\phi$ being dominant-strategy incentive compatible. ∎
PROOF OF LEMMA 4
Proof.
Clearly, $P$ implements $\phi$. Towards a contradiction, suppose there exists some $\hat{P}$ that implements $\phi$ and dominates $P$. Let $\hat{P}^{2}$ be equal to $\hat{P}$ restricted to type profiles in $\hat{\bm{\Theta}}$ so $\hat{P}^{2}$ implements $\phi|_{\hat{\bm{\Theta}}}$. Similarly, let $\hat{P}^{1}$ be equal to $\hat{P}$ truncated to end and output $x^{\prime}$ whenever the auctioneer knows that $\bm{\theta}\in\hat{\bm{\Theta}}$, implementing $\phi^{\prime}$. The relationship between $\hat{P},\hat{P}^{1},\hat{P}^{2}$ is the same between the relationship between $P,P^{1},P^{2}$.
Note that the following are equivalent:
  1. 1.
    $(\theta_{i},\bm{\theta}_{-i})\not\sim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$;
  2. 2.
    $\theta_{i}$ and $\theta_{i}^{\prime}$ take different actions along the path of play of $P$ at $\bm{\theta}_{-i}$;
  3. 3.
    $\theta_{i}$ and $\theta_{i}^{\prime}$ take different actions along the path of play of $P$ at $\bm{\theta}_{-i}$ either before or after the auctioneer can conclude that $\bm{\theta}\in\hat{\bm{\Theta}}$;
  4. 4.
    $(\theta_{i},\bm{\theta}_{-i})\not\sim_{(P^{1},A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$ or $(\theta_{i},\bm{\theta}_{-i})\not\sim_{(P^{2},A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$.
Similarly, the following are also equivalent:
  1. 1.
    $(\theta_{i},\bm{\theta}_{-i})\not\sim_{(P,i)}(\theta_{i},\bm{\theta}_{-i}^{\prime})$;
  2. 2.
    $\bm{\theta}_{-i}$ and $\bm{\theta}_{-i}^{\prime}$ induce different experiences along the path of play of $P$ at $\theta_{i}$;
  3. 3.
    $\bm{\theta}_{-i}$ and $\bm{\theta}_{-i}^{\prime}$ induce different experiences along the path of play of $P$ at $\theta_{i}$ either before or after the auctioneer can conclude that $\bm{\theta}\in\hat{\bm{\Theta}}$;
  4. 4.
    $(\theta_{i},\bm{\theta}_{-i})\not\sim_{(P^{1},i)}(\theta_{i},\bm{\theta}_{-i}^{\prime})$ or $(\theta_{i},\bm{\theta}_{-i})\not\sim_{(P^{2},i)}(\theta_{i},\bm{\theta}_{-i}^{\prime})$.
As such, letting $A(P)$ denote the set of contextual privacy violations and $B(P)$ denote the set of bidder privacy violations from a protocol $P$, we have
$A(P)=A(P^{1})\cup A(P^{2});B(P)=B(P^{1})\cup B(P^{2})$
and
$A(\hat{P})=A(\hat{P}^{1})\cup A(\hat{P}^{2});B(\hat{P})=B(\hat{P}^{1})\cup B(\hat{P}^{2}).$
Furthermore, all these unions are disjoint unions. As $P^{1}$ is on the frontier, $A(\hat{P}^{1})\setminus A(P^{1})$ and $B(\hat{P}^{1})\setminus B(P^{1})$ are non-empty. Similarly, $P^{2}$ being on the frontier implies that $A(\hat{P}^{2})\setminus A(P^{2})$ and $B(\hat{P}^{2})\setminus B(P^{2})$ are non-empty as well.
It suffice to show that
$A(\hat{P})\setminus A(P)=(A(\hat{P}^{1})\cup A(\hat{P}^{2}))\setminus(A(P^{1})\cup A(P^{2}))$
and
$B(\hat{P})\setminus B(P)=(B(\hat{P}^{1})\cup B(\hat{P}^{2}))\setminus(B(P^{1})\cup B(P^{2}))$
are non-empty as well. Towards a contradiction, suppose they were not non-empty. Then, no $x\in A(\hat{P}^{2})\setminus A(P^{2})$ can be in $A(\hat{P})\setminus A(P)$, which implies that $x$ must be in $A(P^{1})$. However, violations in $\hat{P}^{2}$ can only arise within $\hat{\bm{\Theta}}$ (since $\hat{P}^{2}$ is fundamentally a protocol over $\bm{\hat{}}{\Theta}$) but violations in $P^{1}$ cannot arise purely within $\hat{\bm{\Theta}}$ (by assumption in the Lemma), a contradiction. A similar argument can be repeated for bidder privacy violations. ∎
PROOF OF LEMMA 5
Proof.
Consider any two bidders $i,j$. Choose $\theta_{i}<\theta_{j}^{\prime}<\theta_{i}^{\prime}<\theta_{j}$ and some $\bm{\theta}_{-ij}$ with $\theta_{k}\leq\theta_{i}$ for all $k\neq i,j$. Let $\bm{\theta}_{-i}=(\theta_{j},\bm{\theta}_{-ij})$ and $\bm{\theta}_{-i}^{\prime}=(\theta_{j}^{\prime},\bm{\theta}_{-ij})$.
Then $\phi(\theta_{i},\bm{\theta}_{-i})=\phi(\theta_{i}^{\prime},\bm{\theta}_{-i})$ because both cases lead to bidder $j$ winning at their bid $b_{j}(\theta_{j})$. We also have $\phi(\theta_{i},\bm{\theta}_{-i})\sim_{\Omega_{i}}\phi(\theta_{i},\bm{\theta}_{-i}^{\prime})$ because bidder $i$ loses the auction in both cases. Finally, $\phi(\theta_{i},\bm{\theta}_{-i}^{\prime})\neq\phi(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$, because bidder $i$ loses the auction in the first case and wins the auction in the second case. We conclude that $\phi^{FPA}$ violates the indistinguishable corners condition for bidder $i$. ∎
PROOF OF PROPOSITION 3
Proof.
We show that any protocol which implements $\phi^{FPA}$ and produces no contextual privacy violations must be the descending auction protocol. Given possible type profiles $\tilde{\bm{\Theta}}$, only one bidder $i$ satisfies the condition that there exists a threshold $\hat{\theta}_{i}\in\tilde{\Theta}_{i}$ such that $\theta_{i}\geq\hat{\theta}_{i}$ and $\bm{\theta}_{-i}\in\tilde{\bm{\Theta}}_{-i}$ implies bidder $i$ wins the auction (recall that ties are broken in favor of the lowest-indexed bidder so this bidder is unique). If there were multiple such bidders for which this condition holds, we immediately obtain a contradiction since it could be that both bidders have values above their thresholds but it cannot be that both bidders win the auction.
If a protocol queries any bidder $j\neq i$, then the auctioneer distinguishes some types $\theta_{j},\theta_{j}^{\prime}\in\tilde{\bm{\Theta}}_{j}$, but $\phi(\hat{\theta}_{i},\theta_{j},\bm{\theta}_{-ij})=\phi(\hat{\theta}_{i},\theta_{j}^{\prime},\bm{\theta}_{-ij})$, and therefore produces a contextual privacy violation. Hence any fully contextually private protocol can only ask $i$ to distinguish their type. By choosing the minimal threshold $\hat{\theta}_{i}$, the auctioneers learns the minimal necessary information about bidder $i$. We conclude that any protocol with no contextual privacy violations must be the descending protocol. In particular, no protocol privacy-dominates the descending protocol. ∎
PROOF OF PROPOSITION 4
Proof.
It suffices to show that any minimally bidder-informative protocol $P$ implementing $\phi^{FPA}$ must distinguish all types, that is, $(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$ for all $\theta_{i},\theta_{i}^{\prime}\in\Theta_{i}$ and all $\bm{\theta}_{-i}\in\bm{\Theta}_{-i}$.
If $\phi(\theta_{i},\bm{\theta}_{-i})\neq\phi(\theta_{i}^{\prime},\bm{\theta}_{-i})$, then any protocol implementing $\phi$ satisfies this condition. Hence it suffices to show this condition holds when $\phi(\theta_{i},\bm{\theta}_{-i})=\phi(\theta_{i}^{\prime},\bm{\theta}_{-i})$. To do so, we will construct a violation of the indistinguishable corners condition using this type profile.
Without loss of generality, suppose $\theta_{i}<\theta_{i}^{\prime}$. For any type profiles with $\phi^{FPA}(\theta_{i},\bm{\theta}_{-i})=\phi^{FPA}(\theta_{i}^{\prime},\bm{\theta}_{-i})$, we must have that bidder $i$ loses the auction in both cases (otherwise, their payment would differ in the two cases). We may then construct some type profile $\theta_{-i}^{\prime}\in\Theta_{-i}$ such that
$\theta_{i}\leq\max_{j\neq i}\{\theta_{j}^{\prime}\}\leq\theta_{i}^{\prime}$
with one inequality holding strictly. [^19] The choice of which inequality to make strict depends on the tie-breaking rule. Then we have that $\phi(\theta_{i},\theta_{-i}^{\prime})$ also results in $i$ losing the auction, and hence $\phi(\theta_{i},\theta_{-i})\sim_{\Omega_{i}}\phi(\theta_{i},\theta_{-i}^{\prime})$. However, we have that $\phi(\theta_{i}^{\prime},\theta_{-i}^{\prime})$ results in $i$ winning the auction, and hence $\phi(\theta_{i},\theta_{-i}^{\prime})\neq\phi(\theta_{i}^{\prime},\theta_{-i}^{\prime})$.
Combining these conditions, we conclude that $\phi^{FPA}$ violates the indistinguishable corners condition at $\{\theta_{i},\theta_{i}^{\prime}\}\times\{\bm{\theta}_{-i},\bm{\theta}_{-i}^{\prime}\}$. Hence by Lemma˜2, $G$ must satisfy $(\theta_{i},\theta_{-i})\nsim_{(P,A)}(\theta_{i}^{\prime},\theta_{-i})$ as desired. ∎
PROOF OF LEMMA 7
Proof.
We will construct an indistinguishable corners condition violation for bidder $i$. Consider any three bidders $i,j,k$ and let $\theta_{i}<\theta_{j}^{\prime}<\theta_{i}^{\prime}<\theta_{j}<\theta_{k}$. Choose $\bm{\theta}_{-ijk}$ with $\theta_{\ell}\leq\theta_{i}$ for all $\ell\neq i,j,k$. Thus, bidder $k$ always wins the auction but their payment depends on bidders $i$ and $j$. Take $\bm{\theta}_{-i}=(\theta_{j},\theta_{k},\bm{\theta}_{-ijk})$ and $\bm{\theta}_{-i}=(\theta_{j},\theta_{k},\bm{\theta}_{-ijk})^{\prime}$.
Then, $\phi(\theta_{i},\bm{\theta}_{-i})=\phi(\theta_{i}^{\prime},\bm{\theta}_{-i})$ as both output bidder $\ell$ winning at a price of $\theta_{j}$. Next, $\phi(\theta_{i},\bm{\theta}_{-i})\sim_{\Omega_{i}}\phi(\theta_{i},\bm{\theta}_{-i}^{\prime})$ while $\phi(\theta_{i},\bm{\theta}_{-i})\neq\phi(\theta_{i},\bm{\theta}_{-i}^{\prime})$ as bidder $i$ loses in both cases but the final price is different ($\theta_{j}$ versus $\theta_{j}^{\prime}$). Finally, $\phi(\theta_{i},\bm{\theta}_{-i}^{\prime})\neq\phi(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$ as the final price is different ($\theta_{j}^{\prime}$ versus $\theta_{i}^{\prime}$). As such, $\phi^{SPA}$ has an indistinguishable corners condition violation for bidder $i$ at $\{\theta_{i},\theta_{i}^{\prime}\}\times\{\bm{\theta}_{-i},\bm{\theta}_{-i}^{\prime}\}$ but does not have a corners condition violation for bidder $i$ at those type profiles. ∎
PROOF OF PROPOSITION 6
Proof.
In the ascending auction, the auctioneer learns all losers’ values but does not learn the winner’s value. Hence the ascending auction reveals strictly less to the auctioneer than the sealed-bid second-price auction. In both formats, a losing bidder learns only that she lost the auction. In both protocols, the winning bidder learns the second-highest bidder’s value based on their payment. Hence both protocols reveal the same information to bidders (that is, they are both minimally bidder-informative), and the ascending auction privacy dominates the sealed-bid second-price auction. ∎
PROOF OF PROPOSITION 7
Proof.
As described in the main text, the ascending auction protocol is minimally bidder-informative, reveals the value of all losing bidders to the auctioneer, and reveals nothing about the winner’s value to the auctioneer except that the winner’s value is above the second-highest value. It therefore suffices to show that in any minimally bidder-informative protocol $P$ which implements $\phi^{SPA}$, the auctioneer must learn the value of all losing bidders. For brevity, in this proof we use $\phi$ to refer to $\phi^{SPA}$.
Formally, it suffices to show that for any distinct $\theta_{i},\theta_{i}^{\prime}\in\Theta_{i}$ and any $\bm{\theta}_{-i}$ satisfying $\theta_{i},\theta_{i}^{\prime}\leq\max_{j\neq i}\theta_{j}$, we have $(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$. Without loss of generality, we may consider $\theta_{i}<\theta_{i}^{\prime}$. Also note that if $\phi(\theta_{i},\bm{\theta}_{-i})\neq\phi(\theta_{i}^{\prime},\bm{\theta}_{-i})$, then $(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$ because $G$ implements $\phi$. So it remains to consider the case when $\phi(\theta_{i},\bm{\theta}_{-i})=\phi(\theta_{i}^{\prime},\bm{\theta}_{-i})$.
We will find a type profile $\bm{\theta}_{-i}^{\prime}$ that yields a violation of the indistinguishable corners condition at $\{\theta_{i},\theta_{i}^{\prime}\}\times\{\bm{\theta}_{-i},\bm{\theta}_{-i}^{\prime}\}$. By Lemma 2 and the fact that $G$ is minimally bidder-informative, the auctioneer thus must distinguish between $(\theta_{i},\bm{\theta}_{-i})$ and $(\theta_{i}^{\prime},\bm{\theta}_{-i})$, implying $(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$ as desired. Indeed, consider any $\bm{\theta}_{-i}^{\prime}$ satisfying $\theta_{i}\leq\max_{j\neq i}\theta_{j}^{\prime}\leq\theta_{i}^{\prime}$, with one inequality holding strictly. [^20] The choice of which inequality to hold strictly will depend on the tie-breaking priority of bidder $i$. Then bidder $i$ loses the auction at type profiles $(\theta_{i},\bm{\theta}_{-i})$ and $(\theta_{i},\bm{\theta}_{-i}^{\prime})$, hence $(\theta_{i},\bm{\theta}_{-i})\sim_{\Omega_{i}}(\theta_{i},\bm{\theta}_{-i}^{\prime})$. However, bidder $i$ wins the auction at type profile $(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$, hence $\phi(\theta_{i},\bm{\theta}_{-i}^{\prime})\neq\phi(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$. Combining these three conditions, we conclude that $\phi^{SPA}$ violates the indistinguishable corners condition at $\{\theta_{i},\theta_{i}^{\prime}\}\times\{\bm{\theta}_{-i},\bm{\theta}_{-i}^{\prime}\}$. Hence by Lemma˜2, $(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$. ∎
DEFINITION OF ASCENDING-JOIN PROTOCOL (haupt2025contextually)
Let $\tilde{\bm{\Theta}}\subset\bm{\theta}$ denote the set of potential types remaining. Let $x=\min\{\phi^{SPA}(\theta):\theta\in\tilde{\bm{\Theta}}$ denote the tentative outcome, where $X$ is ordered lexicographically by price then winner identity. Given $\tilde{\bm{\Theta}}$, query agents in order $i=1,2,\dots n$ to ask: “Can you rule out outcome $x$ with your type information?”. Formally, the query separates $\hat{\bm{\Theta}}_{i}=\{\tilde{\theta}_{i}\in\tilde{\Theta}_{i}:\forall\;\tilde{\bm{\theta}}_{-i}\in\tilde{\bm{\Theta}}_{-i},\;\phi(\theta_{i},\tilde{\bm{\theta}}_{-i})\neq x\}$ from $\tilde{\bm{\Theta}}_{i}\setminus\hat{\bm{\Theta}}_{i}$. If agent $i$ reports $\theta_{i}\in\hat{\bm{\Theta}}_{i}$, then $\tilde{\bm{\Theta}}$ is updated and the procedure begins again from bidder 1. If no bidder can rule out outcome $x$, then the protocol terminates and chooses outcome $x$.
PROOF OF PROPOSITION 8
Proof.
Let $P_{AJ}$ denote the ascending-join protocol. By haupt2025contextually, the ascending-join protocol is maximally contextually private, so there does not exist a protocol $P$ which implements $\phi^{SPA}$ and for which $\mathcal{P}(\sim_{(P,A)})$ is strictly coarser than $\mathcal{P}(\sim_{(P_{AJ},A)})$. [^21] Formally, this follows from the statement in the appendix of haupt2025contextually that $P_{AJ}$ is minimally relatively informative among protocols implementing $\phi^{SPA}$. Hence any protocol $P$ which privacy-dominates the ascending-join protocol must satisfy $\mathcal{P}(\sim_{(P_{AJ},A)})=\mathcal{P}(\sim_{(P,A)})$. It therefore suffices to show that if protocol $P$ satisfies $\mathcal{P}(\sim_{(P_{AJ},A)})=\mathcal{P}(\sim_{(P,A)})$, then $P$ must partition bidder types in the same sequence as $P_{AJ}$. This then implies that each bidder observes the same history in $P$ as in $P_{AJ}$. As such, $P$ reveals the same information to bidders as $P_{AJ}$ so $P$ does not privacy dominate $P_{AJ}$. To do this, we show that in order to prevent additional contextual privacy violations that are not produced by ascending-join, each subsequent question must be the next query in the ascending-join protocol.
We say a bidder is active if they are one of the two bidders with minimum index who have not dropped out of the auction. At any point in the ascending-join protocol, we are in one of two cases. Either two bidders are active, or one bidder has just dropped out. We say a bidder $i$ is active at price $p$ if she is one of the two active bidders and the designer has distinguished that $\theta_{i}\geq p$. Any protocol begins with the first two bidders active at $p=0$. We show the auctioneer cannot deviate from the ascending-join protocol without producing an additional contextual privacy violation. For the sake of contradiction, suppose otherwise, so under $P$, the auctioneer does not distinguish any type profiles that she did not distinguish under $P_{AJ}$.
At any stage, the protocol dictates that the auctioneer sends an information set to some agent. We say that the protocol distinguishes values of agent $j$ if she sends a message $I_{j}$ to $j$ for which there exists $\theta_{j},\theta_{j}^{\prime}\in\Theta_{j}$ such that $S(\theta_{j})(I_{j})\neq S(\theta_{j}^{\prime})(I_{j})$. We say that the protocol distinguishes values of agent $j$ strictly above $p$ if she sends a message $I_{j}$ to $j$ for which there exists $\theta_{j},\theta_{j}^{\prime}>p$ with $S(\theta_{j})(I_{j})\neq S(\theta_{i}^{\prime})(I_{j})$. [^22] We define distinguishes values of agent $j$ strictly below $p$ analogously, with $\theta_{j},\theta_{j}^{\prime}<p$. Note that at any point in the protocol, $P$ cannot distinguish values for any bidder who has already dropped out; this produces additional contextual privacy violations not present in $P_{AJ}$. We first prove the result for the case of symmetric $\Theta_{i}=\Theta$ for all $i$, which we may without loss of generality index by $\Theta_{i}=\{1,2,\dots,\bar{\theta}\}$ for all $i$. The case of asymmetric $\Theta_{i}$ follows straightforwardly. [^23] The only changes when $\Theta_{i}$ are asymmetric are as follows. First, to rule out querying a bidder $k>j$ in Case 1 or a bidder $k^{\prime}>k$ in Case 2, we may construct a type profile in which the two active bidders do not have the highest possible value, but which still renders the information from the queried bidder redundant. Second, the incremented price $p+1$ instead becomes $\min\{\theta_{i}\in\Theta_{i}:\theta_{i}>p\}$ for the appropriate bidder $i$.
Case 1: The first case is when two bidders are active. Either both bidders are active at some $p$, or one bidder is active at $p$ and the other bidder is active at $p+1$. Refer to the two active bidders as $i$ and $j$. Consider the next information set in the protocol. First, the protocol cannot distinguish values for any bidder $k>j$, because if $\theta_{i}=\theta_{j}=\bar{\theta}$ (the maximum value), then bidder $k$ is not pivotal, and such a query would distinguish a pair of values that are not distinguished in $P_{AJ}$. Second, for any bidder the auctioneer cannot distinguish values strictly above $p+1$: this violates unconditional winner privacy if all other bidders have values at most $p+1$, unlike the ascending-join protocol. The only remaining possibility is to query one of the two active bidders to distinguish whether their type is weakly greater than $p+1$. This is exactly the subsequent query of the ascending-join protocol.
Case 2: The second case is when some bidder $j$ has just dropped out. Index the remaining active bidder as $i$, where we have $\tilde{\bm{\Theta}}_{i}=\{\theta_{i}:\theta_{i}\geq p\}$ for some price $p$. The next bidder to be queried, under $P_{AJ}$, is $k=\max\{i,j\}+1$. [^24] If there are no remaining bidders to be queried, that is, if $\max\{i,j\}=n$, then the outcome has been determined. Again, any further queries reveals information to the auctioneer that she does not learn in $P_{AJ}$. Then:
  1. 1.
    The auctioneer cannot query a bidder $k^{\prime}>k$ to distinguish her values: if $\theta_{i}=\theta_{k}=\bar{\theta}$, then agent $k^{\prime}$ is never pivotal and this distinguishes types that $P_{AJ}$ does not.
  2. 2.
    The auctioneer cannot query bidder $i$. If all other bidders have value less than $p$, then $i$ must win with a price already determined. In particular, no further information from $i$ is necessary, so any such query distinguishes types in $\bm{\Theta}_{i}$ that $P_{AJ}$ does not.
  3. 3.
    The auctioneer cannot query bidder $k$ to distinguish values strictly below $p+1$. This distinction never affects the outcome, because the auction price is already at least $p$, so such a query distinguishes types that $P_{AJ}$ does not.
  4. 4.
    The auctioneer cannot query bidder $k$ to distinguish values strictly above $p$. If all other bidders have a value of at most $p$, and $\theta_{k}>p$, then $k$ wins the auction regardless. So such a query again distinguishes types that $P_{AJ}$ does not.
The only remaining possibility is for the auctioneer to query bidder $k$ to partition her types into the two sets $\{\theta_{k}:\theta_{k}>p\}$ and $\{\theta_{k}:\theta_{k}\leq p\}$. This is exactly the subsequent query of the ascending-join protocol. ∎