### v2

# Private From Whom? Minimal Information Leakage in Auctions

Abstract
In many auctions, bidders may be reluctant to reveal private information to the auctioneer and other bidders. Among deterministic bilateral communication protocols, reducing what bidders learn requires increasing what the auctioneer learns. A protocol implementing a given social choice rule is on the Privacy Frontier if no alternative protocol reveals less to both bidders and the auctioneer. For first-price auctions, the descending protocol and the sealed-bid protocol are both on the Privacy Frontier. For second-price auctions, the ascending protocol and the ascending-join protocol of Haupt and Hitzig (2025) are both on the Privacy Frontier, but the sealed-bid protocol is not. We provide sufficient conditions for a protocol to be on the Privacy Frontier and devise alternative protocols on the Privacy Frontier for first-price auctions that allow the designer to flexibly trade off between privacy from bidders and the auctioneer.
Keywords: Extensive-Form Mechanism Design, Privacy, Auctions
JEL Codes: D44, D47, D82.
1 Introduction
Bidders participating in auctions may value keeping their information private from both the auctioneer and other bidders. For example, bidders may be concerned that their bids reveal their technology level to rivals or reduce their leverage in post-auction negotiations. [^2] As Rothkopf et al. (1990) describe, “People of our acquaintance with experience in conducting business are reluctant to reveal their true costs or valuations. … It could reveal to potential competitors the extent to which a firm’s technology was superior. Most important, it could reveal to others with whom the firm must subsequently negotiate precisely how much it can yield. … Successful oil lease bidders must deal with drilling contractors, rig owners, and so forth. Successful coal lease bidders must deal with equipment suppliers, railroads, and coal purchasers. Successful construction contract bidders must deal with subcontractors and labor unions.” In the context of spectrum auctions, Milgrom and Segal (2020) write, “The practical significance of UWP [Unconditional Winner Privacy] is that it may alleviate winners’ concerns about misuse of the revealed information, encourage participation by bidders who find it costly to figure out their exact bid, and allow the auctioneer to conceal winners’ politically sensitive windfalls.” Protecting bidders’ private information may encourage bidder participation and prevent the auctioneer from learning sensitive information. [^3] McMillan (1994) describes a New Zealand spectrum auction in which the auction revealed a large gap between the winner’s willingness to pay and her actual payment, resulting in public criticism of the government. A government auctioneer may wish to avoid criticism by reducing the amount she learns about bidders’ values. To what extent can different auction formats protect bidders’ private information from both the auctioneer and other bidders?
Consider a designer choosing between a sealed-bid first-price auction and a descending auction. In a sealed-bid first-price auction, the auctioneer observes every bidder’s exact bid, including losing bids. A descending auction seems to better protect bidders’ privacy: the auction ends when the highest bidder claims the item, so the auctioneer no longer observes exact losing bids. However, a descending auction reveals more information to bidders: each losing bidder now infers that the winning bid is exactly the price at which the auction stops. In contrast, in a sealed-bid first-price auction bidders only learn whether they won or lost.
What drives this tradeoff? Suppose an auctioneer is running a descending auction and a bidder claims the object. The outcome is already determined, so further questions from the auctioneer are unnecessary. But if the auctioneer stops asking questions to some bidder, that bidder infers that a competitor won the auction at the last price they were offered. In general, protecting bidders’ privacy from the auctioneer requires the auctioneer to not ask bidders questions about their type when the outcome is identical regardless of their answer. However, whether a question is necessary depends on what the auctioneer already knows about other bidders’ types: By conditioning which questions to ask on their current information, the auctioneer reveals other bidders’ private information.
This stark tradeoff holds more generally. Consider a designer implementing a given choice rule, which maps from a type space to a set of possible outcomes. [^4] In the single-unit auction setting, a choice rule describes the allocation of the good and agents’ transfers as a function of bidder’ types. The first-price auction choice rule and second-price auction choice rule are two examples. We focus on the single-unit auction setting, but our model and impossibility theorem apply to arbitrary choice rules. To implement the choice rule, the designer chooses a deterministic bilateral communication protocol among an auctioneer and bidders. Under a general condition on the choice rule, implementation necessitates either (1) the auctioneer learning information about every bidder that is not pivotal to the outcome or (2) bidders learning about other bidders’ types beyond what the outcome alone reveals. Consequently, the designer must choose how information leakages are distributed between the auctioneer and bidders. We analyze this understudied additional dimension of privacy—bidders learning about other bidders’ private information.
Given this tradeoff, some privacy violations are inevitable. However, some protocols can still be improved upon from a privacy perspective. Consider now a designer choosing between a sealed-bid second-price auction and an ascending auction. In a sealed-bid second-price auction, the auctioneer learns all bidders’ values; losing bidders learn that the winning bidder’s value is higher than their own; and the winning bidder learns the second-highest value via the price they pay. In an ascending auction, the auctioneer no longer learns the winning bidder’s valuation (the auction ends when the second-to-last bidder drops out), but what bidders learn remains unchanged. [^5] Losing bidders still learn only that the winning bidder’s value is higher than their own; the winning bidder still learns the second-highest value via the price they pay. The ascending auction privacy dominates the sealed-bid second-price auction: It reveals strictly less information to the auctioneer and the same information to bidders. In general, a protocol lies on the Privacy Frontier if it is not privacy dominated by any protocol implementing the same underlying choice rule.
We apply our framework to study the implementation of first-price auctions and second-price auctions. For the first-price auction choice rule, both the sealed-bid first-price auction and the descending auction lie on the Privacy Frontier. We then propose two new protocols on the Privacy Frontier. The descending then sealed-bid protocol trades off privacy from bidders and the auctioneer at different ranges of bidder values. The descending-holdout protocol protects one selected bidder’s privacy from both bidders and the auctioneer.
To implement the second-price auction choice rule, the ascending-join protocol, proposed by Haupt and Hitzig (2025), reveals minimal information to the auctioneer. [^6] Intuitively, the ascending-join protocol orders the bidders and runs an ascending auction with two bidders at a time, until one bidder drops out, then brings the next bidder into the ascending auction. This prevents the auctioneer from learning the value of later bidders, if later bidders have low values and would lose the auction. We elaborate on this protocol in Section 6. However, other protocols reveal less information to bidders. In particular, the ascending-join protocol does not privacy dominate the standard ascending protocol, which reveals minimal information to bidders. We show both the ascending and the ascending-join auctions are on the privacy frontier, but the sealed-bid second-price auction is not.
We then generalize our techniques by providing sufficient conditions to determine whether a given protocol to lie on the Privacy Frontier. These conditions allow a designer to evaluate a protocol only using model primitives and a description of what information the protocol reveals. Our sufficient conditions apply to any type space and choice rule, enabling analysis beyond the structure of auction settings.
Our results give a designer tools to balance revealing information to the auctioneer with revealing information to bidders. The designer’s preferred choice may depend on the specific interaction taking place, which we do not take a stand on. In some settings, such as a one-time auction for a procurement contract, preventing bidders from learning about rival bidders’ technological capabilities may be more important. In other settings, such as repeated auctions for a good (in which the auctioneer may set future reserve prices based on bidders’ values), preserving privacy from the auctioneer may be more important. [^7] Such settings could include repeated auctions for online ads or natural resources.
Our framework of deterministic bilateral communication restricts the usage of alternative tools such as cryptography or an external mediator. [^8] Formally, our designer does not rely on a trusted randomization device as in Izmalkov et al. (2011), and we make no computational hardness assumptions. These alternative tools may be costly to implement and come with their own limitations. Participants must understand or trust that a cryptographic scheme works as intended for the mechanism to be successful. Relying on an external mediator requires bidders to trust that the mediator will not divulge any information. Finally, alternative tools may be inaccessible, become obsolete, or be prohibitively costly for alternative reasons. [^9] In many auctions in practice, the auctioneer and bidders communicate directly without a mediator. These include art and antique auctions (at Sotheby’s or Christie’s), and some auctions for resources such as timber and natural gas. Developments in quantum computing (Cherney (2025)) may undermine the security guarantees of cryptographic schemes. In spectrum auctions, where a graph coloring problem must be solved quickly in every iteration, any increase in time complexity could be prohibitive. With that in mind, our results provide a calculus for deciding when the costs of moving to an alternative scheme are worthwhile. A designer may first wish to understand what privacy protections are feasible in the absence of cryptography before deciding whether to use cryptographic protocols; our work provides this characterization.
The remainder of the paper proceeds as follows. Section˜2 reviews related literature. Section˜3 introduces our model. Section˜4 defines our notion of privacy and states our impossibility theorem, identifying the tradeoff between revealing information to bidders and the auctioneer. Section˜5 and Section˜6 apply our results to first and second-price auctions respectively, situating common auction formats relative to the frontier. Section˜7 provides sufficient conditions for a protocol to lie on the frontier. Section˜8 concludes. All omitted proofs are in Appendix B: Omitted Proofs.
2 Related Literature
Our work is related to the notions of unconditional privacy (which has the strict requirement of zero privacy violations) and contextual privacy (which only analyzes privacy from the auctioneer). Brandt and Sandholm (2008) consider implementing auctions via bilateral communication among bidders without revealing any information, a condition known as unconditional privacy (and considered earlier by Chor and Kushilevitz (1991)). We instead consider settings with an auctioneer and characterize a privacy frontier even when unconditional privacy is unattainable. Milgrom and Segal (2020) and Ausubel (2004) consider the related notion of unconditional winner privacy, which reveals no unnecessary information about the winner’s value; in contrast, we consider the privacy of all bidders. Privacy from the auctioneer is related to relative informativeness (Mackenzie and Zhou (2022), Segal (2007)), but we further analyze what bidders learn from the mechanism. Haupt and Hitzig (2025) characterizes protocols that implement choice rules while revealing minimal information to the auctioneer, which they call contextual privacy. We show that contextual privacy typically conflicts with the desire to reveal minimal information to bidders, leading to a distinct Privacy Frontier.
Other approaches to privacy include Liu (2025), which quantifies the worst-case privacy risk of information structures and studies optimal mechanisms subject to a “budget” of privacy violations. We instead study the maximally privacy-preserving implementation of a given allocation rule. Häfner et al. (2025) models agents who may learn other agents’ actions through channels outside of the mechanism; we study what information must be revealed by the mechanism itself. In contrast with Eilat et al. (2021), who propose a quantitative measure of information that a mechanism reveals to the designer, we study what agents learn about each other during the mechanism. The literature on differential privacy (Dwork (2006), Pai and Roth (2013)) focuses on what a designer learns from a mechanism’s outcome; we broaden the focus to what both the auctioneer and agents learn from playing the mechanism itself.
A broader literature considers protocols that use trusted randomization or computational hardness assumptions. Izmalkov et al. (2011) provide a way to implement any choice rule in the presence of a trusted randomization device that fully preserves privacy. See Alvarez and Nojoumian (2020) for a survey of cryptographic protocols in auctions and see Bogetoft et al. (2009) for applications to an agricultural auction. In contrast, we provide a benchmark for what privacy protections are possible without resorting to cryptographic protocols or trusted randomization.
Other literature discusses the importance of bidders’ privacy in real-world settings, typically nesting a mechanism in a larger model. Rothkopf et al. (1990) consider bidders who are reluctant to reveal their true valuations in Vickrey auctions due to post-auction bargaining. Milgrom and Segal (2020), Ausubel (2004) and McMillan (1994) discuss the importance of protecting winner privacy in spectrum auctions and multi-unit auctions. Dworczak (2020) models an aftermarket which is affected by the information revealed in an auction. We are motivated by similar concerns as this literature, and we provide a framework to compare whose privacy is compromised by a given protocol, and who their information is revealed to.
Finally, our work relates to the broader literature on extensive-form mechanism design. Segal (2013) utilizes extensive-form protocols to study the communication complexity of verifying that a given allocation rule is faithfully executed. Akbarpour and Li (2020), Grigoryan and Möller (2023), Banchio et al. (2025), and Komo et al. (2025) consider an auctioneer or mediator deviating from a specified mechanism. In contrast to this literature, we consider an auctioneer who faithfully executes a specified protocol. Privacy concerns persist even if the auctioneer can commit to the protocol used. Privacy may also be a component of the credibility tradeoff: Whether such deviations are profitable are informed by what the auctioneer learns, while whether such deviations are detectable are disciplined by what bidders learn.
3 Model
Consider a finite set of agents (bidders) $N=\{1,2,...,n\}$, with generic element $i$. Bidder $i$ draws values from type space $\Theta_{i}$, independently of other bidders, and $\bm{\Theta}=\times_{i\in N}\Theta_{i}$. $X$ denotes the set of possible outcomes. Bidder $i$ has utility function $u_{i}:X\times\Theta_{i}\to\mathbb{R}$.
The designer seeks to implement a fixed choice rule $\phi:\bm{\Theta}\to X$. In the auction setting, this rule describes agents’ transfers and allocation as a function of the type profile. Examples include a second-price auction rule or a first-price auction rule. The designer may consider other tradeoffs (such as efficiency, revenue, or desirable incentive properties) when choosing $\phi$. The choice rule may also simply be exogenously imposed for institutional reasons.
For a given choice rule $\phi$, the designer still has considerable flexibility to choose an extensive-form implementation of $\phi$; privacy violations stem from the particular extensive-form implementation chosen. Formally, the designer chooses an extensive-form game $G=(H,i,\{\mathcal{I}_{i}\}_{i\in N},\{A_{i}\}_{i\in N},A,I,g)$ of finite length and perfect recall with notation specified in Table˜1.101010We model the designer as external to the interaction between bidders and the auctioneer. In many settings, the designer that advises organizations on what mechanism to use is not a participant themselves, and simply seeks to identify tradeoffs, bounds on what is inherently feasible, and how to best achieve certain objectives.
Table 1: Notation for Extensive-Form Games
Name Notation Representative Element
Histories $H$ $h$
Player called to play at $h$ $i(h)$
Information sets of player $i$ $\mathcal{I}_{i}$ $I_{i}$
Actions available to $i$ $A_{i}$
Actions available at $I_{i}$ $A(I_{i})$ $a$
Information set at history $h$ $I(h)$
Terminal histories $Z$ $z$
Outcome at terminal node $g(z)$
Interim strategy of player $i$ $\sigma_{i}(\mathcal{I}_{i})$
We limit our attention to deterministic strategy profiles and abstract away from incentives. [^11] Appendix A: Pruning and Incentive Compatibility discusses pruning protocols and establishes that every protocol implementing a dominant strategy incentive compatible choice rule is ex-post incentive compatible. Restricting to deterministic strategies is without loss due to the requirement that a protocol must implement the desired choice rule (with probability one): If two types take the same sequence of actions with positive probability yet the choice rule is still implemented, then these two types never needed to be distinguished. Thus, instead of playing a mixed strategy, the two types could have been pooled, leading to even less information leaked to bidders or the auctioneer than mixing. As such, our results can also be interpreted as a tension between privacy and implementation: If privacy is a sufficiently large concern and bidders rely on randomization, then no deterministic choice rule is implementable. An interim strategy for bidder $i$ is a map $\sigma_{i}:\mathcal{I}_{i}\to A_{i}$ satisfying $\sigma_{i}(I_{i})\in A(I_{i})$ for all $I_{i}$. Let $\Sigma_{i}$ denote the space of interim strategies for bidder $i$. An ex-ante strategy for bidder $i$ is a map $S_{i}:\Theta_{i}\to\Sigma_{i}$ from types to interim strategies. A protocol $P$ is a pair $P=(G,\bm{S})$ where $\bm{S}=\{S_{i}\}_{i\in N}$ is a fixed profile of ex-ante strategies for $G$.
Each type profile $\bm{\theta}$ yields a sequence of information sets and bidder actions at those information sets. Let $I_{i}^{k}(\bm{\theta})$ denote the $k$th information set reached at which bidder $i$ is called to play, and let $a_{i}^{k}(\bm{\theta})$ denote their action chosen at this information set. Let $I^{k}(\bm{\theta})$ denote the $k$th information set reached in the game, and let $a^{k}(\bm{\theta})$ denote the action chosen at this information set. Let $z(\{\sigma_{i}\})_{i\in\mathcal{I}}$ denote the terminal node reached via playing interim strategies $\{\sigma_{i}\}_{i\in\mathcal{I}}$ so $z(\bm{S}(\bm{\theta}))$ is the terminal reached via playing ex-ante strategies $\bm{S}$ at type profile $\bm{\theta}$. Protocol $P$ implements social choice rule $\phi$ if for all $\bm{\theta}$, $\phi(\bm{\theta})=g(z(\bm{S}(\bm{\theta}))).$
We do not restrict the incentive properties of $P=(G,\bm{S})$, which only strengthens our two classes of results. Requiring $P$ to satisfy a particular equilibrium concept only strengthens our impossibility result. When we show protocols lie on the Privacy Frontier, we separately check that they also satisfy desirable incentive properties. [^12] In the second-price auction setting, the given protocols are ex-post incentive compatible or dominant-strategy incentive compatible. In the first-price auction setting, the given protocols form a perfect Bayesian equilibrium. Protocols that lie on the Privacy Frontier, by definition, are not privacy dominated by any protocol: they are therefore not dominated by any protocol satisfying desired incentive properties.
What does the auctioneer observe from the protocol? The auctioneer implements the choice rule by repeatedly querying a bidder, observing his action, and moving to the successor information set; this continues until the auctioneer arrives at a terminal history. This embeds an assumption that the auctioneer cannot “commit to forget” the content of a bidder’s message. Hence the auctioneer observes the full sequence of information sets reached and actions taken during the play of the game: For a given protocol $P$ and type profile $\bm{\theta}$, the auctioneer observes the game history $(I^{k}(\bm{\theta}),a^{k}(\bm{\theta}))_{k=1}^{T}$. Let $o_{A}(P,\bm{\theta})$ denote this observed history under protocol $P$ at type profile $\theta$. We say $\bm{\theta}\sim_{(P,A)}\bm{\theta}^{\prime}$ if $o_{A}(P,\bm{\theta})=o_{A}(P,\bm{\theta}^{\prime})$: If the auctioneer cannot distinguish between type profiles $\bm{\theta}$ and $\bm{\theta}^{\prime}$, then they are equivalent under $\sim_{(P,A)}$. To implement $\phi$, the auctioneer must at least be able to distinguish between type profiles that prescribe different outcomes. Letting $\mathcal{P}(\sim)$ denote the partition induced by an equivalence relation $\sim$, this condition can be stated in the language of partitions: $\mathcal{P}(\sim_{(P,A)})$ must be finer than $\mathcal{P}(\sim_{\phi})$ where $a\sim_{\phi}b$ if and only if $\phi(a)=\phi(b)$. [^13] Elements $a,b\in\Theta$ are defined to be in the same cell of $\mathcal{P}(\sim)$ if and only if $a\sim b$. Partition $\mathcal{P}$ is finer than partition $\mathcal{P}^{\prime}$ if for any cell $X$ in $\mathcal{P}$, there exists a cell $Y$ in $\mathcal{P}^{\prime}$ such that $X\subseteq Y$.
What do bidders observe from the protocol? First, the bidder can distinguish some information from the outcome alone. Let $\Omega_{i}$ be a partition of $X$, with cells denoted by $\omega_{i}$, describing the outcomes that bidder $i$ can distinguish from the outcome alone: in the auction setting, we assume that a bidder observes his allocation and payment at the conclusion of the auction. [^14] Suppose $\{(q_{j},t_{j})\}_{j=1}^{N}\in X$ represents an outcome with allocation $q_{i}\in\{0,1\}$ and transfer $t_{i}$ for bidder $i$. Then $\{(q_{j},t_{j})\}_{j=1}^{N}\sim_{\Omega_{i}}\{(q_{j}^{\prime},t_{j}^{\prime})\}_{j=1}^{N}$ if $q_{i}=q_{i}^{\prime}$ and $t_{i}=t_{i}^{\prime}$. This corresponds to bidder $i$ being unable to inherently distinguish between two different outcomes which both result in them losing the auction. Furthermore, each bidder observes his own sequence of communication with the auctioneer. In total, bidder $i$ observes $(\theta_{i},(I_{i}^{k}(\bm{\theta}),a_{i}^{k}(\bm{\theta}))_{k=1}^{T_{i}},\omega_{i})$: the sequence of the information sets at which he has been called to play, the actions he has taken at those information sets, and what she observes about the final outcome, $\omega_{i}\in\Omega_{i}$. [^15] Under this formulation, the bidder can observe their sequence of play, but not calendar time. This is also the definition of an experience, restricted to the end of the game, in Li (2017). Let $o_{i}(P,\bm{\theta})$ denote bidder $i$’s observation in protocol $P$ when bidders have type profile $\bm{\theta}\in\bm{\Theta}$. We may then define what bidder $i$ learns about other bidders under protocol $P$ by defining the following equivalence relation. We say $\bm{\theta}\sim_{(P,i)}\bm{\theta}^{\prime}$ if $o_{i}(P,\bm{\theta})=o_{i}(P,\bm{\theta}^{\prime})$. That is, the bidder can distinguish between $\bm{\theta}$ and $\bm{\theta}^{\prime}$ in protocol $P$ if and only if they induce different experiences for bidder $i$. Similarly, $\sim_{(P,i)}$ induces a partition $\mathcal{P}(\sim_{(P,i)})$ over $\bm{\Theta}$ for $i$.
4 Privacy
We adopt the language of Haupt and Hitzig (2025) to define privacy from the auctioneer. Some privacy violations may be necessary to implement a given $\phi$, but other violations are not: the auctioneer learns information about a bidder that is not pivotal to the outcome.
Definition 1 (Contextual Privacy).
Protocol $P$ produces a contextual privacy violation for bidder $i$ at type profiles $(\theta_{i},\bm{\theta}_{-i}),(\theta_{i}^{\prime},\bm{\theta}_{-i})$ if
$(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})\text{ but }\phi(\theta_{i},\bm{\theta}_{-i})=\phi(\theta_{i}^{\prime},\bm{\theta}_{-i}).$
Protocols can be partially ordered via inclusion on the type profiles where contextual privacy violations occur. A protocol produces no contextual privacy violations for bidder $i$ if the auctioneer only learns from bidder $i$ exactly what is needed to compute the choice rule. For many choice rules, any protocol produces some contextual privacy violation. However, there are still maximally contextually private protocols; as some violations are inevitable, the set of maximally contextually private protocols is often large, each producing contextual privacy violations at different type profiles.
In contrast, bidder information leakages capture our notion of protecting one bidder’s private information from other bidders. Similar to how the auctioneer must learn some components of bidders’ private information to implement $\phi$, bidders can inherently learn about other bidders’ private information from the outcome of $\phi$. However, information leakages go beyond what bidders can inherently—and inevitably—learn.
Definition 2 (Bidder Information Leakage).
Protocol $P$ produces a bidder information leakage to bidder $i$ at type profiles $(\theta_{i},\bm{\theta}_{-i}),(\theta_{i},\bm{\theta}_{-i}^{\prime})$ if
$\phi(\theta_{i},\bm{\theta}_{-i})\sim_{\Omega_{i}}\phi(\theta_{i},\bm{\theta}_{-i}^{\prime})\text{ but }(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,i)}(\theta_{i},\bm{\theta}_{-i}^{\prime})$
A minimally bidder-informative protocol produces no bidder information leakages to any bidder. For any choice rule $\phi$, one way to attain minimal bidder-informativeness is via a sealed-bid protocol, where each bidder directly reports their type to the auctioneer and the auctioneer chooses the outcome. Such a protocol reveals no information to bidders beyond what they observe of the outcome, because each bidder has only one information set. [^16] Strictly speaking, multiple sealed-bid protocols exist since the auctioneer can query bidders in any arbitrary order. These protocols are all equivalent. In a sealed-bid protocol, the auctioneer perfectly learns bidders’ values, and bidders are minimally informed. As such, sealed-bid protocols are minimally contextually private, foreshadowing our impossibility results in the next section. Sealed-bid protocols also are not the unique minimally bidder-informative protocols: Recall that an ascending second-price auction leaks the same amount of information to bidders as sealed-bid second price auctions. In general, protocols that leak information to bidders are minimally bidder-informative if bidders still learn that information solely from their observations of the outcome.
4.1 Impossibility Theorem
We provide a general condition under which contextual privacy and minimal bidder informativeness conflict: the indistinguishable corners condition.
Definition 3.
A choice rule $\phi$ violates the indistinguishable corners condition for bidder $i$ at type profiles $\{\theta_{i},\theta_{i}^{\prime}\}\times\{\bm{\theta}_{-i},\bm{\theta}_{-i}^{\prime}\}$ if:
  1. 1.
    At $\bm{\theta}_{-i}$, the auctioneer should not distinguish between $\theta_{i}$ and $\theta_{i}^{\prime}$:
    $\phi(\theta_{i},\bm{\theta_{-i}})=\phi(\theta_{i}^{\prime},\bm{\theta_{-i}});$
  2. 2.
    At $\theta_{i}$, bidder $i$ should not distinguish between $\bm{\theta}_{-i}$ and $\bm{\theta}_{-i}^{\prime}$:
    $\phi(\theta_{i},\bm{\theta_{-i}})\sim_{\Omega_{i}}\phi(\theta_{i},\bm{\theta_{-i}}^{\prime});$
  3. 3.
    At $\bm{\theta}_{-i}^{\prime}$, the auctioneer must distinguish between $\theta_{i}$ and $\theta_{i}^{\prime}$:
    $\phi(\theta_{i},\bm{\theta_{-i}}^{\prime})\neq\phi(\theta_{i}^{\prime},\bm{\theta_{-i}}^{\prime}).$
We say that $\phi$ violates the indistinguishable corners condition for bidder $i$ if there exists some type profile where the above holds.
$\alpha$$\alpha$$\beta$$\gamma$$\theta_{i}$$\theta_{i}^{\prime}$$\bm{\theta}_{-i}$$\bm{\theta}_{-i}^{\prime}$
Figure 1: Indistinguishable Corners Condition Violation
Figure˜1 illustrates a violation of the indistinguishable corners condition. $\alpha,\beta,\gamma\in X$ represent outcomes of the choice rule $\phi$. The shaded colors represent which outcomes bidder $i$ can distinguish: in particular, $\alpha\sim_{\Omega_{i}}\beta$.
This is a generalization of the corners condition, which requires the stronger condition that $\phi(\theta_{i},\theta_{j},\bm{\theta}_{-ij})=\phi(\theta_{i},\theta_{j}^{\prime},\bm{\theta}_{-ij})$, instead of $\phi(\theta_{i},\theta_{j},\bm{\theta}_{-ij})\sim_{\Omega_{i}}\phi(\theta_{i},\theta_{j}^{\prime},\bm{\theta}_{-ij})$; each outcome is trivially in the same partition as itself. In many cases, there are strictly more violations of the indistinguishable corners condition than the corners condition. In the first-price auction, there are no violations of the corners condition but many violations of the indistinguishable corners condition (Lemma 2). In the second-price auction, there are violations of the corners condition, but also distinct violations of the indistinguishable corners condition (Lemma 4).
Suppose the designer implements some choice rule $\phi$ that violates the indistinguishable corners condition. The driving force behind our impossibility theorem is that to prevent bidders from learning information about other bidders’ types, the auctioneer must ask superfluous questions, learning unnecessary information. For intuition, consider Figure˜1. Suppose bidder $i$ has type $\theta_{i}$. If other bidders have type profile $\bm{\theta}_{-i}^{\prime}$, then the auctioneer must eventually ask bidder $i$ to distinguish type $\theta_{i}$ and $\theta_{i}^{\prime}$, in order for the auctioneer to determine the outcome. This means that if the auctioneer never asks bidder $i$ to distinguish $\theta_{i}$ and $\theta_{i}^{\prime}$, then bidder $i$ understands the type profile is not $(\theta_{i},\bm{\theta}_{-i}^{\prime})$. This is additional information that bidder $i$ obtains from the protocol being run. An auctioneer can instead provide each bidder strictly less information by running a sealed-bid protocol: in a sealed-bid protocol, bidder $i$ cannot distinguish between $\phi(\theta_{i},\bm{\theta}_{-i})$ and $\phi(\theta_{i},\bm{\theta}_{-i}^{\prime})$. This yields a tradeoff: given type profile $\bm{\theta}_{-i}$, either the auctioneer never distinguishes between $\theta_{i}$ and $\theta_{i}^{\prime}$ (which reveals additional information to bidder $i$), or the auctioneer eventually distinguishes between $\theta_{i}$ and $\theta_{i}^{\prime}$ (which produces a contextual privacy violation). Theorem 1 formalizes this insight.
Theorem 1 (Impossibility).
Suppose $\phi$ violates the indistinguishable corners condition for bidder $i$. No protocol $P$ that implements $\phi$ can satisfy both:
  1. (i)
    $P$ produces no contextual privacy violations for bidder $i$.
  2. (ii)
    $P$ produces no bidder information leakages for any bidder.
The proof of Theorem 1 relies on the following Lemma formalizing the tension between auctioneer and bidder privacy, which follows from the same intuition above. Given some set of auctioneer privacy violations, some instances of bidder information leakages are forced, and vice versa. Lemma 1 is also the workhorse result allowing us to show that certain protocols are on the frontier.
Lemma 1.
Consider $\theta_{i},\theta_{i}^{\prime}\in\Theta_{i}$ and $\bm{\theta}_{-i},\bm{\theta}_{-i}^{\prime}\in\bm{\Theta}_{-i}$. If $(\theta_{i},\bm{\theta}_{-i})\sim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$ and $(\theta_{i},\bm{\theta}_{-i}^{\prime})\nsim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$ then $(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,i)}(\theta_{i},\bm{\theta}_{-i}^{\prime})$ and $(\theta_{i}^{\prime},\bm{\theta}_{-i})\nsim_{(P,i)}(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$. Conversely, if $(\theta_{i},\bm{\theta}_{-i})\sim_{(P,i)}(\theta_{i},\bm{\theta}_{-i}^{\prime})$ then $(\theta_{i},\bm{\theta}_{-i})\sim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$ if and only if $(\theta_{i},\bm{\theta}_{-i}^{\prime})\sim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$.
Theorem˜1 implies that a designer need not search for a protocol that preserves both desirable privacy properties; such a protocol does not exist. [^17] The strength of Theorem 1 is that a designer cannot fully preserve any bidder’s privacy from the auctioneer without compromising privacy between bidders. For example, the ascending-join protocol of Haupt and Hitzig (2025) chooses one bidder to protect from the auctioneer, by producing no contextual privacy violations for the chosen bidder. Theorem 1 implies that the ascending-join protocol is not minimally bidder informative. We also use Theorem˜1 to show that several protocols which preserve privacy from the auctioneer may not be preferred by a privacy-conscious designer. In a setting that violates the indistinguishable corners condition, if a protocol $P$ prevents the auctioneer from learning any unnecessary information about some bidder, Theorem˜1 implies we can produce a protocol that is strictly less bidder informative than $P$. This complicates the privacy argument for using such protocols without additionally considering privacy from whom.
4.2 The Privacy Frontier
In light of Theorem˜1, no one protocol is best at protecting privacy from both the auctioneer from bidders. However, protocols can still be partially ordered (via set inclusion) based on the contextual privacy violations and bidder information leakages they result in. If the bidder leakages to bidder $i$ that protocol $P$ produces are a superset of the bidder leakages to bidder $i$ that protocol $P^{\prime}$ produces, then bidder $i$ learns a finer partition of the type space under $P$ than he learns under $P^{\prime}$. A similar definition holds for what the auctioneer learns.
Definition 4.
Protocol $P$ reveals (strictly) more to bidder $i$ than protocol $P^{\prime}$ if $\mathcal{P}(\sim_{(P,i)})$ is (strictly) finer than $\mathcal{P}(\sim_{(P^{\prime},i)})$.
Definition 5.
Protocol $P$ reveals (strictly) more to the auctioneer than protocol $P^{\prime}$ if $\mathcal{P}(\sim_{(P,A)})$ is (strictly) finer than $\mathcal{P}(\sim_{(P^{\prime},A)})$.
One protocol dominates another from a privacy perspective if it has fewer contextual privacy violations and bidder information leakages. If a protocol is not dominated by any other protocol, it lies on the Privacy Frontier.
Definition 6 (Privacy Domination).
A protocol $P$ is (strictly) privacy dominated by protocol $P^{\prime}$ if $P$ reveals more to each bidder than $P^{\prime}$ and $P$ reveals more to the auctioneer than $P^{\prime}$ (with at least one strict).
Definition 7 (Privacy Frontier).
A protocol $P$ which implements $\phi$ lies on the Privacy Frontier for $\phi$ if there does not exist a protocol that implements $\phi$ which strictly privacy dominates $P$.
At the endpoints of this frontier are minimally auctioneer-informative and minimally bidder-informative protocols. Since sets are only partially ordered by inclusion, ordering protocols by bidder or auctioneer privacy is also a partial ordering. As such, the Privacy Frontier will generally have many incomparable protocols on the frontier.
5 First-Price Auctions
In this section, we consider an auctioneer implementing a first-price auction choice rule. As discussed in the introduction, a sealed-bid first-price auction reveals more information to the auctioneer, but a descending auction reveals more information to bidders. We first show that a first-price auction choice rule violates the indistinguishable corners condition for every bidder. With Theorem˜1, this implies that the descending auction is not minimally bidder-informative. We then show that the descending auction and sealed-bid first-price auction both lie on the Privacy Frontier. Finally, we introduce two novel protocols that also lie on the Privacy Frontier.
Given our focus on deterministic protocols, we consider a first-price auction rule with lexicographic tie-breaking. The type spaces are $\Theta_{i}\subset\mathbb{R}_{+}$ for $i\in\{1,\dots,n\}$. Let $\triangleright$ denote the lexicographic order on $\mathbb{R}_{+}\times N$ so the bidder $j$ that has the largest bid in the $\triangleright$ order wins.
Definition 8.
Given strictly increasing bidding functions $b_{i}:\Theta_{i}\to\mathbb{R}_{+}$ for $i\in\{1,\dots n\}$, the first-price auction choice rule $\phi^{FPA}$ is defined by
$\phi^{FPA}(\bm{\theta})=(\phi_{1}^{FPA},\dots,\phi_{n}^{FPA})(\bm{\theta)}=((q_{1},t_{1}),\dots,(q_{n},t_{n}))(\bm{\theta})$
with
$(q_{j},t_{j})(\bm{\theta})=\begin{cases}(1,b_{i}(\theta_{i}))&\text{ if }i=\min\{k:b_{k}(\theta_{k})=\max_{j\in N}b_{j}(\theta_{j})\}\\ (0,0)&\text{ otherwise}\end{cases}$
Defining the first-price auction allocation rule with respect to given bid functions disentangles incentive compatibility and privacy concerns. Whether $\phi^{FPA}$ is incentive-compatible depends on the choice of bid functions $\{b_{i}\}$, while for any $\{b_{i}\}$, the same privacy concerns persist. [^18] We also make an interleaving assumption: for all $i\in N$ and all $\theta_{i},\theta_{i}^{\prime}\in\Theta_{i}$ with $\theta_{i}<\theta_{i}^{\prime}$, there exists $j\in N$ and $\theta_{j}\in\Theta_{j}$ such that $(b_{i}(\theta_{i}),i)\triangleleft(b_{j}(\theta_{j}),j)\triangleleft(b_{i}(\theta_{i}^{\prime}),i)$. This assumption is motivated by the first-price auction setting: no bidder can reduce their bid without affecting the outcome at some realized type profile. For example, symmetric type spaces and symmetric bidding functions satisfy this assumption. We show that $\phi^{FPA}$ violates the indistinguishable corners condition for every bidder $i$, implying that our impossibility theorem applies for each bidder.
Lemma 2.
Suppose $\Theta_{i}=\Theta$ and $b_{i}=b$ for all $i$, with $|\Theta|>3$. [^19] This condition ensures that there is sufficient overlap between bidders’ type spaces. Weaker assumptions would also be sufficient: as we show in the Appendix, the symmetry assumption is not necessary for violations. Then $\phi^{FPA}$ violates the indistinguishable corners condition for every bidder.
To see why $\phi^{FPA}$ violates the indistinguishable corners, consider the pair of values $\theta_{i},\theta_{i}^{\prime}\in\Theta_{i}$ with $\theta_{i}<\theta_{i}^{\prime}$. If the highest bid among other bidders is some $b_{j}(\theta_{j})>b_{i}(\theta_{i}^{\prime})$, then bidder $i$ is not pivotal to the outcome at these types: she will lose the auction regardless. However, if the highest bid among other bidders is some $b_{j}(\theta_{j})\in(b_{i}(\theta_{i}),b_{i}(\theta_{i}^{\prime}))$, then bidder $i$ is pivotal to the outcome at these types: whether she has value $\theta_{i}$ or $\theta_{i}^{\prime}$ determines the winner of the auction. Finally, if bidder $i$ loses the auction, all she observes is that she loses the auction: this is why the indistinguishable corners condition is violated, despite the corners condition not being violated. [^20] The existence of a protocol that produces no contextual privacy violations (the descending auction protocol) is consistent with the absence of corners condition violations in $\phi^{FPA}$. However, because $\phi^{FPA}$ violates the indistinguishable corners condition, we may apply Theorem 1. Given that our impossibility theorem applies, what protocols are on the Privacy Frontier? Two natural candidates are the descending protocol and the sealed-bid protocol.
Definition 9.
The descending auction protocol for $\phi^{FPA}$ successively queries bidders as follows. Let $\tilde{\bm{\Theta}}=\bigtimes_{i=1}^{n}\tilde{\Theta}_{i}$ denote the potential type profiles remaining at each stage. At each stage, there exists bidder $i$ and a minimum threshold $b_{i}(\hat{\theta}_{i})\in\{b_{i}(\theta_{i})\}_{\theta_{i}\in\tilde{\Theta}_{i}}$ for which $\bm{\theta}_{-i}\in\tilde{\bm{\Theta}}_{-i}$ and $b_{i}(\theta_{i})\geq b_{i}(\hat{\theta}_{i})$ implies that $i$ wins the good. Ask bidder $i$ if their bid is at least $b_{i}(\hat{\theta}_{i})$. If so, award the object to bidder $i$ at price $b_{i}(\hat{\theta}_{i})$; otherwise, update $\tilde{\bm{\Theta}}$ and continue.
The descending protocol only reveals the winner’s value and that losers’ values were below the winner’s value to the auctioneer so it produces no contextual privacy violations.
Lemma 3.
(Haupt and Hitzig (2025)) The descending auction protocol for a first-price auction produces no contextual privacy violations for any bidder.
By our impossibility theorem (Theorem˜1) and Lemma˜2, this implies that the descending auction protocol is not minimally bidder-informative: When a losing bidder $i$ observes that the clock has stopped at some $b_{i}(\theta)$, they can immediately infer that the winner has value $\theta$.
Corollary 1.
The descending auction protocol is not minimally bidder-informative. Any protocol on the Privacy Frontier for $\phi^{FPA}$ either reveals strictly more information to bidders than the sealed-bid first-price auction protocol, or reveals strictly more information to the auctioneer than the descending auction protocol.
The second sentence of the corollary holds for the same reason: no auction format can simultaneously be minimally bidder-informative and produce no contextual privacy violations. Despite this, the descending protocol is on the frontier, being the only fully contextually private protocol. Sealed-bid auctions are also on the frontier (at the other end), being the only minimally bidder-informative protocol.
Proposition 1.
The descending protocol lies on the Privacy Frontier for $\phi^{FPA}$.
Proposition 2.
The sealed-bid protocol lies on the Privacy Frontier for $\phi^{FPA}$.
Although the sealed-bid protocol is minimally bidder-informative, this does not imply that it lies on the Privacy Frontier. [^21] For example, in the second-price auction setting, the sealed-bid auction does not lie on the Privacy Frontier. Instead, we must show that all information learned by the auctioneer in the sealed-bid protocol cannot be omitted without giving bidders additional information. We use Lemma˜1 to show that preserving minimal bidder-informativeness requires the auctioneer to fully distinguish bidders’ types.
5.1 Descending Then Sealed-Bid Protocol
We now introduce two new protocols to implement $\phi^{FPA}$. One protocol which reveals more information to bidders when bidders have high values, but reveals more information to the auctioneer when bidders have low values is the descending-then-sealed-bid protocol, which proceeds as follows:
  1. 1.
    Run a descending auction until some predetermined threshold.
  2. 2.
    If some bidder claimed the object in stage 1, the auction has finished.
  3. 3.
    Otherwise, run a sealed bid auction.
To illustrate, suppose there are two bidders, $A$ and $B$, each with values in $\{1,2,3,4,5\}$. Figure˜2 plots $\phi^{FPA}$, where $i,v$ denotes bidder $i$ winning at a price of $b_{i}(v)$.
12345Bidder AA12345Bidder BBA,1A,2A,3A,4A,5B,2B,3B,4B,5
Figure 2: Choice rule for $\phi^{FPA}$
Let $\hat{X}$ be the set of outcomes corresponding to the winner’s value being less than or equal to $3$, so $(\phi^{FPA})^{-1}(\hat{X})=\hat{\bm{\Theta}}=\{1,2,3\}\times\{1,2,3\}$. [^22] In the multi-bidder case, we can take $\hat{X}$ to still be the set of outcomes corresponding to the price being less than or equal to some threshold $p$, leading $\hat{\Theta}=\prod_{i}\{\theta_{i}:b_{i}(\theta_{i})<p\}$. Construct $\phi^{\prime}$ by setting $\phi^{\prime}(\bm{\theta})=\hat{x}$ for all $\bm{\theta}\in\hat{\bm{\Theta}}$. Equivalently, suppose values of $1,2$, and $3$ were homogenized into one type, $\leq 3$.” Then $\phi^{\prime}$ has the same structure as a first-price auction on this simpler type space. Splitting up the original implementation problem into implementing $\phi^{\prime}$ on $\bm{\Theta}$ and $\phi$ on $\hat{\bm{\Theta}}$ yields the decomposition in Figure˜3.
[…]
Figure 3: Splitting $\phi^{FPA}$
Suppose the designer cares about privacy from bidders more than privacy from the auctioneer at low valuations (perhaps because post-auction competition is more intense at these valuations) but cares more about privacy from the auctioneer than privacy from bidders at high valuations (perhaps because bidders with especially high valuations in this auction are more likely to win future auctions, which the auctioneer can use to their advantage). The designer may then choose to use a descending protocol on $\phi^{\prime}$ with the simpler type space and a sealed-bid protocol on $\phi$ restricted to $\hat{\bm{\Theta}}$. This produces the following “descending then sealed-bid” protocol:
12345Bidder AA12345Bidder BBA,4A,5B,4B,5A,1A,2A,2B,2A,3A,3A,3B,3B,3
Figure 4: Partition induced by descending then sealed-bid protocol
Definition 10.
For a threshold price $p$, let $\hat{\Theta}_{j}:=\{\theta_{j}\in\Theta_{j}:b_{j}(\theta_{j})\leq p\}$ and let $\hat{\bm{\Theta}}=\bigtimes_{j}\hat{\Theta}_{j}$. Define $\phi^{\prime}:\bm{\Theta}\to X\cup\{\hat{x}\}$ as $\phi^{\prime}(\theta)=\hat{x}$ if $\theta\in\hat{\bm{\Theta}}$ and $\phi^{\prime}(\theta)=\phi(\theta)$ otherwise. The descending then sealed-bid protocol runs the descending protocol until the auctioneer either determines the outcome or determines $\theta\in\hat{\bm{\Theta}}$. If $\theta\in\hat{\bm{\Theta}}$, the auctioneer then runs a sealed-bid protocol to determine the outcome.
Proposition 3.
The descending then sealed-bid protocol lies on the Privacy Frontier for $\phi^{FPA}$.
By choosing $\hat{\Theta}$, the designer can change when the auction switches from the descending stage to the sealed-bid stage. If one descending then sealed-bid protocol descends for longer than another, then it preserves more privacy from the auctioneer at the cost of bidders learning more. At the two extremes of this tradeoff are the descending protocol and sealed-bid protocol themselves.
5.2 Descending-Holdout Protocol
In some settings, it may be especially important to protect one particular agent’s privacy, and less important to protect other agents’ privacy. [^23] For example, one bidder may have particularly sensitive data or technological information that affects their bid. As another example, one bidder may be a long-lived agent concerned about repeated interactions with third parties, while other bidders are short-lived agents bidding in only one interaction. With this motivation, we introduce the descending-holdout protocol to implement $\phi^{FPA}$. The descending-holdout protocol reveals less information about one chosen bidder than both the sealed-bid first-price protocol and the descending protocol. In contrast with either of these existing auction formats, the designer can now protect the chosen agent’s privacy from both the auctioneer and other bidders.
The descending-holdout protocol fixes one bidder $k$ to be held out. The auctioneer first solicits sealed bids from all bidders $j\neq k$, and selects the highest bidder as the tentative winner $i\in N$. The auctioneer then runs a descending auction for bidder $k$ with a secret reserve price of $b_{i}(\theta_{i})$. If bidder $k$ claims the item at a price above $b_{i}(\theta_{i})$, bidder $k$ wins the item. Otherwise, bidder $i$ wins the item.
Definition 11.
The descending-holdout protocol fixes bidder $k$ as the “holdout” bidder. Each bidder $j\neq k$ has one information set with actions $b_{j}(\Theta_{j})$. The tentative winner is bidder $i=\operatorname*{\arg\!\max}_{j\neq k}\{(b_{j}(\theta_{j}),j)\}$.
The auctioneer then sequentially asks bidder $k$ if she will claim the item for a price of $b_{k}(\tilde{\theta}_{k})$, with $\tilde{\theta}_{k}$ beginning at $\max\{\Theta_{k}\}$ and descending in each step. The protocol ends when either $k$ claims the item or the tentative winner beats the holdout bidder: $(b_{i}(\tilde{\theta}_{i}),i)\triangleleft(b_{k}(\theta_{k}),k)$. If $k$ claims the item, she wins the item at price $b_{k}(\tilde{\theta}_{k})$. Otherwise, bidder $i$ wins the item at price $b_{i}(\theta_{i})$.
Figure˜5 displays the partition of information the auctioneer learns, $\mathcal{P}(\sim_{P,A})$, in the descending-holdout protocol with Bidder B as the holdout bidder. We then show that the descending-holdout protocol lies on the Privacy Frontier.
12345Bidder AA12345Bidder BBA,4A,5B,4B,4B,4B,5B,5B,5B,5A,1A,2B,2A,3B,3B,3
Figure 5: Partition induced by descending-holdout protocol
Proposition 4.
The descending-holdout protocol lies on the Privacy Frontier for $\phi^{FPA}$.
The descending-holdout protocol highlights the tradeoffs that a designer makes in choosing which bidder’s information to protect. Compared to the sealed-bid protocol, the descending-holdout protocol reveals less information about the holdout bidder $k$ to the auctioneer: the auctioneer only learns bidder $k$’s exact value if $k$ wins the auction. Compared to the descending protocol, the descending-holdout protocol reveals less information about bidder $k$ to other bidders: if bidder $k$ wins the auction, other bidders still only learn that some other bidder won the auction, and in particular, they do not learn bidder $k$’s exact winning bid. In this sense, the descending-holdout protocols reveals minimal necessary information about bidder $k$ to both the auctioneer and other bidders. [^24] In addition, the descending-holdout protocol satisfies Bayesian incentive compatibility for all bidders. For all bidders $j\neq k$, the protocol is strategically equivalent to a sealed-bid first-price auction. For bidder $k$, the protocol is strategically equivalent to a descending auction. Finally, note that our core tradeoff still holds: The descending-holdout protocol protects bidder $k$’s information from both the auctioneer and other bidders, but the auctioneer and bidder $k$ learn about other bidders’ values.
Figure˜6 depicts a simplified two-dimensional representation of the Privacy Frontier for first-price auctions. Descending auctions are minimally auctioneer informative, while sealed-bid auctions are minimally bidder informative but also maximally auctioneer informative. Descending-then-sealed-bid auctions can interpolate between the two extremes based on when the sealed-bid portion of the protocol begins. Finally, descending-holdout auctions interpolate between descending and sealed-bid protocols in a different manner and cannot be ordered against descending-then-sealed-bid auctions as they treat bidders differentially.
Privacy from AuctioneerPrivacy from BiddersDescendingSealed-bidDescending then Sealed-bidDescending-Holdout
Figure 6: Privacy Frontier for First-Price Auctions
6 Second-Price Auctions
Consider an auctioneer implementing a second-price auction. The second-price auction choice rule violates the indistinguishable corners condition, so our impossibility theorem applies. The ascending auction and the ascending-join auction lie on the Privacy Frontier, and the sealed-bid second-price auction is privacy dominated by the ascending auction.
Definition 12.
The second-price auction $\phi^{SPA}$ is defined by
$\phi^{SPA}(\bm{\theta})=(\phi_{1}^{SPA},\dots,\phi_{n}^{SPA})(\bm{\theta)}=((q_{1},t_{1}),\dots,(q_{n},t_{n}))(\bm{\theta})$
with
$(q_{j},t_{j})(\bm{\theta})=\begin{cases}(1,\max_{j\neq i}\theta_{j})&\text{ if }i=\min\{k:\theta_{k}=\max_{j\in N}\theta_{j}\}\\ (0,0)&\text{ otherwise}\end{cases}$
Unlike the first-price auction, the second-price auction violates the standard corners condition. [^25] Our impossibility theorem still has implications beyond Haupt and Hitzig (2025), even in settings with only violations of the standard corners condition. In such a setting, the designer can construct a protocol (in this case, the ascending-join protocol) which produces no contextual privacy violations for one protected bidder; our impossibility theorem shows that such a protocol cannot be minimally bidder-informative. $\phi^{SPA}$ also has indistinguishable corners conditions violations which are not corners conditions violations. Intuitively, whenever a bidder is relevant in determining the price but not who receives the good, they become pivotal between outcomes indistinguishable to them.
Lemma 4.
Suppose $\Theta_{i}=\Theta$ for all $i$ and $|\Theta|>5$. Then for each bidder, there exists type profiles that violate the indistinguishable corners condition for $\phi^{SPA}$ but do not violate the corners condition.
Two protocols that the designer could consider using are the sealed-bid protocol and the ascending protocol. The ascending protocol increases the price offered to each bidder until all but one of the bidders drops out.
Definition 13.
The ascending auction protocol has information sets $\mathcal{I}_{i}=\{I_{\theta_{i}}\}_{\theta_{i}\in\Theta_{i}}$ for bidder $i$, so each information set is identified with a clock price (however, note that there is no publicly observable clock). $A(I)=\{\text{Stay},\text{Quit}\}$ for all $I\in\mathcal{I}_{i}$. Let $p(i)$ denote the highest information set that bidder $i$ has reached in the auction. At each stage, the bidder with the lowest $p(i)$ still in the auction is called to play. If the bidder has value $\theta_{i}\leq p(i)$, he quits. Otherwise, $p(i)$ is incremented. The last remaining bidder wins the good at the highest dropout price among the other bidders.
The ascending-join protocol from Haupt and Hitzig (2025) instead reveals strictly less information to the auctioneer than a sealed-bid auction (for any $N$) or ascending auction (for $N>2$). The ascending-join protocol begins by increasing the clocks for two bidders until one bidder drops out, then successively asks new bidders whether their value is above the current second-highest value. We defer the formal definition of the ascending-join protocol to the Appendix. This protocol is one way to minimize what the auctioneer learns about losers’ values. However, it reveals more information to bidders than the sealed-bid auction: if a bidder’s first query is “Is your value above pp?” the bidder immediately learns that the second-highest value among all previous bidders is $p$. The ascending-join protocol therefore fails to be minimally bidder-informative; this also follows immediately from our impossibility theorem by noting that there is no contextual privacy violation for the final bidder who joins the auction.
Proposition 5.
For $N>2$, the ascending-join auction is not minimally bidder-informative. In particular, the ascending-join auction does not privacy dominate the sealed-bid second-price auction.
This result shows that a privacy-conscious designer need not prefer the ascending-join auction format. If privacy from bidders is more important than privacy from the auctioneer, she may instead prefer the sealed-bid second-price auction protocol. However, the sealed-bid protocol is in turn privacy dominated by the ascending auction, and thus does not lie on the Privacy Frontier. That is, a designer who solely values privacy for all parties involved should always weakly prefer using the ascending auction instead of the sealed-bid second-price auction.
Proposition 6.
The sealed-bid second-price auction is privacy dominated by the ascending auction.
The ascending auction may initially appear to reveal more to the winning bidder than the sealed-bid auction does. However, the only additional information that the ascending auction reveals to the winner is the second-highest value. Since the winner learns this information via their payment regardless of the protocol used, the ascending auction reveals the same information to bidders as the sealed-bid auction does. We then show that the ascending protocol and the ascending-join protocol lie on the Privacy Frontier.
Proposition 7.
The ascending auction lies on the Privacy Frontier for $\phi^{SPA}$.
As described above, the ascending auction is also minimally bidder-informative. However, as Proposition˜6 illustrates, a minimally bidder-informative protocol need not lie on the Privacy Frontier. The ascending auction reveals all losing bidders’ values to the auctioneer, and reveals no unnecessary information about the winning bidder to the auctioneer. The difficulty in the proof is therefore to show that in any minimally bidder-informative protocol which implements $\phi^{SPA}$, the auctioneer must learn the value of all losing bidders.
In our proof, we use Lemma˜1 to show that the auctioneer is forced to distinguish all losers’ values. Concretely, consider a type profile $\bm{\theta}_{-i}$ and $\theta_{i},\theta_{i}^{\prime}\in\Theta_{-i}$ such that bidder $i$ loses when she has value $\theta_{i}$ or $\theta_{i}^{\prime}$. We construct a type profile $\bm{\theta}_{-i}^{\prime}$ such that bidder $i$ is pivotal at type profile $\bm{\theta}_{-i}^{\prime}$, that is, that these type profiles violate the indistinguishable corners condition. Then by Lemma˜1, to maintain minimal bidder informativeness, the auctioneer must distinguish $\theta_{i}$ and $\theta_{i}^{\prime}$ at type profile $\bm{\theta}$. A generalization of this approach yields our sufficiency conditions in Section˜7. Finally:
Proposition 8.
The ascending-join protocol lies on the Privacy Frontier for $\phi^{SPA}$.
This statement is straightforward to show, though non-trivial because we must show that the designer cannot reorder the queries in the ascending-join protocol to obtain the same partition $\mathcal{P}(\sim_{(P,A)})$ while reducing the information revealed to bidders. In sum, Figure˜7 represents the Privacy Frontier for $\phi^{SPA}$.
Privacy from AuctioneerPrivacy from BiddersMinimally Bidder-Informativeascending-joinascendingsealed-bid
Figure 7: Privacy Frontier for Second-Price Auctions
7 Sufficient Conditions for the Privacy Frontier
How can a designer verify that a given protocol lies on the Privacy Frontier? This is challenging because the set of protocols implementing a choice rule is typically vast. However, we provide sufficient conditions for a protocol to lie on the privacy frontier. These conditions apply to the implementation of any choice function, even beyond the structure of an auction setting.
The strength of these conditions is that they are stated only in terms of model primitives and equivalence relations $\sim_{(P,A)}$ and $\sim_{(P,i)}$. That is, if the designer is evaluating a given protocol, she only needs to know what the auctioneer learns and what bidders learn from the protocol. She can then use our conditions to verify that the protocol is on the frontier, without further analysis of the protocol’s extensive form.
We state our conditions in two lemmas. Our first condition guarantees that any protocol which privacy dominates protocol $P$ must reveal the same information to the auctioneer as $P$. The logic is analogous to the indistinguishable corners condition. Suppose that in protocol $P$, the auctioneer distinguishes type profiles $(\theta_{i},\bm{\theta}_{-i})$ and $(\theta_{i}^{\prime},\bm{\theta}_{-i})$. When the condition of Lemma˜5 is satisfied, if any other protocol $\hat{P}$ does not have the auctioneer distinguish $(\theta_{i},\bm{\theta}_{-i})$ and $(\theta_{i}^{\prime},\bm{\theta}_{-i})$, then some bidder $i$ can now distinguish types in $\hat{P}$ that he could not distinguish in $P$. As a result, all information the auctioneer learns in $P$ is necessary to prevent bidders from learning additional information.
Lemma 5.
Let $P$ be a protocol that implements choice rule $\phi$. Suppose that for all $i\in N$, all $\theta_{i},\theta_{i}^{\prime}\in\Theta_{i}$ and all $\bm{\theta}_{-i}\in\bm{\Theta}_{i}$, if $(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$, at least one of the following holds:
  1. (a)
    $\phi(\theta_{i},\bm{\theta}_{-i})\neq\phi(\theta_{i}^{\prime},\bm{\theta}_{-i})$
  2. (b)
    there exists $\bm{\theta}_{-i}^{\prime}\in\bm{\Theta}_{-i}$ such that $\phi(\theta_{i},\bm{\theta}_{-i}^{\prime})\neq\phi(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$ and $(\theta_{i},\bm{\theta}_{-i})\sim_{(P,i)}(\theta_{i},\bm{\theta}_{-i}^{\prime})$
  3. (c)
    there exists $\bm{\theta}_{-i}^{\prime}\in\bm{\Theta}_{-i}$ such that $\phi(\theta_{i},\bm{\theta}_{-i}^{\prime})\neq\phi(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$ and $(\theta_{i}^{\prime},\bm{\theta}_{-i})\sim_{(P,i)}(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$.
If protocol $\hat{P}$ privacy dominates $P$, then $\sim_{(\hat{P},A)}$ is identical to $\sim_{(P,A)}$.
We then state an analogous result for the information that a protocol reveals to bidders. For a given description of what the auctioneer learns from $P$, any protocol which reveals that information to the auctioneer must also reveal certain information to bidders.
Lemma 6.
Let $P$ be a protocol that implements choice rule $\phi$. Consider any $i\in N$. Suppose that for all $\theta_{i}\in\Theta_{i}$ and all $\bm{\theta}_{-i},\bm{\theta}_{-i}^{\prime}\in\bm{\Theta}_{-i}$, if $(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,i)}(\theta_{i},\bm{\theta}_{-i}^{\prime})$, at least one of the following holds:
  1. (a)
    $\phi(\theta_{i},\bm{\theta}_{-i})\nsim_{\Omega_{i}}\phi(\theta_{i},\bm{\theta}_{-i}^{\prime})$
  2. (b)
    there exists $\theta_{i}^{\prime}\in\Theta_{i}$ such that $(\theta_{i},\bm{\theta}_{-i})\sim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$ and $(\theta_{i},\bm{\theta}_{-i}^{\prime})\nsim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$
  3. (c)
    there exists $\theta_{i}^{\prime}\in\Theta_{i}$ such that $(\theta_{i},\bm{\theta}_{-i}^{\prime})\sim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$ and $(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$.
If protocol $\hat{P}$ privacy dominates $P$ and $\sim_{(\hat{P},A)}$ is identical to $\sim_{(P,A)}$, then $\sim_{(\hat{P},i)}$ is identical to $\sim_{(P,i)}$.
If both conditions hold for protocol $P$, then protocol $P$ lies on the Privacy Frontier.
Theorem 2.
Suppose protocol $P$ and choice rule $\phi$ satisfy the conditions of Lemma˜5 and satisfy the conditions of Lemma˜6 for all $i\in N$. Then $P$ lies on the Privacy Frontier for $\phi$.
Proof.
(Theorem˜2.) Suppose protocol $\hat{P}$ privacy dominates $P$. By Lemma˜5, $\sim_{(\hat{P},A)}$ is identical to $\sim_{(P,A)}$. Then by Lemma˜6, $\sim_{(\hat{P},i)}$ is identical to $\sim_{(P,i)}$ for all $i\in N$. Hence $\hat{P}$ does not strictly privacy dominate $P$, and $P$ lies on the Privacy Frontier for $\phi$. ∎
Theorem˜2 is a general tool for verifying if a protocol is on the Privacy Frontier. In the auction setting, we use Theorem˜2 to verify that our proposed protocols—the descending then sealed-bid protocol and the descending-holdout protocol—are on the Privacy Frontier. It is also useful for designers analyzing other novel protocols or protocols in other settings, as it does not rely on the structure of the auction setting.
8 Conclusion
We consider a designer who constructs a deterministic bilateral communication protocol among an auctioneer and multiple bidders to implement a given choice rule. Information leakages to the auctioneer or bidders may have unintended and negative consequences in a variety of economic settings. However, the designer faces an inherent tradeoff between preventing the auctioneer from unnecessarily learning bidders’ private information and preventing bidders from unnecessarily learning each others’ private information. Protocols can be partially ordered based on how much information they reveal to the auctioneer or to other bidders; protocols for which no alternative protocols reveal less to both the auctioneer and to other bidders lie on the Privacy Frontier.
For first-price auctions, the sealed-bid protocol (which reveals minimal information to bidders) and the descending protocol (which reveals minimal information to the auctioneer) lie on the frontier. We introduce the descending then sealed-bid protocol to trade off between these two extremes in different ranges of bidder values. We also introduce the descending-holdout protocol to protect the privacy of one selected bidder. For second-price auctions, the ascending protocol and the ascending-join protocol lie on the frontier. However, our privacy considerations have bite: The sealed-bid protocol is not on the frontier for second-price auctions. Finally, we provide sufficient conditions for a designer to verify that any protocol lies on the Privacy Frontier. These tradeoffs persist beyond auction settings: our impossibility theorem holds for arbitrary choice rules. Our results highlight that the criterion of simply preserving privacy is incomplete without first answering the question of privacy from whom.
Several avenues for further research remain. Future work could utilize our sufficient conditions to establish a sharp characterization of all protocols on the frontier. To illustrate the difficulties in characterizing the frontier, consider the effects of removing a single question from a protocol. This reduces the information revealed to the auctioneer. However, it may be infeasible if future questions condition on bidders’ responses to the question. Furthermore, it may increase or decrease the information revealed to bidders, depending on whether bidders are being asked the same question for different realizations of the type profile. More generally, the space of all protocols implementing a fixed choice rule can be complex.
Future work could also connect privacy to other concerns in the extensive-form mechanism design literature. If bidders learn more along the path of play of a protocol, the protocol is less private from bidders. However, this could result in (1) bidders needing less contingent reasoning, resulting in a protocol that is strategically simpler or (2) bidders having more information to detect potential auctioneer deviations. This interplay between privacy and other economic forces is important yet underdeveloped.
Appendix A: Pruning and Incentive Compatibility
This appendix comments on the incentive properties of protocols. If the choice rule the designer seeks to implement is dominant-strategy incentive compatible, then it is without loss to restrict to protocols $P=(G,\bm{S})$ which are ex-post incentive compatible (Nagel and Saitto, 2024). Recall that $g(z)$ is the outcome at terminal history $z$ and $z(\bm{S})$ is the terminal history reached at strategies $S$.
Definition 14 (Incentive Compatibility).
A choice rule $\phi$ is dominant-strategy incentive compatible if it is a dominant strategy for each bidder to truthfully reveal their type: for all bidders $i$, $\theta_{i},\theta_{i}^{\prime}\in\Theta_{i},\bm{\theta}_{-i}\in\bm{\Theta}_{-i}$,
$u_{i}(\phi(\theta_{i},\bm{\theta}_{-i}),\theta_{i})\geq u_{i}(\phi(\theta_{i}^{\prime},\bm{\theta}_{-i}),\theta_{i}).$
A protocol $P=(G,\{S_{i}(\theta_{i})\}_{i\in N})$ is ex-post incentive compatible if it is a dominant strategy for each bidder to follow the designer’s strategy recommendation: for all bidders $i$, $\theta_{i}\in\Theta_{i},\bm{\theta}_{-i}\in\bm{\Theta}_{-i},\sigma_{i}^{\prime}\in\Sigma_{i}$,
$u_{i}(g(z(S_{i}(\theta_{i}),\bm{S}_{-i}(\bm{\theta}_{-i}))),\theta_{i})\geq u_{i}(g(z(\sigma_{i}^{\prime},\bm{S}_{-i}(\bm{\theta}_{-i}))),\theta_{i}).$
We will restrict attention to protocols where each history is reached at some type profile. Protocols where some histories are never reached at any type profile can be “pruned” to produce an alternative protocol such that the preceding condition is satisfied without changing privacy. Pruning also simplifies strategic analysis: At non-pruned protocols, deviations to histories that are never relevant may be profitable, but the designer can simply prune these branches of the game tree.
Definition 15 (Pruning).
[…]
  1. 1.
[…]
  2. 2.
    $\tilde{\mathcal{I}}_{i}=\{I\cap\tilde{H}\}_{I\in\mathcal{I}_{i}}$ for each $i\in N$.
  3. 3.
[…]
Given $P=(G,\bm{S})$ let $\tilde{P}=(\tilde{G},\bm{S})$ denote the protocol where $G$ is pruned and $\bm{S}$ remains unchanged.
Since $P$ implements $\phi$ on-path and $P=\tilde{P}$ for each realized path of play, $\tilde{P}$ implements $\phi$ as well. As such, it is without loss for the purpose of implementation to restrict to pruned protocols. It is also without loss for the purpose of analyzing privacy. Privacy violations are determined by on-path behavior so pruning off-path histories has no effect.
Proposition 9 (Pruning Preserved Privacy).
Let $\tilde{P}=(\tilde{G},\bm{S})$ be the protocol attained from pruning $P=(G,\bm{S})$. Then:
  1. 1.
    Auctioneer Privacy Equivalence: $\bm{\theta}\sim_{(P,A)}\bm{\theta}^{\prime}$ if and only if $\bm{\theta}\sim_{(\tilde{P},A)}\bm{\theta}^{\prime}$.
  2. 2.
    Bidder Privacy Equivalence: For each bidder $i$, $\bm{\theta}\sim_{(P,i)}\bm{\theta}^{\prime}$ if and only if $\bm{\theta}\sim_{(\tilde{P},i)}\bm{\theta}^{\prime}$.
This result is of independent interest beyond enabling analysis of incentive compatibility. It allows the designer to significantly simplify the search for protocols on the frontier to protocols where each history is reached at some path of play; in particular it places a finite bound on the maximal length of protocols which do not ask redundant questions.
In any pruned protocol, each history is reached on-path for some type, so deviations in the game $G$ are essentially deviations where a bidder pretends to be a different type. As such, incentive compatibility of the choice rule is able to rule out all possible deviations.
Proposition 10 (Incentive Compatibility of Pruned Protocols).
Suppose $P=(G,\bm{S})$ is pruned and implements $\phi$. If $\phi$ is dominant-strategy incentive compatible, then $P$ is dominant-strategy incentive compatible.
Protocols do not necessarily preserve Bayesian incentive compatibility, however. A protocol could provide certain bidders with pertinent information before they make a choice. As such, interim beliefs may be different under the protocol implementing the mechanism compared to its normal form game, so Bayesian incentive compatibility does not translate. This may provide another novel justification for dominant strategy incentive compatibility: Dominant strategy incentive compatibility is robust to the details of how a mechanism is actually run whereas Bayesian incentive compatibility is not. Furthermore, our impossibility theorem only gets stronger if Bayesian incentive compatibility is required as it restricts the set of potential protocols the designer may use. In our application to first-price auctions, for which the choice rule $\phi^{FPA}$ is not dominant-strategy incentive compatible, we show that certain auctions are not privacy dominated by any protocol implementing $\phi^{FPA}$. This immediately implies that these auctions are not privacy dominated by any incentive-compatible protocol implementing $\phi^{FPA}$. For additional discussion, see Fadel and Segal (2009) investigating the communication costs of requiring incentive compatibility in extensive-form protocols.
Appendix B: Omitted Proofs
Impossibility Theorem
PROOF OF LEMMA 1
Proof.
(Lemma˜1) Because $(\theta_{i},\bm{\theta}_{-i}^{\prime})\nsim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$, there exists an information set $I_{i}$ on the path of play for $(\theta_{i},\bm{\theta}_{-i}^{\prime})$ at which $S_{i}(\theta_{i})(I_{i})\neq S_{i}(\theta_{i}^{\prime})(I_{i})$. Because $(\theta_{i},\bm{\theta}_{-i})\sim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$, at every information set $I_{i}$ on the path of play for $(\theta_{i},\bm{\theta}_{-i})$, we have $S_{i}(\theta_{i})(I_{i})=S_{i}(\theta_{i}^{\prime})(I_{i})$. Hence there exists some information set $I_{i}$ on the path of play for $(\theta_{i},\bm{\theta}_{-i}^{\prime})$ which is not on the path of play for $(\theta_{i},\bm{\theta}_{-i})$, and therefore $o_{i}(\theta_{i},\bm{\theta}_{-i})\neq o_{i}(\theta_{i},\bm{\theta}_{-i}^{\prime})$. By definition, $(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,i)}(\theta_{i},\bm{\theta}_{-i}^{\prime})$. By an identical argument, there exists some information set on the path of play for $(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$ which is not on the path of play for $(\theta_{i}^{\prime},\bm{\theta}_{-i})$, and therefore $(\theta_{i}^{\prime},\bm{\theta}_{-i})\nsim_{(P,i)}(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$. The second sentence in the lemma follows from the contrapositive. ∎
PROOF OF THEOREM 1
Proof.
The proof proceeds in three steps. First, we characterize when bidder information leakages occur. Second, we use this to show that contextual privacy violations must occur if bidder information leakages do not. Finally, we connect these results to the statement of Theorem 1.
Lemma 7.
Suppose $\phi$ violates the indistinguishable corners condition for bidder $i$ at $\{\theta_{i},\theta_{i}^{\prime}\}\times\{\bm{\theta}_{-i},\bm{\theta}_{-i}^{\prime}\}$. If protocol $P$ is minimally bidder-informative and implements $\phi$, then
$(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i}).$
Proof.
(Lemma˜7) In any minimally bidder-informative protocol, bidders learn nothing except what they observe of the outcome. Now suppose $G$ is minimally bidder-informative and implements $\phi$. Consider any $\theta_{i},\theta_{i}^{\prime}\in\Theta_{i}$ and $\bm{\theta}_{-i},\bm{\theta}_{-i}^{\prime}\in\bm{\Theta}_{-i}$ which violate the indistinguishable corners condition in Definition˜3. Because protocol $P$ implements $\phi$, we have $(\theta_{i},\bm{\theta}_{-i}^{\prime})\nsim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$. By the indistinguishable corners condition, $\phi(\theta_{i},\bm{\theta}_{-i})\sim_{\Omega_{i}}\phi(\theta_{i},\bm{\theta}_{-i}^{\prime})$, and hence by $P$ being minimally bidder-informative, $o_{i}(\theta_{i},\bm{\theta}_{-i})=o_{i}(\theta_{i},\bm{\theta}_{-i}^{\prime})$. By the contrapositive of Lemma˜1, we conclude that $(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$. ∎
Suppose $\phi$ violates the indistinguishable corners condition at some type profiles $\{\theta_{i},\theta_{i}^{\prime}\}\times\{\bm{\theta}_{-i},\bm{\theta}_{-i}\}$. Consider any minimally bidder-informative protocol $P$ that implements $\phi$. By Lemma˜7, $(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$. But $\phi(\theta_{i},\bm{\theta}_{-i})=\phi(\theta_{i}^{\prime},\bm{\theta}_{-i})$ so $P$ produces a contextual privacy violation for bidder $i$. ∎
First-Price Auctions
PROOF OF LEMMA 2
Proof.
Consider any two bidders $i,j$. Choose $\theta_{i}<\theta_{j}^{\prime}<\theta_{i}^{\prime}<\theta_{j}$ and some $\bm{\theta}_{-ij}$ with $\theta_{k}\leq\theta_{i}$ for all $k\neq i,j$. Let $\bm{\theta}_{-i}=(\theta_{j},\bm{\theta}_{-ij})$ and $\bm{\theta}_{-i}^{\prime}=(\theta_{j}^{\prime},\bm{\theta}_{-ij})$.
Then $\phi(\theta_{i},\bm{\theta}_{-i})=\phi(\theta_{i}^{\prime},\bm{\theta}_{-i})$ because both cases lead to bidder $j$ winning at their bid $b_{j}(\theta_{j})$. We also have $\phi(\theta_{i},\bm{\theta}_{-i})\sim_{\Omega_{i}}\phi(\theta_{i},\bm{\theta}_{-i}^{\prime})$ because bidder $i$ loses the auction in both cases. Finally, $\phi(\theta_{i},\bm{\theta}_{-i}^{\prime})\neq\phi(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$, because bidder $i$ loses the auction in the first case and wins the auction in the second case. ∎
Note that the symmetry assumption is not necessary to create indistinguishable corners violations. Indeed, suppose $|\Theta_{j}|\geq 3$ for all $j$. Let $\theta_{j}^{(-1)}$ and $\theta_{j}^{(-2)}$ denote the lowest and second-lowest values, respectively, in $\Theta_{j}$. Let $i=\operatorname*{\arg\!\max}_{j\in N}\{(b_{j}(\theta_{j}^{(-2)}),j)\}$, with the maximum taken with respect to the $\triangleleft$ operator. By our interleaving assumptions, we may choose $\bm{\theta_{-i}}$ such that $(b_{i}(\theta_{i}^{(-2)}),i)\triangleleft\max_{j\neq i}\{(b_{j}(\theta_{j}),j)\}$. Again, by our interleaving assumption, we may choose $\bm{\theta_{-i}}^{\prime}$ such that $(b_{i}(\theta_{i}^{(-1)}),i)\triangleleft\max_{j\neq i}\{(b_{j}(\theta_{j}^{\prime}),j)\}\triangleleft(b_{i}(\theta_{i}^{(-2)}),i)$. Then by the same logic as in the proof of Lemma˜2, the profiles $\{\theta_{i}^{(-1)},\theta_{i}^{(-2)}\}\times\{\bm{\theta}_{-i},\bm{\theta}_{-i}^{\prime}\}$ form an indistinguishable corners violation.
PROOF OF PROPOSITION 1
Proof.
We show that any protocol which implements $\phi^{FPA}$ and produces no contextual privacy violations must be the descending auction protocol. Given possible type profiles $\tilde{\bm{\Theta}}$, only one bidder $i$ satisfies the condition that there exists a threshold $\hat{\theta}_{i}\in\tilde{\Theta}_{i}$ such that $\theta_{i}\geq\hat{\theta}_{i}$ and $\bm{\theta}_{-i}\in\tilde{\bm{\Theta}}_{-i}$ implies bidder $i$ wins the auction (recall that ties are broken in favor of the lowest-indexed bidder so this bidder is unique). If there were multiple such bidders for which this condition holds, we immediately obtain a contradiction since it could be that both bidders have values above their thresholds but it cannot be that both bidders win the auction.
If a protocol queries any bidder $j\neq i$, then the auctioneer distinguishes some types $\theta_{j},\theta_{j}^{\prime}\in\tilde{\bm{\Theta}}_{j}$, but $\phi(\hat{\theta}_{i},\theta_{j},\bm{\theta}_{-ij})=\phi(\hat{\theta}_{i},\theta_{j}^{\prime},\bm{\theta}_{-ij})$, and therefore produces a contextual privacy violation. Hence any fully contextually private protocol can only ask $i$ to distinguish their type. By choosing the minimal threshold $\hat{\theta}_{i}$, the auctioneers learns the minimal necessary information about bidder $i$. We conclude that any protocol with no contextual privacy violations must be the descending protocol. In particular, no protocol privacy dominates the descending protocol. ∎
PROOF OF PROPOSITION 2
Proof.
It suffices to show that any minimally bidder-informative protocol $P$ implementing $\phi^{FPA}$ must distinguish all types, that is, $(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$ for all $\theta_{i},\theta_{i}^{\prime}\in\Theta_{i}$ and all $\bm{\theta}_{-i}\in\bm{\Theta}_{-i}$.
If $\phi(\theta_{i},\bm{\theta}_{-i})\neq\phi(\theta_{i}^{\prime},\bm{\theta}_{-i})$, then any protocol implementing $\phi$ satisfies this condition. Hence it suffices to show this condition holds when $\phi(\theta_{i},\bm{\theta}_{-i})=\phi(\theta_{i}^{\prime},\bm{\theta}_{-i})$. To do so, we will construct a violation of the indistinguishable corners condition using this type profile.
Without loss of generality, suppose $\theta_{i}<\theta_{i}^{\prime}$. For any type profiles with $\phi^{FPA}(\theta_{i},\bm{\theta}_{-i})=\phi^{FPA}(\theta_{i}^{\prime},\bm{\theta}_{-i})$, we must have that bidder $i$ loses the auction in both cases (otherwise, their payment would differ in the two cases). We may then construct some type profile $\theta_{-i}^{\prime}\in\Theta_{-i}$ such that
$(\theta_{i},i)\triangleleft\max_{j\neq i}\{(\theta_{j}^{\prime},j)\}\triangleleft(\theta_{i}^{\prime},i).$
Then $\phi(\theta_{i},\theta_{-i}^{\prime})$ also results in $i$ losing the auction, and hence $\phi(\theta_{i},\theta_{-i})\sim_{\Omega_{i}}\phi(\theta_{i},\theta_{-i}^{\prime})$. However, $\phi(\theta_{i}^{\prime},\theta_{-i}^{\prime})$ results in $i$ winning the auction, and hence $\phi(\theta_{i},\theta_{-i}^{\prime})\neq\phi(\theta_{i}^{\prime},\theta_{-i}^{\prime})$.
Combining these conditions, we conclude that $\phi^{FPA}$ violates the indistinguishable corners condition at $\{\theta_{i},\theta_{i}^{\prime}\}\times\{\bm{\theta}_{-i},\bm{\theta}_{-i}^{\prime}\}$. Hence by Lemma˜7, $G$ must satisfy $(\theta_{i},\theta_{-i})\nsim_{(P,A)}(\theta_{i}^{\prime},\theta_{-i})$ as desired. ∎
PROOF OF PROPOSITION 3
Proof.
As described in the text, the descending-then-sealed-bid protocol first runs the descending protocol on choice problem $\phi^{\prime}$; if this protocol outputs $\hat{x}$, it then runs the sealed-bid protocol on $\hat{\bm{\Theta}}$. Let $P$ denote this protocol.
For the sake of contradiction, suppose some protocol $\hat{P}$ implements $\phi^{FPA}$ and reveals less information to both the auctioneer and other bidders. We first show that the auctioneer must distinguish the same type profiles with $\hat{P}$ as with $P$. In particular, to avoid revealing information to bidders in $\hat{\bm{\Theta}}$, the auctioneer must fully distinguish agents’ types in $\hat{\bm{\Theta}}$. We then show that bidders must learn the same information in $\hat{P}$ as in $P$. Thus, no protocol $\hat{P}$ can possibly dominate $P$.
Lemma 8 (Equivalent Auctioneer Privacy Violations).
$\sim_{(\hat{P},A)}$ is identical to $\sim_{(P,A)}$.
Proof.
By assumption, $\hat{P}$ has less contextual privacy violations than $P$. As such, it suffices to show that anything the auctioneer learns under $P$, they also learn under $\hat{P}$. Formally, for any $(\theta_{i},\bm{\theta}_{-i})$ and $(\theta_{i}^{\prime},\bm{\theta}_{-i})$, if $(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$ then $(\theta_{i},\bm{\theta}_{-i})\nsim_{(\hat{P},A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$.
Suppose $(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$. If $(\theta_{i},\bm{\theta}_{-i})\notin\hat{\bm{\Theta}}$ or $(\theta_{i}^{\prime},\bm{\theta}_{-i})\notin\hat{\bm{\Theta}}$, it is straightforward to verify that $\phi^{FPA}(\theta_{i},\bm{\theta}_{-i})\neq\phi(\theta_{i}^{\prime},\bm{\theta}_{-i})$: in the descending auction, the auctioneer learns only what she needs to know. Hence $(\theta_{i},\bm{\theta}_{-i})\nsim_{(\hat{P},A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$ since the auctioneer must tell these type profiles apart under any protocol that implements $\phi^{FPA}$.
Otherwise, $(\theta_{i},\bm{\theta}_{-i}),(\theta_{i}^{\prime},\bm{\theta}_{-i})\in\hat{\bm{\Theta}}$. We apply the same techniques from the proof of Proposition˜2. Without loss of generality, suppose $\theta_{i}<\theta_{i}^{\prime}$. If $\phi^{FPA}((\theta_{i},\bm{\theta}_{-i}))\neq\phi^{FPA}((\theta_{i}^{\prime},\bm{\theta}_{-i}))$, we are done. Otherwise, bidder $i$ loses the auction in both cases. [^26] If bidder $i$ wins the auction in both cases, his payment differs because $b_{i}$ is strictly increasing, and hence the outcome differs. We may then construct a type profile $\bm{\theta}_{-i}^{\prime}\in\hat{\Theta}_{-i}$ such that
$(b_{i}(\theta_{i}),i)\triangleleft\max_{j\neq i}\{(b_{j}(\theta_{j}^{\prime}),j)\}\triangleleft(b_{i}(\theta_{i}^{\prime}),i).$
Then $\phi^{FPA}(\theta_{i},\bm{\theta}_{-i}^{\prime})$ also results in $i$ losing the auction, so $\phi^{FPA}(\theta_{i},\bm{\theta}_{-i})\sim_{\Omega_{i}}\phi^{FPA}(\theta_{i},\bm{\theta}_{-i}^{\prime})$. However, $\phi^{FPA}(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$ results in $i$ winning the auction, so $\phi^{FPA}(\theta_{i},\bm{\theta}_{-i}^{\prime})\neq\phi^{FPA}(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$. We conclude that $\phi^{FPA}$ violates the indistinguishable corners condition at $\{\theta_{i},\theta_{i}^{\prime}\}\times\{\bm{\theta}_{-i},\bm{\theta}_{-i}^{\prime}\}$. Because $\hat{P}$ dominates $P$ and $(\theta_{i},\bm{\theta}_{-i})\sim_{(P,i)}(\theta_{i},\bm{\theta}_{-i}^{\prime})$, we have $(\theta_{i},\bm{\theta}_{-i})\sim_{(\hat{P},i)}(\theta_{i},\bm{\theta}_{-i}^{\prime})$. By Lemma˜1, $(\theta_{i},\bm{\theta}_{-i})\nsim_{(\hat{P},A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$. ∎
Lemma 9 (Equivalent Bidder Information Leakages).
$\sim_{(\hat{P},i)}$ is identical to $\sim_{(P,i)}$ for all bidders $i$.
Proof.
Using Lemma 8, it remains to show that the set of auctioneer privacy violations being the same implies that $P$ and $\hat{P}$ have same set of bidder privacy violations. Once again, since $\hat{P}$ leaks less information to bidders than $P$ by assumption, it suffice to show that every information leakage under $P$ is also a leakage under $\hat{P}$. That is, given $\sim_{(\hat{P},A)}$ is identical to $\sim_{(P,A)}$, we must show that $\sim_{(\hat{P},i)}$ is identical to $\sim_{(P,i)}$ for all bidders $i$. In particular, we must show that for all $(\theta_{i},\bm{\theta}_{-i})$ and $(\theta_{i},\bm{\theta}_{-i}^{\prime})$, if $(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,i)}(\theta_{i},\bm{\theta}_{-i}^{\prime})$ then $(\theta_{i},\bm{\theta}_{-i})\nsim_{(\hat{P},i)}(\theta_{i},\bm{\theta}_{-i}^{\prime})$.
Suppose $(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,i)}(\theta_{i},\bm{\theta}_{-i}^{\prime})$. There are three cases to consider.
  1. 1.
    Case 1: $(\theta_{i},\bm{\theta}_{-i}),(\theta_{i},\bm{\theta}_{-i}^{\prime})\in\hat{\Theta}$. Then because $P$ is minimally bidder informative in $\hat{\Theta}$, we have $(\theta_{i},\bm{\theta}_{-i})\nsim_{\Omega_{i}}(\theta_{i},\bm{\theta}_{-i}^{\prime})$. Hence $(\theta_{i},\bm{\theta}_{-i})\nsim_{(\hat{P},i)}(\theta_{i},\bm{\theta}_{-i}^{\prime})$.
  2. 2.
    Case 2: $(\theta_{i},\bm{\theta}_{-i})\in\hat{\Theta}$ and $(\theta_{i},\bm{\theta}_{-i}^{\prime})\notin\hat{\Theta}$. Consider any other $\theta_{i}^{\prime}\in\hat{\Theta}_{i}$ with $\theta_{i}\neq\theta_{i}^{\prime}$. Then $(\theta_{i}^{\prime},\bm{\theta}_{-i})\in\hat{\Theta}$ so Case 1 applies and $(\theta_{i},\bm{\theta}_{-i})\nsim_{(\hat{P},A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$. However, $(\theta_{i},\bm{\theta}_{-i}^{\prime})\sim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$, because an agent other than $i$ won the object in the descending stage. Since $\hat{P}$ leads to less bidder information leakages, we have $(\theta_{i},\bm{\theta}_{-i}^{\prime})\sim_{(\hat{P},A)}(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$. By Lemma˜1, $(\theta_{i},\bm{\theta}_{-i})\nsim_{(\hat{P},i)}(\theta_{i},\bm{\theta}_{-i}^{\prime})$.
  3. 3.
    Case 3: $(\theta_{i},\bm{\theta}_{-i}),(\theta_{i},\bm{\theta}_{-i}^{\prime})\notin\hat{\Theta}$. If $(\theta_{i},\bm{\theta}_{-i})\nsim_{\Omega_{i}}(\theta_{i},\bm{\theta}_{-i}^{\prime})$ then directly $(\theta_{i},\bm{\theta}_{-i})\nsim_{(\hat{P},i)}(\theta_{i},\bm{\theta}_{-i}^{\prime})$ and we are done. Otherwise, agent $i$ loses the auction in both cases. Furthermore, there exists $\theta_{i}^{\prime}$ such that (without loss of generality in re-labeling $\bm{\theta}_{-i}\bm{\theta}_{-i}^{\prime})$, bidder $i$ would have won if their type was $\theta_{i}^{\prime}$: [^27] Since $(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,i)}(\theta_{i},\bm{\theta}_{-i}^{\prime})$, bidder $i$ is offered the object at some price at least once in at least one of the type profiles.
    $\max_{j\neq i}\{(b_{j}(\theta_{j}),j)\}\triangleleft(b_{i}(\theta_{i}^{\prime}),i)\triangleleft\max_{j\neq i}\{(b_{j}(\theta_{j}^{\prime}),j)\}.$
    Then $(\theta_{i},\bm{\theta}_{-i}^{\prime})\sim_{(\hat{P},A)}(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$, as agent $i$ loses the auction in both cases, but $(\theta_{i},\bm{\theta}_{-i})\nsim_{(\hat{P},A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$, as agent $i$ wins in one case and loses in the other case. By Lemma˜1, $(\theta_{i},\bm{\theta}_{-i})\nsim_{(\hat{P},i)}(\theta_{i},\bm{\theta}_{-i}^{\prime})$.
We conclude that $\hat{P}$ reveals the same information to the auctioneer and reveals the same information to each bidder as $P$, hence $P$ lies on the Privacy Frontier. ∎
PROOF OF PROPOSITION 4
Proof.
Let $P$ denote the descending-holdout protocol, with $k$ denoting the holdout bidder. We use $\phi$ to refer to $\phi^{FPA}$. Suppose $\hat{P}$ privacy dominates $P$. We first use Lemma˜5 to show $\sim_{(\hat{P},A)}$ is identical to $\sim_{(P,A)}$. We then use Lemma˜6 to show $\sim_{(\hat{P},i)}$ is identical to $\sim_{(P,i)}$ for all bidders $i$.
Claim 1: $\sim_{(\hat{P},A)}$ is identical to $\sim_{(P,A)}$.
It suffices to check the conditions of Lemma˜5.
For bidder $k$, if $(\theta_{k},\bm{\theta}_{-k})\nsim_{(P,A)}(\theta_{k}^{\prime},\bm{\theta}_{-k})$ then $\phi(\theta_{k},\bm{\theta}_{-k})\neq\phi(\theta_{k}^{\prime},\bm{\theta}_{-k})$. (This is because if $k$ wins in both cases, the price she pays is different. If $k$ loses the auction in both cases, the auctioneer only learns that $k$ loses the auction.)
For bidder $i\neq k$, the auctioneer perfectly learns $\theta_{i}$ in protocol $P$. The logic is then similar to that of the sealed-bid first-price auction. Without loss of generality, suppose $\theta_{i}<\theta_{i}^{\prime}$. If $\phi(\theta_{i},\bm{\theta}_{-i})\neq\phi(\theta_{i}^{\prime},\bm{\theta}_{-i})$, we are done. Otherwise, bidder $i$ loses the auction in both cases. [^28] If bidder $i$ wins the auction in both cases, his payment differs because $b_{i}$ is strictly increasing, and hence the outcome differs. We may then construct a type profile $\bm{\theta}_{-i}^{\prime}\in\hat{\Theta}_{-i}$ such that
$(b_{i}(\theta_{i}),i)\triangleleft\max_{j\neq i}\{(b_{j}(\theta_{j}^{\prime}),j)\}\triangleleft(b_{i}(\theta_{i}^{\prime}),i).$
Then $\phi(\theta_{i},\bm{\theta}_{-i}^{\prime})$ also results in $i$ losing the auction, so $\phi(\theta_{i},\bm{\theta}_{-i})\sim_{(P,i)}\phi(\theta_{i},\bm{\theta}_{-i}^{\prime})$. However, $\phi^{FPA}(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$ results in $i$ winning the auction, so $\phi(\theta_{i},\bm{\theta}_{-i}^{\prime})\neq\phi(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$. This satisfies condition (b).
By Lemma˜5, $\sim_{(\hat{P},A)}$ is identical to $\sim_{(P,A)}$.
Claim 2: $\sim_{(\hat{P},i)}$ is identical to $\sim_{(P,i)}$ for all $i\in N$.
By Claim 1, $\sim_{(\hat{P},A)}$ is identical to $\sim_{(P,A)}$. Hence it suffices to check the conditions of Lemma˜6.
For bidders $i\neq k$, bidder $i$ learns nothing from the protocol except the outcome (he has only one information set), so if $(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,i)}(\theta_{i},\bm{\theta}_{-i}^{\prime})$ we have $(\theta_{i},\bm{\theta}_{-i})\nsim_{\Omega_{i}}(\theta_{i},\bm{\theta}_{-i}^{\prime})$.
For the holdout bidder $i=k$, suppose $(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,i)}(\theta_{i},\bm{\theta}_{-i}^{\prime})$. Let $j:=\operatorname*{\arg\!\max}_{l\neq i}\{(b_{l}(\theta_{l}),l)\}$ and $j^{\prime}:=\operatorname*{\arg\!\max}_{l\neq i}\{(b_{l}(\theta_{l}^{\prime}),l)\}$ denote the highest other bidders in $\bm{\theta}_{-i}$ and $\bm{\theta}_{-i}^{\prime}$, respectively. Then $(b_{j}(\theta_{j}),j)\neq(b_{j^{\prime}}(\theta_{j^{\prime}}^{\prime}),j^{\prime})$, and without loss of generality we may consider $(b_{j}(\theta_{j}),j)\triangleleft(b_{j^{\prime}}(\theta_{j^{\prime}}^{\prime}),j^{\prime})$.
If $(b_{j^{\prime}}(\theta_{j^{\prime}}^{\prime}),j^{\prime})\triangleleft(b_{i}(\theta_{i}),i)$ then $(\theta_{i},\bm{\theta}_{-i})\sim_{(P,i)}(\theta_{i},\bm{\theta}_{-i}^{\prime})$, because agent $i$ wins the item in both cases before reaching the reserve price. If $(b_{j}(\theta_{j}),j)\triangleleft(b_{i}(\theta_{i}),i)\triangleleft(b_{j^{\prime}}(\theta_{j^{\prime}}^{\prime}),j^{\prime})$ then $(\theta_{i},\bm{\theta}_{-i})\nsim_{\Omega_{i}}(\theta_{i},\bm{\theta}_{-i}^{\prime})$, because agent $i$ wins in $(\theta_{i},\bm{\theta}_{-i})$ and loses in $(\theta_{i},\bm{\theta}_{-i}^{\prime})$. The remaining case is when $(b_{i}(\theta_{i}),i)\triangleleft(b_{j}(\theta_{j}),j)$. Because $(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,i)}(\theta_{i},\bm{\theta}_{-i}^{\prime})$, bidder $i$ receives a different set of questions in $(\theta_{i},\bm{\theta}_{-i})$ and $(\theta_{i},\bm{\theta}_{-i}^{\prime})$. That is, the auction ends later in $(\theta_{i},\bm{\theta}_{-i})$ than in $(\theta_{i},\bm{\theta}_{-i}^{\prime})$. Hence there exists $\theta_{i}^{\prime}\in\Theta_{i}$ such that
$(b_{j}(\theta_{j}),j)\triangleleft(b_{i}(\theta_{i}^{\prime}),i)\triangleleft(b_{j^{\prime}}(\theta_{j^{\prime}}^{\prime}),j^{\prime}).$
We have $(\theta_{i},\bm{\theta}_{-i}^{\prime})\sim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$, because bidder $i$ loses in both cases, so the auction ends before the auctioneer distinguishes $\theta_{i}$ from $\theta_{i}^{\prime}$. However, $(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$, because bidder $i$ loses the auction with value $\theta_{i}$ and wins the auction with value $\theta_{i}^{\prime}$. This satisfies condition (c) of the lemma.
We conclude that $\hat{P}$ reveals the same information to auctioneers and to bidders as $P$, hence $P$ lies on the privacy frontier. ∎
Second-Price Auctions
PROOF OF LEMMA 4
Proof.
We will construct an indistinguishable corners condition violation for bidder $i$. Consider any three bidders $i,j,k$ and let $\theta_{i}<\theta_{j}^{\prime}<\theta_{i}^{\prime}<\theta_{j}<\theta_{k}$. Choose $\bm{\theta}_{-ijk}$ with $\theta_{\ell}\leq\theta_{i}$ for all $\ell\neq i,j,k$. Thus, bidder $k$ always wins the auction but their payment depends on bidders $i$ and $j$. Take $\bm{\theta}_{-i}=(\theta_{j},\theta_{k},\bm{\theta}_{-ijk})$ and $\bm{\theta}_{-i}^{\prime}=(\theta_{j}^{\prime},\theta_{k},\bm{\theta}_{-ijk})$.
Then, $\phi(\theta_{i},\bm{\theta}_{-i})=\phi(\theta_{i}^{\prime},\bm{\theta}_{-i})$ as both output bidder $k$ winning at a price of $\theta_{j}$. Next, $\phi(\theta_{i},\bm{\theta}_{-i})\sim_{\Omega_{i}}\phi(\theta_{i},\bm{\theta}_{-i}^{\prime})$ while $\phi(\theta_{i},\bm{\theta}_{-i})\neq\phi(\theta_{i},\bm{\theta}_{-i}^{\prime})$ as bidder $i$ loses in both cases but the final price is different ($b_{j}(\theta_{j})$ versus $b_{j}(\theta_{j}^{\prime})$). Finally, $\phi(\theta_{i},\bm{\theta}_{-i}^{\prime})\neq\phi(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$ as the final price is different ($b_{j}(\theta_{j}^{\prime})$ versus $b_{i}(\theta_{i}^{\prime})$). As such, $\phi^{SPA}$ has an indistinguishable corners condition violation for bidder $i$ at $\{\theta_{i},\theta_{i}^{\prime}\}\times\{\bm{\theta}_{-i},\bm{\theta}_{-i}^{\prime}\}$ but does not have a corners condition violation for bidder $i$ at those type profiles. ∎
PROOF OF PROPOSITION 6
Proof.
In the ascending auction, the auctioneer learns all losers’ values but does not learn the winner’s value. Hence the ascending auction reveals strictly less to the auctioneer than the sealed-bid second-price auction. In both formats, a losing bidder learns only that she lost the auction. In both protocols, the winning bidder learns the second-highest bidder’s value based on their payment. Hence both protocols reveal the same information to bidders (that is, they are both minimally bidder-informative), and the ascending auction privacy dominates the sealed-bid second-price auction. ∎
PROOF OF PROPOSITION 7
Proof.
As described in the main text, the ascending auction protocol is minimally bidder-informative, reveals the value of all losing bidders to the auctioneer, and reveals nothing about the winner’s value to the auctioneer except that the winner’s value is above the second-highest value. It therefore suffices to show that in any minimally bidder-informative protocol $P$ which implements $\phi^{SPA}$, the auctioneer must learn the value of all losing bidders. For brevity, in this proof we use $\phi$ to refer to $\phi^{SPA}$.
Formally, it suffices to show that for any distinct $\theta_{i},\theta_{i}^{\prime}\in\Theta_{i}$ and any $\bm{\theta}_{-i}$ satisfying $\theta_{i},\theta_{i}^{\prime}\leq\max_{j\neq i}\theta_{j}$, we have $(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$. Without loss of generality, we may consider $\theta_{i}<\theta_{i}^{\prime}$. Also note that if $\phi(\theta_{i},\bm{\theta}_{-i})\neq\phi(\theta_{i}^{\prime},\bm{\theta}_{-i})$, then $(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$ because $G$ implements $\phi$. So it remains to consider the case when $\phi(\theta_{i},\bm{\theta}_{-i})=\phi(\theta_{i}^{\prime},\bm{\theta}_{-i})$.
We will find a type profile $\bm{\theta}_{-i}^{\prime}$ that yields a violation of the indistinguishable corners condition at $\{\theta_{i},\theta_{i}^{\prime}\}\times\{\bm{\theta}_{-i},\bm{\theta}_{-i}^{\prime}\}$. By Lemma˜7 and the fact that $P$ is minimally bidder-informative, the auctioneer thus must distinguish between $(\theta_{i},\bm{\theta}_{-i})$ and $(\theta_{i}^{\prime},\bm{\theta}_{-i})$, implying $(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$ as desired. Indeed, consider any $\bm{\theta}_{-i}^{\prime}$ satisfying $(\theta_{i},i)\triangleleft\max_{j\neq i}(\theta_{j}^{\prime},j)\triangleleft(\theta_{i}^{\prime},i)$, with one inequality holding strictly. [^29] The choice of which inequality to hold strictly will depend on the tie-breaking priority of bidder $i$. Then bidder $i$ loses the auction at type profiles $(\theta_{i},\bm{\theta}_{-i})$ and $(\theta_{i},\bm{\theta}_{-i}^{\prime})$, hence $(\theta_{i},\bm{\theta}_{-i})\sim_{\Omega_{i}}(\theta_{i},\bm{\theta}_{-i}^{\prime})$. However, bidder $i$ wins the auction at type profile $(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$, hence $\phi(\theta_{i},\bm{\theta}_{-i}^{\prime})\neq\phi(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$. Combining these three conditions, we conclude that $\phi^{SPA}$ violates the indistinguishable corners condition at $\{\theta_{i},\theta_{i}^{\prime}\}\times\{\bm{\theta}_{-i},\bm{\theta}_{-i}^{\prime}\}$. Hence by Lemma˜7, $(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$. ∎
Definition of Ascending-Join Protocol (Haupt and Hitzig (2025))
Let $\tilde{\bm{\Theta}}\subset\bm{\theta}$ denote the set of potential types remaining. Let $x=\min\{\phi^{SPA}(\theta):\theta\in\tilde{\bm{\Theta}}\}$ denote the tentative outcome, where $X$ is ordered lexicographically by price then winner identity. Given $\tilde{\bm{\Theta}}$, query agents in order $i=1,2,\dots n$ to ask: “Can you rule out outcome $x$ with your type information?”. Formally, the query separates $\hat{\bm{\Theta}}_{i}=\{\tilde{\theta}_{i}\in\tilde{\Theta}_{i}:\forall\;\tilde{\bm{\theta}}_{-i}\in\tilde{\bm{\Theta}}_{-i},\;\phi(\theta_{i},\tilde{\bm{\theta}}_{-i})\neq x\}$ from $\tilde{\bm{\Theta}}_{i}\setminus\hat{\bm{\Theta}}_{i}$. If agent $i$ reports $\theta_{i}\in\hat{\bm{\Theta}}_{i}$, then $\tilde{\bm{\Theta}}$ is updated and the procedure begins again from bidder 1. If no bidder can rule out outcome $x$, then the protocol terminates and chooses outcome $x$.
PROOF OF PROPOSITION 8
Proof.
Let $P_{AJ}$ denote the ascending-join protocol. By Haupt and Hitzig (2025), the ascending-join protocol is maximally contextually private, so there does not exist a protocol $P$ which implements $\phi^{SPA}$ and for which $\mathcal{P}(\sim_{(P,A)})$ is strictly coarser than $\mathcal{P}(\sim_{(P_{AJ},A)})$. [^30] Formally, this follows from the statement in the appendix of Haupt and Hitzig (2025) that $P_{AJ}$ is minimally relatively informative among protocols implementing $\phi^{SPA}$. Hence any protocol $P$ which privacy dominates the ascending-join protocol must satisfy $\mathcal{P}(\sim_{(P_{AJ},A)})=\mathcal{P}(\sim_{(P,A)})$. It therefore suffices to show that if protocol $P$ satisfies $\mathcal{P}(\sim_{(P_{AJ},A)})=\mathcal{P}(\sim_{(P,A)})$, then $P$ must partition bidder types in the same sequence as $P_{AJ}$. This then implies that each bidder observes the same history in $P$ as in $P_{AJ}$. As such, $P$ reveals the same information to bidders as $P_{AJ}$ so $P$ does not privacy dominate $P_{AJ}$. To do this, we show that in order to prevent additional contextual privacy violations that are not produced by ascending-join, each subsequent question must be the next query in the ascending-join protocol.
We say a bidder is active if they are one of the two bidders with minimum index who have not dropped out of the auction. At any point in the ascending-join protocol, we are in one of two cases. Either two bidders are active, or one bidder has just dropped out. We say a bidder $i$ is active at price $p$ if she is one of the two active bidders and the designer has distinguished that $\theta_{i}\geq p$. Any protocol begins with the first two bidders active at $p=0$. We show the auctioneer cannot deviate from the ascending-join protocol without producing an additional contextual privacy violation. For the sake of contradiction, suppose otherwise, so under $P$, the auctioneer does not distinguish any type profiles that she did not distinguish under $P_{AJ}$.
At any stage, the protocol dictates that the auctioneer sends an information set to some agent. We say that the protocol distinguishes values of agent $j$ if she sends a message $I_{j}$ to $j$ for which there exists $\theta_{j},\theta_{j}^{\prime}\in\Theta_{j}$ such that $S(\theta_{j})(I_{j})\neq S(\theta_{j}^{\prime})(I_{j})$. We say that the protocol distinguishes values of agent $j$ strictly above $p$ if she sends a message $I_{j}$ to $j$ for which there exists $\theta_{j},\theta_{j}^{\prime}>p$ with $S(\theta_{j})(I_{j})\neq S(\theta_{j}^{\prime})(I_{j})$. [^31] We define distinguishes values of agent $j$ strictly below $p$ analogously, with $\theta_{j},\theta_{j}^{\prime}<p$. Note that at any point in the protocol, $P$ cannot distinguish values for any bidder who has already dropped out; this produces additional contextual privacy violations not present in $P_{AJ}$. We first prove the result for the case of symmetric $\Theta_{i}=\Theta$ for all $i$, which we may without loss of generality index by $\Theta_{i}=\{1,2,\dots,\bar{\theta}\}$ for all $i$. The case of asymmetric $\Theta_{i}$ follows straightforwardly. [^32] The only changes when $\Theta_{i}$ are asymmetric are as follows. First, to rule out querying a bidder $k>j$ in Case 1 or a bidder $k^{\prime}>k$ in Case 2, we may construct a type profile in which the two active bidders do not have the highest possible value, but which still renders the information from the queried bidder redundant. Second, the incremented price $p+1$ instead becomes $\min\{\theta_{i}\in\Theta_{i}:\theta_{i}>p\}$ for the appropriate bidder $i$.
Case 1: The first case is when two bidders are active. Either both bidders are active at some $p$, or one bidder is active at $p$ and the other bidder is active at $p+1$. Refer to the two active bidders as $i$ and $j$. Consider the next information set in the protocol. First, the protocol cannot distinguish values for any bidder $k>j$, because if $\theta_{i}=\theta_{j}=\bar{\theta}$ (the maximum value), then bidder $k$ is not pivotal, and such a query would distinguish a pair of values that are not distinguished in $P_{AJ}$. Second, for any bidder the auctioneer cannot distinguish values strictly above $p+1$: this violates unconditional winner privacy if all other bidders have values at most $p+1$, unlike the ascending-join protocol. The only remaining possibility is to query one of the two active bidders to distinguish whether their type is weakly greater than $p+1$. This is exactly the subsequent query of the ascending-join protocol.
Case 2: The second case is when some bidder $j$ has just dropped out. Index the remaining active bidder as $i$, where we have $\tilde{\bm{\Theta}}_{i}=\{\theta_{i}:\theta_{i}\geq p\}$ for some price $p$. The next bidder to be queried, under $P_{AJ}$, is $k=\max\{i,j\}+1$. [^33] If there are no remaining bidders to be queried, that is, if $\max\{i,j\}=n$, then the outcome has been determined. Again, any further queries reveals information to the auctioneer that she does not learn in $P_{AJ}$. Then:
  1. 1.
    The auctioneer cannot query a bidder $k^{\prime}>k$ to distinguish her values: if $\theta_{i}=\theta_{k}=\bar{\theta}$, then agent $k^{\prime}$ is never pivotal and this distinguishes types that $P_{AJ}$ does not.
  2. 2.
    The auctioneer cannot query bidder $i$. If all other bidders have value less than $p$, then $i$ must win with a price already determined. In particular, no further information from $i$ is necessary, so any such query distinguishes types in $\bm{\Theta}_{i}$ that $P_{AJ}$ does not.
  3. 3.
    The auctioneer cannot query bidder $k$ to distinguish values strictly below $p+1$. This distinction never affects the outcome, because the auction price is already at least $p$, so such a query distinguishes types that $P_{AJ}$ does not.
  4. 4.
    The auctioneer cannot query bidder $k$ to distinguish values strictly above $p$. If all other bidders have a value of at most $p$, and $\theta_{k}>p$, then $k$ wins the auction regardless. So such a query again distinguishes types that $P_{AJ}$ does not.
The only remaining possibility is for the auctioneer to query bidder $k$ to partition her types into the two sets $\{\theta_{k}:\theta_{k}>p\}$ and $\{\theta_{k}:\theta_{k}\leq p\}$. This is exactly the subsequent query of the ascending-join protocol. ∎
Sufficient Conditions for the Privacy Frontier
PROOF OF LEMMA 5
Proof.
We first show that in describing the type profiles that the auctioneer distinguishes, it suffices to restrict our attention to pairs of type profiles which differ in only one agent’s type.
Lemma 10.
Let $P$ and $\hat{P}$ be two protocols. Suppose that for all $i\in N$, all $\theta_{i},\theta_{i}^{\prime}\in\Theta_{i}$ and all $\bm{\theta}_{-i}\in\bm{\Theta}_{i}$, we have $(\theta_{i},\bm{\theta}_{-i})\sim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$ if and only if $(\theta_{i},\bm{\theta}_{-i})\sim_{(\hat{P},A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$. Then $\sim_{(P,A)}$ and $\sim_{(\hat{P},A)}$ are identical.
Proof.
(Lemma˜10) We may find a sequence $\{\bm{\theta}^{(1)},\bm{\theta}^{(2)},\dots,\bm{\theta}^{(K)}\}$ which transforms $\bm{\theta}$ into $\bm{\theta}^{\prime}$ by altering one agent’s type in each step. That is, $\bm{\theta}^{(1)}=\bm{\theta}$ and $\bm{\theta}^{(K)}=\bm{\theta}^{\prime}$, where for all $k\in\{1,\dots,K-1\}$ there exists agent $i(k)\in N$ and $\theta_{i(k)},\theta_{i(k)}^{\prime}\in\Theta_{i}$ such that $\bm{\theta}^{(k)}=(\theta_{i(k)},\bm{\theta}^{(k)}_{-i(k)})$ and $\bm{\theta}^{(k+1)}=(\theta_{i(k)}^{\prime},\bm{\theta}^{(k)}_{-i(k)})$. In particular, we may find some sequence for which each $i(k)$ is distinct: that is, the sequence alters each agents’ type at most once.
If $\bm{\theta}^{(k)}\sim_{(P,A)}\bm{\theta}^{(k+1)}$ for all $k\in\{1,\dots,K-1\}$ then by transitivity of $\sim_{(P,A)}$ we have $\bm{\theta}\sim_{(P,A)}\bm{\theta}^{\prime}$.
Otherwise, there exists some $k\in\{1,\dots,K-1\}$ for which $\bm{\theta}^{(k)}\nsim_{(P,A)}\bm{\theta}^{(k+1)}$. As the protocol $P$ can only query one bidder at a time, in the set of all pairs $\bm{\theta}^{(k)},\bm{\theta}^{(k+1)}$ for which $\bm{\theta}^{(k)}\nsim_{(P,A)}\bm{\theta}^{(k+1)}$, there exists some pair that protocol $P$ could distinguish at the earliest point in time (i.e., earliest in the predecessor order), because each information set in a protocol only distinguishes the types of one agent. Let $k$ index this pair. Before distinguishing these profiles, the auctioneer could not distinguish $\bm{\theta}$ from $\bm{\theta}^{(k)}$ and could not distinguish $\bm{\theta}^{(k+1)}$ from $\bm{\theta}^{\prime}$. This question distinguishes $\bm{\theta}^{(k)}$ from $\bm{\theta}^{(k+1)}$ and distinguishes no other pairs in the sequence, hence the auctioneer can distinguish $\bm{\theta}$ and $\bm{\theta}^{\prime}$ after this question. That is, $\bm{\theta}\nsim_{(P,A)}\bm{\theta}^{\prime}$.
Suppose $P$ and $\hat{P}$ satisfy the conditions of the lemma. Then for all $k\in\{1,\dots,K-1\}$ we have $\bm{\theta}^{(k)}\sim_{(P,A)}\bm{\theta}^{(k+1)}$ if and only if $\bm{\theta}^{(k)}\sim_{(\hat{P},A)}\bm{\theta}^{(k+1)}$. Hence $\bm{\theta}\sim_{(P,A)}\bm{\theta}^{\prime}$ if and only if $\bm{\theta}\sim_{(\hat{P},A)}\bm{\theta}^{\prime}$. ∎
We may then prove Lemma˜5 because the given pairs of types that the auctioneer must distinguish are necessarily pinned down by pairs of type profiles that the bidder cannot distinguish between. To reveal less information to the auctioneer requires revealing more information to the bidder. Formally:
$\hat{P}$ privacy dominates $P$, so $(\theta_{i},\bm{\theta}_{-i})\sim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$ implies $(\theta_{i},\bm{\theta}_{-i})\sim_{(\hat{P},A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$. Hence it suffices to show that $(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$ implies $(\theta_{i},\bm{\theta}_{-i})\nsim_{(\hat{P},A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$. There are three cases corresponding to which of the conditions in the Lemma statement hold:
  • In case (a), $(\theta_{i},\bm{\theta}_{-i})\nsim_{(\hat{P},A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$ because $\hat{P}$ implements $\phi$.
  • In case (b), $(\theta_{i},\bm{\theta}_{-i}^{\prime})\nsim_{(\hat{P},A)}(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$ because $\hat{P}$ implements $\phi$ and $\phi(\theta_{i},\bm{\theta}_{-i}^{\prime})\neq\phi(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$. Furthermore, $(\theta_{i},\bm{\theta}_{-i})\sim_{(\hat{P},i)}(\theta_{i},\bm{\theta}_{-i}^{\prime})$ because $\hat{P}$ privacy dominates $P$. By Lemma˜1, $(\theta_{i},\bm{\theta}_{-i})\nsim_{(\hat{P},A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$.
  • Case (c) is identical to case (b) with the roles of $\theta_{i}$ and $\theta_{i}^{\prime}$ reversed.
We have now shown that $\sim_{(P,A)}$ and $\sim_{(\hat{P},A)}$ are identical for any pair of type profiles that differ in only one agent’s type. By Lemma˜10, we conclude that $\sim_{(P,A)}$ and $\sim_{(\hat{P},A)}$ are identical everywhere. ∎
PROOF OF LEMMA 6
Proof.
Similar to the proof of Lemma 5, the proof of Lemma˜6 follows because a fixed set of auctioneer violations pins down a particular set of bidder violations.
$\hat{P}$ privacy dominates $P$, so $(\theta_{i},\bm{\theta}_{-i})\sim_{(P,i)}(\theta_{i},\bm{\theta}_{-i}^{\prime})$ implies $(\theta_{i},\bm{\theta}_{-i})\sim_{(\hat{P},i)}(\theta_{i},\bm{\theta}_{-i}^{\prime})$. Hence it suffices to show that $(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,i)}(\theta_{i},\bm{\theta}_{-i}^{\prime})$ implies $(\theta_{i},\bm{\theta}_{-i})\nsim_{(\hat{P},i)}(\theta_{i},\bm{\theta}_{-i}^{\prime})$. [^34] In the other case, if any two type profiles $\bm{\theta}$ and $\bm{\theta}^{\prime}$ have $\theta_{i}\neq\theta_{i}^{\prime}$, then $\bm{\theta}\nsim_{(P,i)}\bm{\theta}^{\prime}$ and $\bm{\theta}\nsim_{(\hat{P},i)}\bm{\theta}^{\prime}$, because $i$ has different types in the two profiles.
In case (a), $(\theta_{i},\bm{\theta}_{-i})\nsim_{(\hat{P},i)}(\theta_{i},\bm{\theta}_{-i}^{\prime})$ because bidder $i$ can distinguish the two distinct outcomes. In case (b), Lemma˜1 implies $(\theta_{i},\bm{\theta}_{-i})\nsim_{(\hat{P},i)}(\theta_{i},\bm{\theta}_{-i}^{\prime})$. Case (c) is identical to case (b), with $\bm{\theta}_{-i}$ and $\bm{\theta}_{-i}^{\prime}$ swapped. ∎
Pruning and Incentive-Compatibility
PROOF OF PROPOSITION 9
Proof.
First, auctioneer privacy. By definition of $\sim_{(P,A)}$, we have $\bm{\theta}\sim_{(P,A)}\bm{\theta}^{\prime}$ is equivalent to $o_{A}(P,\bm{\theta})=o_{A}(P,\bm{\theta}^{\prime})$. Pruning leaves strategies unchanged and does not remove the on-path history $o_{A}(P,\bm{\theta})$, so $o_{A}(P,\bm{\theta})=o_{A}(P,\bm{\theta}^{\prime})$ holds if and only if $o_{A}(\tilde{P},\bm{\theta})=o_{A}(\tilde{P},\bm{\theta}^{\prime})$. However, $o_{A}(\tilde{P},\bm{\theta})=o_{A}(\tilde{P},\bm{\theta}^{\prime})$ is in turn equivalent to $\bm{\theta}\sim_{(\tilde{P},A)}\bm{\theta}^{\prime}$. Thus $\bm{\theta}\sim_{(P,A)}\bm{\theta}^{\prime}$ if and only if $\bm{\theta}\sim_{(\tilde{P},A)}\bm{\theta}^{\prime}$.
Next, bidder privacy. Suppose $\bm{\theta}\sim_{(P,i)}\bm{\theta}^{\prime}$. This happens if and only if:
  1. 1.
    $\theta_{i}=\theta^{\prime}_{i}$;
  2. 2.
    $\phi(\bm{\theta})\sim_{\Omega_{i}}\phi(\bm{\theta^{\prime}})$;
  3. 3.
    bidder $i$ observes the same sequence of information sets and takes the same actions along the path of play of $\bm{\theta}$ and $\bm{\theta}^{\prime}$ under protocol $P$.
Since conditions one and two do not depend on the game $G$ or $\tilde{G}$ being played, it suffice to focus on condition three. Let $\{h_{k}(\bm{\theta})\}_{k}$ be the sequence of histories at type profile $\bm{\theta}$ and let $\{h_{k}(\bm{\theta})\}_{k:i(h_{k})=i}$ be the subsequence of histories for which bidder $i$ is called to play. Then, condition three can be stated as
$\{I(h_{k}(\bm{\theta}))\}_{k:i(h_{k})=i}=\{I(h_{k}(\bm{\theta}^{\prime}))\}_{k:i(h_{k})=i}.$ (1)
For any on-path histories $h,h^{\prime}\in\tilde{H}$, we have $I(h)=I(h^{\prime})\Leftrightarrow\tilde{I}(h)=\tilde{I}(h^{\prime})$, as $\tilde{I}$ is simply a restriction of $I$ to on-path histories. Thus, 1 is equivalent to
$\{\tilde{I}(h_{k}(\bm{\theta}))\}_{k:i(h_{k})=i}=\{\tilde{I}(h_{k}(\bm{\theta}^{\prime}))\}_{k:i(h_{k})=i}$ (2)
which is then equivalent to bidder $i$ observing the same sequence of information sets along the path of play of $\bm{\theta}$ and $\bm{\theta}^{\prime}$ under $\tilde{P}$. Since the strategies $S$ are unchanged under pruning, the actions bidder $i$ plays are similarly unchanged. Thus, if bidder $i$ observes the same sequence of information sets and takes the same actions along the path of play of $\bm{\theta}$ and $\bm{\theta}^{\prime}$ under protocol $P$, they will also observe the same sequence of information sets and takes the same actions along the path of play of $\bm{\theta}$ and $\bm{\theta}^{\prime}$ under protocol $\tilde{P}$. ∎
PROOF OF PROPOSITION 10
Proof.
Suppose $\phi$ is dominant-strategy incentive compatible. Towards a contradiction, suppose $P$ is not dominant-strategy incentive compatible. Then, there exists $i,\theta_{i},\bm{\theta_{-i}},\sigma_{i}^{\prime}$ such that
$u_{i}(G(S_{i}(\theta_{i}),\bm{S}_{-i}(\bm{\theta}_{-i})),\theta_{i})<u_{i}(G(\sigma_{i}^{\prime},\bm{S}_{-i}(\bm{\theta}_{-i})),\theta_{i}).$ (3)
As $G$ is pruned, the sequence of histories reached under $(\sigma_{i}^{\prime},\bm{S}_{-i}(\bm{\theta}_{-i}))$ must have been reachable via $\bm{S}(\bm{\theta}^{\prime})$ for some $\bm{\theta}^{\prime}$. Without loss, we can take $\bm{\theta}^{\prime}_{-i}=\bm{\theta}_{-i}$; furthermore, $\sigma_{i}^{\prime}=S_{i}(\theta_{i}^{\prime})$. As $P$ implements, $\phi$, we have that
[…]
Plugging this back into (3) gives
$u_{i}(\phi(\theta_{i},\bm{\theta}_{-i}),\theta_{i})<u_{i}(\phi(\theta_{i}^{\prime},\bm{\theta}_{-i}),\theta_{i})$
which contradicts $\phi$ being dominant-strategy incentive compatible. ∎